### v1

# Fundamental limits of weak learnability in high-dimensional multi-index models

Abstract
Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural networks. This paper examines the theoretical boundaries of learnability in this hypothesis class, focusing particularly on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples is $n=\alpha d$ is proportional to the covariate dimension $d$. Our findings unfold in three parts: (i) first, we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!>\!0$; (ii) second, in the case where the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace consisting of directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$. The critical threshold $\alpha_{c}$ marks the presence of a computational phase transition, in the sense that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$. In a limited but interesting set of really hard directions —akin to the parity problem— $\alpha_{c}$ is found to diverge. Finally, (iii) we demonstrate that interactions between different directions can result in an intricate hierarchical learning phenomenon, where some directions can be learned sequentially when coupled to easier ones. Our analytical approach is built on the optimality of approximate message-passing algorithms among first-order iterative methods, delineating the fundamental learnability limit across a broad spectrum of algorithms, including neural networks trained with gradient descent.
1 Introduction
A fundamental property of neural networks is their ability to learn features from data and adapt to relevant structures in high-dimensional noisy data. However, our mathematical understanding of this mechanism remains limited. A popular model for studying this question is multi-index models. Multi-index functions are a class of statistical models encoding the inductive bias that the relevant directions for prediction depend only on a low-dimensional subspace of the covariates ${\bf x}\in\mathbb{R}^{d}$:
$y=g({{\bm{{W}}}}{\bm{x}}),\qquad{{\bm{{W}}}}\in\mathbb{R}^{p\times d},\quad k\in[p].$ (1)
They define a rich class of hypotheses (Li, 1991), containing many widely studied functions in the statistical learning and theoretical computer science literature. The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z)\!=\!z\!+\!\sqrt{\Delta}Z$, with $Z\!\sim\!{\cal N}(0,1)$. In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, the terminology single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is employed when the link function is known and only the weights are learned. In this work, we use both interchangeability., $g(z)$ can be an arbitrary, possibility stochastic, link function, including some widely studied problems such as phase retrieval $g(z)\!=\!z^{2}$ and the perceptron $g(z)={\rm sign}(z)$. Some example of multi-index models include the Two-layer neural network ($p>1$), given by a link function $g(z_{1},\dots,z_{p})=\sum_{k=1}^{p}a_{k}\sigma(z_{k})$; Polynomial functions ($p>1$), given by any linear combination of products of $z_{1},\dots,z_{p}$, for example $g(z_{1},z_{2},z_{3})=1+z_{1}^{2}z_{2}^{2}-2z_{1}+z_{2}z_{3}^{3}$; and s-sparse parities ($p>1$) that can be embedded in a multi-index model by taking a link function $g(z_{1},\dots,z_{p})=\prod_{k\in I}{\rm sign}(z_{k})$ for any subset $I\subset[p]$ of size $|I|=s\leq p$. Training a multi-index model typically translates to a non-convex optimization problem. Several authors have thus used multi-index models as a test-bed for understanding the behavior of neural nets and gradient-descent in non-convex, high-dimensional contexts, e.g. (Saad and Solla, 1995; 1996; Arous et al., 2021; Abbe et al., 2022; 2023; Veiga et al., 2022; Ba et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2023; Damian et al., 2023; Bietti et al., 2023; Moniri et al., 2023; Berthier et al., 2024).
To serve as a useful benchmark, it is necessary to have an idea of the fundamental limit of learnability in such models. This translates to the question of how many observations from model Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability. This can be studied both statistically (within the class of all, including exponential, algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms). In the single-index case ($p=1$), weak learnability has been heavily studied under probabilistic assumptions for the weights and data distribution (e.g. i.i.d. Gaussian or uniformly in the sphere). The statistical threshold for learnability has been characterised by Barbier et al. (2019) when the covariate dimension $d$ is large. Optimal computational thresholds for the class of first-order iterative algorithms were derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020a) in the same regime. Similar results were also proven for other computational models: Damian et al. (2022) provided a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. Damian et al. (2024b) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
Aside from the particular case of committee machines (Aubin et al., 2019; Diakonikolas et al., 2020; Goel et al., 2020; Chen et al., 2022), results for general multi-index models $p>1$ are scarce, as they crucially depend on the way different directions are coupled by the link function. The goal of the present work is precisely to close this gap for the class of Gaussian multi-index models in the proportional, high-dimensional regime, and to provide a classification of how hard it is to learn feature directions from data in multi-index models. More precisely, our main contributions are:
  • We analyse the computational limits of weak learnability (as defined in Def. 2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\to\infty$ at fixed ratio $\alpha=\nicefrac{{n}}{{d}}$. Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural networks trained with gradient descent.
  • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples in multi-index models: trivial directions can be learned with a single AMP iteration for any $\alpha>0$. When no such direction exists, we show there might be easy directions requiring $O(\log d)$ iterations and at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition.  [^2] Note that $\alpha_{c}$ may or may not coincide with the statistically optimal threshold. In no way do we imply the existence or absence of a computational-to-statistical gap. We provide a sharp formula for the critical sample complexity $\alpha_{c}$ above which easy directions are weakly learnable. No efficient iterative algorithm can succeed for $\alpha<\alpha_{c}$. The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!O(d)$ samples.
  • We discuss how interactions among directions can lead to hierarchical learning phenomena. Interestingly, even hard directions can be learned when coupled to easier ones through a hierarchical process, which is reminiscent of the staircase phenomenon for one-pass SGD (Abbe et al., 2023). However, the class of functions that are weakly learnable hierarchically with AMP differs from SGD and is, in particular, considerably larger. We fully characterise this class of Hierarchical Iterative Denoising Staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
Further related work —
With its origins on the classical projection pursuit method (Friedman and Tukey, 1974; Friedman and Stuetzle, 1981), there is an extensive literature dedicated to designing and analysing efficient algorithms to train multi-index models models, such as Isotronic Regression for single-index (Brillinger, 1982; Kalai and Sastry, 2009; Kakade et al., 2011) and Sliced Inverse Regression in the multi-index case (Dalalyan et al., 2008; Yuan, 2011; Fornasier et al., 2012; Babichev and Bach, 2018).
The case $p=1$ has seen a lot of interest recently. In terms of gradient descent, Arous et al. (2021) has shown that if the link function $g$ is known, one-pass SGD achieves weak recovery in $n=\Theta(d^{\ell-1})$ steps, where $\ell$, known as the information exponent, is the first non-zero Hermite coefficient of the link function $g$. In the non-parametric setting —where $g$ is unknown— Berthier et al. (2024) has shown that a large-width two-layer neural network trained under one-pass SGD can learn an ”easy” single-index target ($\ell=1$) in $n=\Theta(d)$ steps. This is to be contrasted with full-batch GD, which can achieve weak recovery in $\Theta(1)$ steps with sample complexity $n=\Theta(d)$ even for particular problems with $\ell>1$ (Dandi et al., 2024b) (which is, indeed, the statistical optimal rate (Barbier et al., 2019; Mondelli and Montanari, 2019)). Similar results have also been proven for abstract computational models. Damian et al. (2022) has proven a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. More recently, Damian et al. (2024b) has proven a similar result for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$ and hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent. This turns out to be equivalent to the optimal computational weak recovery threshold of (Barbier et al., 2019; Mondelli and Montanari, 2019; Luo et al., 2019; Maillard et al., 2020b).
While the situation is less understood, the multicase has also whitenessed a surge of recent interest. In particular, it has been used to understand the behavior of gradiend descent algorithm in neural networks. For instance, Abbe et al. (2022; 2023) showed that a certain class of staircase functions can be learned by large-width two-layer networks trained under one-pass SGD with sample complexity $n=\Theta(d)$. This is in stark contrast to (embedded) $s$-sparse parities, which require $n\geq d^{s-1}$ samples (Blum et al., 2003). Abbe et al. (2022; 2023) introduced the leap exponent, a direction-wise generalisation of the information exponent for multi-index models, and studied the class of staircase functions which can be efficiently learned with one-pass SGD. Bietti et al. (2023) showed that under a particular gradient flow scheme preserving weight orthogonality (a.k.a. Stiefel gradient flow), training follow a saddle-to-saddle dynamics, with the characteristic time required to escape a saddle given by the leap exponent. Interestingly, it was then shown that the limit discovered in these set of works could be bypassed by slightly different algorithm, for instance by smoothing the landscapes Damian et al. (2024a) or reusing batches multi-time Dandi et al. (2024b). The latter paper, in particular, showing that gradient descent could learn efficiently a larger class of multi-index models that previously believed to be possible. These findings highlight the need of a strict understanding of the limit learnability of these models.
Our approach to establish the limit of computational learnability is based on the study of the approximate message passing algorithm (AMP). Originating from the cavity method in physics Mézard (1989); Kabashima (2008), AMP Donoho et al. (2009), and its generalized version GAMP Rangan (2011) are powerful iterative algorithm to study these high-dimensional setting. These algorithm are widely believed to be optimal between all polynomial algorithms for such high-dimensional problems Zdeborová and Krzakala (2016); Deshpande and Montanari (2015); Bandeira et al. (2018); Aubin et al. (2019); Bandeira et al. (2022). In fact, they are provably optimal among all iterative first-order algorithms Celentano et al. (2020); Montanari and Wu (2022), a very large class of methods that include gradient descent.
While these algorithm were studied in great detail for the single index models, see e.g. Barbier et al. (2019); Aubin et al. (2020b; a), and are at the roots of the spectral method that underline the learnability phase transition in this case Mondelli and Montanari (2019); Maillard et al. (2020b), much less is known in the multi-index case, with the exception of Aubin et al. (2019), who showed that particular instances of multi-index models known as committee machines can be learned with $n=\Theta(d)$ samples.
Different from our approach are (Diakonikolas et al., 2020; Goel et al., 2020; Chen et al., 2022), who focused also on peculiar form of the multi-index models (using combination of Relu) and derived worst case bound (in terms of the hardest possible function). Instead, we focus on the typical-case learnability of a given, explicit, multi-index model.
Notation
We denote by $\mathcal{S}^{+}_{p}$, the cone of positive-semi-definite matrices in $\mathbb{R}^{p}$ and by $\succ$ the associated partial ordering. For ${{\bm{M}}}\in\mathcal{S}^{+}_{p}$, we denote by $\sqrt{M}$ the matrix-square root. |||||$\cdot$||| | ⋅ | | denotes the operator norm and $||\cdot||_{F}$ the Frobenius norm.
2 Settings and definitions
As motivated in the introduction, our main focus in this work will be to study subspace identifiability in the class of Gaussian multi-index models.
Definition 1 (Gaussian multi-index models).
Given a covariate ${\bm{x}}\sim\mathcal{N}(0,\nicefrac{{1}}{{d}}{\bm{I}}_{d})$, we define the class of Gaussian multi-index models as likelihoods of the type:
$\mathbb{E}[y|{\bm{x}}]=g({{\bm{{W}}}}^{\star}{\bm{x}})$ (2)
where $g:\mathbb{R}^{p}\to\mathbb{R}$ is the link function and ${{\bm{{W}}}}^{\star}\in\mathbb{R}^{p\times d}$ is a weight matrix with i.i.d. rows ${\bm{w}}^{\star}_{k}\sim\mathcal{N}(0,{\bm{I}}_{d})$. Note that this uniquely defines a joint distribution $p({\bm{x}},y)$ over $\mathbb{R}^{d+1}$.
Given $n$ i.i.d. samples $({\bm{x}}_{i},y_{i})_{i\in[n]}$ drawn as per Definition 1, we are interested in investigating the computational bottlenecks of estimating ${{\bm{{W}}}}^{\star}$ from the samples $({\bm{x}}_{i},y_{i})_{i\in[n]}$. Note that reconstructing ${{\bm{{W}}}}^{\star}$ or a permutation of its rows is equivalent from the perspective of the likelihood Equation 2. Therefore, in this work, we will be interested in weak subspace learnability, which corresponds to obtaining an estimation of the subspace spanned by ${{\bm{{W}}}}^{\star}$ better than a random estimator. This can be defined in an invariant way as follows
Definition 2 (Weak subspace recovery).
Let $V^{\star}\subset\mathbb{R}^{p}$ denote a subspace spanned by vectors representing components along ${{\bm{{W}}}}^{\star}$ such that each $\mathbf{v}\in V^{\star}$ maps to a vector $\mathbf{v}_{d}$ in $\mathbb{R}^{d}$ through the map $\mathbf{v}_{d}=({{\bm{{W}}}}^{\star})^{\top}\mathbf{v}$. Given an estimator $\hat{{{\bm{{W}}}}}\in\mathbb{R}^{p\times d}$ of ${{\bm{{W}}}}^{\star}$, we have weak recovery of a $V^{\star}$ if:
$\underset{{\bm{v}}\in V^{\star},\norm{{\bm{v}}}=1}{\inf}\norm{\frac{\hat{{{\bm{{W}}}}}({{\bm{{W}}}}^{\star})^{\top}{\bm{v}}}{d}}=\Theta_{d}(1).$ (3)
with high probability as $d\to\infty$.
Our main tool for characterising the computational bottlenecks in the Gaussian multi-index problem is an Approximate Message Passing (AMP) Algorithm 1 tailored to our problem. Note that Algorithm 1 only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and the application of a problem dependent function ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$. Therefore, it belongs to the class of first-order algorithms with linear running time on $O(nd+p^{3})$. The key property of AMP is that for a well-chosen ${\bm{g}}_{\rm out}$, it is provably optimal within the class of first-order methods (Celentano et al., 2020). This means that establishing learnability for the optimal AMP Algorithm 1 implies a computational lower bound in the class of first-order methods, which includes popular machine learning algorithms such as gradient descent (SG). For the Gaussian multi-index estimation problem 1, the optimal ${\bm{g}}_{\rm out}$ is simply given by the optimal denoiser of an effective $p$-dimensional problem $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
[…]
where the conditional expectation is defined through the output channel $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$. Here ${\bm{\omega}}$ can be interpreted as an estimate of the pre-activations. The above denoiser can also be derived as an exact high-dimensional approximation of the Belief Propagation (BP) algorithm for the estimation of the marginals of the Bayesian posterior distribution:
$p({{\bm{{W}}}}|{\bm{X}},{\bm{y}})\propto\prod\limits_{i=1}^{n}\delta(y_{i}-g({{\bm{{W}}}}{\bm{x}}_{i}))\prod_{k=1}^{p}\mathcal{N}({\bm{w}}_{k}|0,{\bm{I}}_{d}).$ (5)
Algorithm 1 was derived for Gaussian multi-index models by Aubin et al. (2019), although it has been analysed only for the particular link function $g({\bm{z}})={\rm sign}(\sum_{k\in[p]}{\rm sign}(z_{k}))$, known as the committee machine. Our goal in this work is to leverage Algorithm 1 in order to provide a sharp classification of which link functions $g$ are computationally challenging to learn.
Algorithm 1 Multi-index AMP
  Input: Data ${\bm{X}}\in\mathbb{R}^{n\times d}$, ${\bm{y}}\in\mathbb{R}^{n}$
  Initialize $\hat{{{\bm{{W}}}}}^{t=0}\in\mathbb{R}^{p\times d}$, $\hat{{\bm{C}}}_{k}^{t=0}\in\mathcal{S}_{p}^{+}$ for $k\in[d]$, ${\bm{g}}^{t=0}\in\mathbb{R}^{n\times p}$.
  for $t\leq T$ do
     /* Update likelihood mean and variance
     ${\bm{V}}_{i}^{t}=\sum_{k=1}^{d}X^{2}_{ik}\hat{{\bm{C}}}_{k}\in\mathbb{R}^{p\times p}$;  $\mathbf{\omega}_{i}^{t}=\sum_{k=1}^{d}X_{ik}\hat{{{\bm{{W}}}}}_{k}^{t}-{\bm{V}}_{i}^{t}{\bm{g}}_{i}^{t-1}\in\mathbb{R}^{p}$, $i\in[n]$;
     ${\bm{g}}_{i}^{t}=g_{\rm{\rm out}}(y_{i},{\bm{\omega}}_{i}^{t},{\bm{V}}_{i}^{t})\in\mathbb{R}^{p}$ ;  $\partial{\bm{g}}_{i}^{t}=\partial_{{\bm{\omega}}}g_{\rm{\rm out}}(y_{i},{\bm{\omega}}_{i}^{t},{\bm{V}}_{i}^{t})\in\mathbb{R}^{p\times p}$ ; $i\in[n]$
     /* Update prior first and second moments
     ${\bm{A}}_{k}^{t}=-\sum_{i=1}^{n}X_{ik}^{2}\partial{\bm{g}}^{t}_{i}\in\mathbb{R}^{p\times p}$ ;   $\bm{b}_{k}^{t}=\sum_{i\in[n]}X_{ik}{\bm{g}}_{i}^{t}+{\bm{A}}_{k}^{t}\hat{{{\bm{{W}}}}}_{k}^{t}$;  $k\in[d]$
     $\hat{{{\bm{{W}}}}}_{k}^{t+1}=({\bm{I}}_{p}+{\bm{A}}_{k}^{t})^{-1}\bm{b}_{k}^{t}\in\mathbb{R}^{p}$ ;  $\hat{{\bm{C}}}_{k}^{t+1}=({\bm{I}}_{p}+{\bm{A}}_{k}^{t})^{-1}\in\mathbb{R}^{p\times p}$,  $k\in[d]$
  end for
  Return: Estimators $\hat{{{\bm{{W}}}}}_{\rm amp}\in\mathbb{R}^{p\times d},\hat{{\bm{C}}}_{\rm amp}\in\mathbb{R}^{d\times p\times p}$
Our analysis is based on two remarkable properties that make AMP a particularly useful tool for studying high-dimensional estimation. The first property is that for any $t<O(\log{d})$ ), in the high-dimensional limit $d\to\infty$ the performance of AMP can be tracked without actually running the algorithm.
This result, known as state evolution, makes AMP mathematically tractable in high-dimensions (Bayati and Montanari, 2011). The second property is its optimality with respect to Bayesian estimation. We discuss this further in Appendix A.
For multi-index models, the state evolution equations were derived by Aubin et al. (2020b) and rigorously proven by Gerbelot and Berthier (2023):
Lemma 2.1 (State evolution (Aubin et al., 2019; Gerbelot and Berthier, 2023)).
Let $({\bm{x}}_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d. samples from the multi-index model defined in 1. Consider running Algorithm 1 from initial condition $\hat{{{\bm{{W}}}}}^{0}\in\mathbb{R}^{p\times d}$, such that the initial overlap $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ converges in probability to a limit ${{\bm{M}}}^{0}$ as $d\rightarrow\infty$. Denote by $\hat{{{\bm{{W}}}}}^{t}$ the resulting estimator at time $0\leq t\leq T$. Then, in the high-dimensional limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$ and $T<O(\log{d})$ ), the asymptotic mean-squared error on the label prediction is given by:
$\lim\limits_{n,d\to\infty}\mathbb{E}[(y-g(\hat{{{\bm{{W}}}}}^{t}{\bm{x}}))^{2}]=\mathbb{E}[(Y^{t}-g({\bm{Z}}))^{2}],$ $0\leq t\leq T$ (6)
where the expectation is taken over the following effective estimation process:
$Y^{t}=g\left(({\bm{I}}_{p}-{{\bm{M}}}^{t})^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{{{\bm{M}}}^{t}}^{\nicefrac{{1}}{{2}}}{\bm{\xi}}\right),\qquad{\bm{Z}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ (7)
with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ independently from ${\bm{Z}}$, and ${{\bm{M}}}^{t}$ is given by iterating the following state evolution equations from initial condition ${{\bm{M}}}^{0}$:
${{\bm{M}}}^{t+1}\!$ $=\!F({{\bm{M}}}^{t})\!\coloneqq{G}\!\left(\!\alpha\,\mathbb{E}\Big{[}{\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}\!\Big{]}\right).$ (8)
where $G({\bm{X}})=({\bm{I}}_{p}+{\bm{X}})^{-1}{\bm{X}}$ and the expectation is taken over the effective process Equation 7. Furthermore, ${{\bm{M}}}^{t}$ is constrained to lie in $\mathcal{S}^{+}_{p}$ due to Nishimori-identity (section G.3 in (Aubin et al., 2019)). The limiting overlap $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ converge in probability to ${{\bm{M}}}^{t}$.
3 The trivial subspace ($\alpha_{c}=0$)
The state evolution Equation 8 maps the problem of characterising the computational bottlenecks of first-order methods for high-dimensional Gaussian multi-index models to the study of the deterministic, $p$-dimensional dynamical system ${{\bm{M}}}^{t+1}=F({{\bm{M}}}^{t})$. A starting point is identifying its fixed points and their basins of attraction. In the absence of any prior information on ${{\bm{{W}}}}^{\star}$ aside from its distribution, one cannot do better than taking $\hat{{\bm{w}}}_{k}^{t=0}\sim\mathcal{N}(0,{\bm{I}}_{d})$ with $k\in[p]$ independently at random from the prior. With high-probability, at initialisation, the elements of the overlap matrix $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ are $\Theta_{d}(d^{-\nicefrac{{1}}{{2}}})$ numbers. The asymptotic overlap for an uninformed initial condition is thus identically zero ${{\bm{M}}}^{0}=0$, a null-rank matrix. If ${{\bm{M}}}^{0}=0$ is not a fixed point, then ${{\bm{M}}}^{1}\succ 0$, implying the weak recovery of a subspace of dimension $k={\rm rank}({{\bm{M}}}^{1})>0$ with just a single step of AMP.
Lemma 3.1 (Existence of uninformed fixed point).
${{\bm{M}}}=0\in\mathbb{R}^{p\times p}$ is a fixed point of Equation 8 if and only if the following condition holds almost surely over $Y$:
${\bm{g}}_{\rm out}(y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})=\mathbb{E}[{\bm{Z}}|Y=y]=\mathbf{0},$ (9)
This implies that as long as the conditional expectation above is not zero almost surely, AMP weakly learns a non-empty subspace immediately in the first iteration for any number of samples $n=\Theta(d)$. For this reason, we refer to this subspace as a “trivial subspace”:
Definition 3 (Trivial subspace).
Let $H^{\star}_{T}\subseteq\mathbb{R}^{p}$ denote the subspace spanned by $\mathbf{v}\in\mathbb{R}^{p}$ such that:
${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\mathbb{E}[\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle|Y=y]=0,$ (10)
almost surely over $Y$ Equation 7. The trivial subspace $T^{\star}$ is defined as the orthogonal complement of $H^{\star}$.
Note that for single-index models ($p=1$), the condition in Equation 9 reduces to the one derived by (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020b). Interestingly, this is exactly the same condition appearing in Damian et al. (2024b) for weak learnability of single-index models in the SQ model; see eq. (3) therein. To make Lemma 3.1 concrete, let us look at a few examples. A detailed derivation of these examples is discussed in the Appendix C.
  1. 1.
    For single-index models ($p=1$), $T^{*}$ is one dimensional if and only if $g$ is non-even, e.g. $g(z)={\rm He}_{3}(z)$. This follows from requiring that $g_{\rm out}(y,0,1)\neq 0$ for at least one value of $y$. In particular, on any open interval where $g_{\rm out}$ is invertible we have $g_{\rm out}=g^{-1}$.
  2. 2.
    For a linear multi-index model, $g({\bm{z}})=\sum_{i=1}^{p}z_{i}$, $T^{\star}$ is spanned by $\mathbf{1}_{p}\in\mathbb{R}^{p}$ (all-one vector).
  3. 3.
    For a committee $g({\bm{z}})\!=\!\sum_{i=1}^{p}{\rm sign}(z_{i})$, the trivial subspace $T^{\star}$ is again 1-d, spanned by $\mathbf{1}\!\in\!\mathbb{R}^{p}$.
  4. 4.
    For monomials $g({\bm{z}})=z_{1}\dots z_{p}$, the trivial subspace $T^{\star}$ is non-empty if and only if $p=1$.
  5. 5.
    For staircase functions (Abbe et al., 2023):
    $g({\bm{z}})=z_{1}+z_{1}z_{2}+z_{1}z_{2}z_{3}+\dots+z_{1}\dots z_{p}$ (11)
    The trivial subspace is $T^{\star}=\mathbb{R}^{p}$ and is spanned by the canonical basis. In other words, AMP learns all the directions with a single step for any $\alpha>0$.
Interestingly, Lemma 3.1 can also be related to computational models based on queries, such as SQ learning (Kearns, 1998): the denoiser ${\bm{g}}_{\rm out}$ can indeed be interpreted as a non-linear transformation on the labels $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$. From this perspective, the statement on the condition for the existence of a non-empty trivial subspace Equation 10 translates to the following condition:
$\mathbb{E}\left[{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=\mathbb{E}\left[\mathbb{E}[(\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|Y=y])^{2}\right]\neq 0,$ (12)
where ${\bm{v}}\in T^{\star}$. Indeed, this can be interpreted as a statistical query of the type $\mathbb{E}[\varphi(y)\psi(x)]$ with label pre-processing $\varphi={\bm{g}}_{\rm out}$. The fact that this linear correlation in the transformed labels is non-vanishing implies that one can weakly recover $\mathbf{v}$ through a tailored spectral method (Mondelli and Montanari, 2019). In fact the denoiser ${\bm{g}}_{\rm out}$ is the optimal such transformation in the sense that when ${\bm{g}}_{\rm out}$ fails to obtain a linear correlation along $\mathbf{v}$, i.e when $\mathbf{v}\in H^{\star}$, then no transformation can:
Lemma 3.2.
For any $\mathbf{u}\in H^{\star}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{u}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=0$ (13)
Note that this is a non-asymptotic statement about the denoising function ${\bm{g}}_{\rm out}$, and the expectation is with respect to the distribution of the labels $y$ (not the effective problem).
Proof.
The above is a direct consequence of the Tower law of expectation. Specifically, we have:
$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right].$
The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}W^{\star}\mathbf{v}=\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$. ∎
This implies that when $T^{\star}$ is empty, running one-pass SGD after applying any transformation $\mathcal{T}$ fails to obtain weak recovery with $O(d)$ samples. The situation can be very different when reusing batches: Dandi et al. (2024b) showed that full batch GD implicitly applies a transformation $\mathcal{T}$ to the labels, in a mechanism similar to the one discussed above. However, this transformation may not be optimal. For instance, while AMP learns the trivial subspaces of both $g(z)={\rm He}_{3}(z)$ and staircase functions (11) in a single-step, GD requires 2-steps for the first and $p$ steps for the latter (Dandi et al., 2024b).
4 Computational phase transitions ($\alpha_{c}>0$)
What happens when the trivial subspace $T^{\star}$ from Definition 3 is empty? In this section, we discuss precisely this case, showing that for some link functions $g$ there exists a critical sample complexity threshold $\alpha_{c}>0$ above which some directions become learnable by iterating AMP. To contrast with the trivial subspace, we refer to these as the easy directions. While the trivial subspace is characterised by the existence of the fixed point ${{\bm{M}}}=0$, the easy subspace will be determined by its stability, which crucially depends on the sample complexity $\alpha=n/d$. The stability of ${{\bm{M}}}=\mathbf{0}$ can be obtained by linearising the state evolution (8) around $\delta{{\bm{M}}}$ with $\norm{\delta{{\bm{M}}}}\approx 0$
$F({{\bm{M}}})\approx\alpha\mathcal{F}(\delta{{\bm{M}}})+\mathcal{O}(\norm{\delta{{\bm{M}}}}^{2})$ (14)
where $\mathcal{F}(\delta{{\bm{M}}})$ is a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$:
$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\mathbb{E}_{Y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,0,{\bm{I}}_{p})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,0,{\bm{I}}_{p})^{\top}\right]\,,$ (15)
This can be summarized as follows:
Lemma 4.1.
(Stability of the uninformed fixed point). If ${{\bm{M}}}=\mathbf{0}\in\mathbb{R}^{p\times p}$ is a fixed point, then it is an unstable fixed point of Equation 8 if and only if $n>\alpha_{c}d$, where the critical sample complexity $\alpha_{c}$ is:
$\frac{1}{\alpha_{c}}=\sup_{{{\bm{M}}}\in\mathcal{S}^{+}_{p}}\norm{\mathcal{F}({{\bm{M}}})}_{F},$ (16)
with $\|\cdot\|_{F}$ the Frobenius norm, and the extremum being achieved at a unique extremizer ${{\bm{M}}}^{\star}\in\mathcal{S}^{+}_{p}$. Moreover, if $\norm{\mathcal{F}({{\bm{M}}})}_{F}=0$, then ${{\bm{M}}}=\mathbf{0}$ is stable a fixed point for any $n=\Theta(d)$.
The above result follows from a generalization of the Perron-Frobenius theorem to the cone $\mathcal{S}^{+}_{p}$; see Section B.2 for a proof. A consequence of this result is that for $\alpha<\alpha_{c}$, even with infinitesimally small informed initialisation AMP will converge back to the trivial fixed point and fail to recover any subspace, while for $\alpha>\alpha_{c}$, it develops a growing overlap along a non-empty subspace starting with arbitrarily small initialization. This marks the onset of a computational phase transition for weak learnability in the first-order methods class. Note that the computational weak learnability threshold $\alpha_{c}$ can be equal to the information-theoretical one (as, for instance, in ”real” phase retrieval (Mondelli and Montanari, 2019; Maillard et al., 2020a)), and by no means we imply the existence (or lack of thereof) of a computational-to-statistical gap (Zdeborová and Krzakala, 2016). The easy subspace spanned by the directions that can be learned with $n>\alpha_{c}d$ samples can be characterised as follows:
Definition 4 (Easy subspace $E^{\star}$).
Let $H^{*}_{E}$ be the subspace spanned by directions $\mathbf{v}\in\mathbb{R}^{p}$ such that:
$\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})\mathbf{v}=0,$ (17)
almost surely over $Y$. We define the easy subspace $E^{\star}$ as the orthogonal complement of $H^{\star}_{E}$.
Lemma 4.1 implies that for $\alpha>\alpha_{c}$ AMP Algorithm 1 initialised from small but non-zero ${{\bm{M}}}^{0}$ will learn some directions in the easy subspace $E^{\star}$. Surprisingly, we can show that for $\alpha>\alpha_{c}$ AMP Algorithm 1 spans the full easy subspace $E^{\star}$ starting from an arbitrarily small but extensive initial correlation.
Theorem 4.2.
Suppose that $T^{\star}={0}$ and the AMP algorithm 1 is initialised at $\hat{{{\bm{{W}}}}}^{0}$ such that ${{\bm{M}}}_{d}^{0}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}=\epsilon{\bm{A}}$, with $\norm{{\bm{A}}}_{F}=1$ and ${\bm{A}}$ being full-rank. Then, with high probability as $d\rightarrow\infty$, for $\alpha\geq\alpha_{c}$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}^{\star}$ for $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{\star}$ is as defined in Lemma 4.1. For $\alpha<\alpha_{c}$ however, ${{\bm{M}}}^{t}_{d}=0$ is asympotically stable i.e. there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}}\rightarrow 0$ as $t\rightarrow\infty$. Furthermore, there exists an $\alpha>0$ and a $\delta>0$ such that is ${{\bm{M}}}_{d}^{t}\succ\delta M_{E^{*}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$ iterations, where $M_{E^{\star}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}$.
Heuristically, random initialization is equivalent to setting $\epsilon=O(\nicefrac{{1}}{{\sqrt{d}}})$ in the above theorem. Therefore, we expect that weak-recovery along $E^{\star}$ can be achieved in $O(\log d)$ iterations. As for Lemma 3.1, the expression for the weak recovery threshold Equation 16 generalises the single-index expression from (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020a). Indeed, for $p=1$, we have $\partial_{\omega}g_{\rm out}(y,0,1)=\mathbb{E}[{\rm He}_{2}(Z)|Y=y]$, and therefore
[…]
which is exactly eq. (11) in Barbier et al. (2019). In other words, as long as the link function $g$ has a non-zero second-order coefficient in the Hermite basis, it is learnable by Algorithm 1 with $n>\alpha_{c}d$ samples. Optimality of AMP then implies that no first-order method can achieve non-vanishing correlation with ${\bm{w}}^{\star}$ when $n<\alpha_{c}d$ as $d\to\infty$.
Figure 1: Numerical illustration of the weak learnability phase transition for the 2-sparse parity $g(z_{1},z_{2})\!=\!{\rm sign}(z_{1}z_{2})$ that has a phase transition at $\alpha_{c}(2)=\pi^{2}/4$. The overlap shows how well the directions $z_{1}$ and $z_{2}$ are recovered. Given the permutation symmetry in (19), we show here and in all the subsequent figures the optimal permutation of the overlap matrix elements reached by AMP. The solid black line is the prediction from the theory. Crosses are averages over $72$ runs of AMP Algorithm 1 with $d\!=\!500$.
We now illustrate Lemma 4.1 in a few examples of interest.
  1. 1.
    The monomial $g({\bm{z}})=z_{1}\dots z_{p}$ with $p>1$ can always be learned with $\alpha>\alpha_{c}(p)$ large enough (Chen and Meka, 2020). For instance, we have $\alpha_{c}(2)\approx 0.5937$, $\alpha_{c}(3)\approx 3.725$ and $\alpha_{c}(4)\approx 4.912$. In Appendix C we derive an analytical formula for $\alpha_{c}(p)$ for arbitrary $p$, and show that $\alpha_{c}(p)\sim p^{1.2}$ for large $p$.
  2. 2.
    The embedding of the sparse parity functions:
    $g({\bm{z}})=\prod_{k=1}^{p}{\rm sign}(z_{k})\,.$ (19)
    As this is invariant under permutations of the indices, Lemma 4.1 implies the existence of a computational phase transition. In Section C.3 we compute analytically the critical value $\alpha_{c}(p)$. For $p=1$ this problem is equivalent to phase retrieval, for which we have $\alpha_{c}(1)=\nicefrac{{1}}{{2}}$. For $p=2$ we show in the appendix that $\alpha_{c}(2)=\nicefrac{{\pi^{2}}}{{4}}$, while $\alpha_{c}(p)=+\infty$ for $p\geq 3$. Figure 1 illustrates this discussion for the 2-sparse parity problem, comparing theoretical prediction Equation 8 with finite-size runs of Algorithm 1.
Figure 2: Hierarchical weak learnability for the staircase function $g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$. (Left): Overlaps with the first direction $|M_{11}|$ ( blue), and with the second and third one $\nicefrac{{1}}{{2}}(M_{22}+M_{33})$ ( red) as a function of the sample complexity $\alpha=\nicefrac{{n}}{{d}}$, with solid lines denoting state evolution curves Equation 8, and crosses/dots finite-size runs of AMP Algorithm 1 with $d=500$ and averaged over $72$ seeds. All other overlaps are zero (black). The two black dots indicate the critical thresholds at $\alpha_{1}\approx 0.575$ and $\alpha_{2}=\pi^{2}/4$. (Right) Corresponding generalization error as a function of the sample complexity. Details on the numerical implementation are discussed in Appendix D.
The sparsity parity function is a classic example of a computationally hard function in theoretical computer science (Blum et al., 1994; 2003). [^3] Strictly speaking, the noiseless parity is efficiently learnable by Gaussian elimination (Blum et al., 1994). Here, we consider computational efficiency in a noise-tolerant sense. Therefore, it should come as no surprise that $\alpha_{c}(p)=\infty$ for $p\geq 3$, providing a concrete example of what a hard subspace (i.e. that cannot be learned with $n=\Theta(d)$ samples) looks like for AMP. Similar to the discussion of trivial directions in Lemma 3.2, we can invoke optimality of AMP to translate our result on the optimal denoiser $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$ to a statement for queries on general label pre-processing transformations:
Lemma 4.3.
For any $\mathbf{v}^{\star}\in H^{\star}_{E}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
$\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\star},\mathbf{x}\rangle)\right]=0$ (20)
Note that this is a non-asymptotic statement about the denoising function ${\bm{g}}_{\rm out}$, and the expectation is with respect to the distribution of the labels $y$ (not the effective problem from Equation 7).
For an AMP-hard function (that is, when $\alpha_{c}=\infty$), this implies that, even after applying any transformation to the output (as allowed in SQ) an algorithm like SGD would require at least $O(d^{2})$ data, or $O(d^{3/2})$ if we allow smoothing as in Damian et al. (2024a).
5 Hierarchical Iterative Denoising
Figure 3: Trajectories of a single finite-size run of AMP with $d=500$ at $\alpha=4$ for $g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$. (Left) Evolution of the overlaps. We display $M_{11}$ in blue, $\nicefrac{{1}}{{2}}(M_{22}+M_{33})$ in red, and the off-diagonal overlaps in black. (Right) Evolution of the generalisation error.
In Section 3 and 4 we have provided a full classification of what class of multi-index models are trivial, easy or hard to learn for a first-order algorithm. This classification was built upon the notion of weak learnability, or the capacity of AMP to learn subspaces starting from a random initial condition. However, it says nothing about about what can be learned after a given subspace is learned, for instance, from possible side information about a particular subspace. We now address this question. Suppose that the estimator $\hat{{{\bm{{W}}}}}^{t}$ has developed an overlap along a subspace belonging to the span of ${{\bm{{W}}}}^{\star}$, resulting in a non-zero overlap ${{\bm{M}}}^{t}\succ 0$. From that point, the main difference with respect to the previous discussion is that the variable ${\bm{\omega}}$ in the linear operator $\mathcal{F}$ defined in Equation 15 becomes non-zero (since it is distributed as ${\bm{\omega}}=\sqrt{{{\bm{M}}}}{\bm{\xi}}$). Crucially, this changes the span of $\mathcal{F}({{\bm{M}}})$ and hence the stability condition in Lemma 4.1. In particular, as we show next, learning some directions might facilitate learning larger subspaces. This phenomenon is reminiscent of the specialisation transition in committee machines (Saad and Solla, 1996; Aubin et al., 2020b) and of the staircase phenomenon for one-pass SGD introduced in (Abbe et al., 2022). Below, we formalise this discussion:
Definition 5.
Let $U\in\mathbb{R}^{p}$ be a subspace of dimensions $k$. We define $H_{E}^{\star}(U)$ to be the subspace spanned by ${\bm{v}}\in U^{\perp}$ such that:
${\bm{v}}^{\top}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}}_{U}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U})=0,$ (21)
almost surely over ${\bm{\xi}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$. We define the “trivially-denoising-coupled” subspace $T^{\star}_{U}$ for $U$ as the orthogonal complement of $H_{T}^{\star}(U)$.
Analogously, let $H^{\star}_{E}(U)$ be the subspace spanned by directions ${\bm{v}}\in U^{\perp}$ such that:
${\bm{v}}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U}){\bm{v}}=0,$ (22)
almost surely over $bxi$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
When ${{\bm{M}}}_{U}$ is additionally a fixed point of $\mathcal{F}_{{{\bm{M}}}}$, we can linearise $\mathcal{F}_{{{\bm{M}}}}$ along the orthogonal complement of $U$. We define the easy-denoising-coupled subspace $E^{\star}_{U}$ for $U$ as the orthogonal complement of $H^{\star}_{E}(U)$. Next, suppose that ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ with $\operatorname{span}({{\bm{M}}}_{U})=U$ is further a fixed-point of $\mathcal{F}_{{{\bm{M}}}}$. Let $\mathcal{F}_{{{\bm{M}}}_{U}}$ denote the linearization of $F({{\bm{M}}})$ along the orthogonal complement $U^{\perp}$ at ${{\bm{M}}}={{\bm{M}}}_{U}$. We define the iterative denoising critical threshold $\alpha_{\text{hid}}({{\bm{M}}}_{U})$ at ${{\bm{M}}}={{\bm{M}}}_{U}$ as:
$\frac{1}{\alpha_{\text{hid}}({{\bm{M}}}_{U})}=\sup_{{{\bm{M}}}^{\perp}\in U^{\perp}}\norm{\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}})}_{F},$ (23)
We denote by ${{\bm{M}}}^{\star}_{U}$ the extremiser achieving the above supremum.
Figure 4: Specialisation as a staircase with $g\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$. The direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is first learned for any $\alpha>0$. The remaining ones are only learned at the specialisation transition occurring at $\alpha\!\approx\!4.3$. We take care of the symmetries as indicated in Appendix D. Crosses denote AMP runs with $d\!=\!100$ averaged over $72$ seeds. The overlap is shown as a function of the sample complexity. Because of the symmetry the elements of $M$ can take one of two values: one on the diagonal and one outside of it. We display $(M_{11}+M_{22}+M_{33})/3$ in blue and $(M_{12}+M_{13}+M_{23})/3$ in red.
The definitions above generalise the notions of trivial and easy subspaces in Definitions 3 and 4 conditionally on a subspace $U$ that has been previously learned, and characterise the directions whose recovery is enabled upon learning the subspace $U$. Concretely, upon developing an initial overlap along $U$, the directions in $T^{\star}_{U},E^{\star}_{U}$ can be recovered analogous to the recovery of $T^{\star},E^{\star}$ starting from random initialization. We formalise this below:
Theorem 5.1.
Let $U\in\mathbb{R}^{p}$ be a fixed subspace. Suppose the AMP algorithm 1 is initialised such that ${{\bm{M}}}_{d}^{0}={{\bm{M}}}_{U}+\epsilon{\bm{A}}$, where $\norm{{\bm{A}}}=1$ and ${{\bm{M}}}_{U}$ is a fixed point of $F({{\bm{M}}})$ in Equation 8 with $\operatorname{span}({{\bm{M}}}_{U})=U$. If $\alpha\geq\alpha_{\text{hid}}({{\bm{M}}}_{U})$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}^{(t)}\succ\delta{{\bm{M}}}^{*}_{U}$ for some $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{*}_{U}$ is defined as in Definition 5. For $\alpha<\alpha_{\text{hid}}({{\bm{M}}}_{U})$ however, there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{\perp}}=0$ as $t\rightarrow\infty$, where ${{\bm{M}}}^{t}_{\perp}$ denotes the projection of ${{\bm{M}}}^{t}$ orthogonal to $U$. Furthermore, suppose that ${\bm{A}}$ is full-rank, then there exists an $\alpha>0$ and a $\delta>0$ such that is ${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}_{E^{\star}_{U}}$, where ${{\bm{M}}}_{E^{\star}_{U}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}_{U}$
When $E^{\star}_{U}$ is non-empty for some subspace $U\subseteq\mathbb{R}^{p}$, we say that the target $y=g({{\bm{{W}}}}{\bm{x}})$ allows learning through Hierarchical Iterative Denoising. A concrete example of a function which displays the hierarchical iterative denoising staircase is a linear combination between hard parity function and an easy polynomial:
$g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ (24)
The sign part is a sparse parity with $p=3$, which cannot be learned with $n=\Theta(d)$ samples, but the quadratic polynomial $z_{1}^{2}$ component in the function allows weak-recovery of the first component i.e. $U={(1,0,0)}$ as long as $\alpha>1/2$. Hence, conditionally on $U$ the effective multi-index model becomes ${\rm sign}(z_{2}z_{3})$, which as discussed in Section 4 is an easy function. Figure 2 illustrates this: first, $z_{1}$ is learned at $\alpha_{1}\approx 0.575$. Then, for larger value when $\alpha>\alpha_{2}$, all directions are learned (see App. (C.4)). Knowing $z_{1}$ makes the hard 3-parity an easy 2-parity problem. We can see the evolution of the overlaps over time in Figure 3.
It is easy to construct a multi-index model where AMP will iterate over any number of such plateaus. For instance, consider the model:
$g({\bm{z}})=z_{1}^{4}+{\rm sign}(z_{1}z_{2}z_{3})+{\rm sign}(z_{3}z_{4}z_{5})+{\rm sign}(z_{5}z_{6}z_{7})+\ldots$ (25)
After the first plateau to learn $z_{1}$, there will be one for $z_{2}$, then for $z_{3}$, for $z_{4}$ and so on.
Another example is provided in Fig. 4 for the neural net $g(z_{1},z_{2},z_{3})\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$. At first only the one-dimensional direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is learned for any $\alpha>0$. This is simply a linear approximation of the function. Only at $\alpha\!\approx\!4.3$ does the network learns the other directions. Known as the the specialisation transition, this has been first discussed by Saad and Solla (1996); Aubin et al. (2020a), and provide a very early example of iterative sequential learning.
Finally, it is important to stress that hierarchical learnable functions are not equivalent to staircase functions (Abbe et al., 2022; 2023). For instance, (25) is learned with $\Theta(d)$, while it requires $O(d^{3})$ samples in the CSQ staircase characterisation.
As we argued in Lemmas 3.2 and 4.3, we believe Hierarchical Iterative Denoising provides the right notion of computational complexity for these models, since it encompasses a larger class of natural algorithms such as batch gradient descent.
6 Conclusion and limitations
This work investigated the computational boundaries of learning high-dimensional Gaussian multi-index models with first-order iterative algorithms. As we hinted in the text (and provided some partial answers), a natural question is how close neural networks are to achieving this optimal performance. Two limitations of our work are: (a) our results hold asymptotically in $d\to\infty$; (b) We assume Gaussian covariates. For the former, numerical experiments support that the predictions are already accurate at fairly moderate sizes. For the latter, a number of universality results justify the scope of high-dimensional Gaussian covariates and extend AMP to more structured designs (Loureiro et al., 2021; Wang et al., 2022; Dudeja et al., 2023; Dandi et al., 2024a).
Acknowledgements
The authors would like to thank Joan Bruna, Theodor Misiakiewicz, Luca Pesce, and Nati Srebro for stimulating discussions. BL acknowledges funding from the Choose France - CNRS AI Rising Talents program. This work was supported by the Swiss National Science Foundation under grants SNSF SMArtNet (grant number 212049) and SNSF OperaGOST (grant number 200390).
Appendix / supplemental material
Appendix A Optimality of AMP
Recall that AMP Algorithm 1 is tailored to estimate the posterior marginals Equation 5. But how efficient is it? A priori, making this comparison requires sampling from the posterior distribution, which is computationally prohibitive in the high-dimensional limit of interest here. The second remarkable property of AMP is that its optimality with respect to the Bayesian posterior Equation 5 can be exactly characterised in the high-dimensional regime.
Lemma A.1 (Bayes-optimal correlation).
Let $(x_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d. samples from the multi-index model defined in 1. Denote by $\hat{{{\bm{{W}}}}}_{\rm bo}=\mathbb{E}[W|X,y]\in\mathbb{R}^{p\times d}$ the mean of the posterior marginals Equation 5. Then, in the high-dimensional asymptotic limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, the asymptotic correlation between the posterior mean and ${{\bm{{W}}}}^{\star}$:
${{\bm{M}}}^{\star}=\lim\limits_{d\to\infty}\mathbb{E}\left[\frac{1}{d}\hat{{{\bm{{W}}}}}_{\rm bo}{{{\bm{{W}}}}^{\star}}^{\top}\right]$ (26)
is the solution of the following $\sup\inf$ problem:
[…]
where $H_{Y}({{\bm{M}}})=\mathbb{E}_{{\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})}[H_{Y}({\bm{m}}|{\bm{\xi}})]$, with $H_{Y}({{\bm{M}}}|{\bm{\xi}})$ the the conditional entropy of the effective $p$-dimensional estimation problem Equation 7.
Note that the state evolution Equation 8 is closely related to the $\sup\inf$ problem in Equation 27. Indeed, remarking that the update function $F$ in Equation 8 is precisely the gradient of the entropy $H_{Y}$ in Equation 7, one can show that state evolution is equivalent to gradient descent in the objective defined by Equation 27 [Aubin et al., 2019]. This non-trivial fact implies that whenever Equation 27 has a single minima, AMP optimally estimates the posterior marginals.
Finally, note that by construction $\hat{{{\bm{{W}}}}}_{\rm bo}$ is the optimal estimator of ${{\bm{{W}}}}^{\star}$ given the data $(X,y)$ (in the MMSE sense). Therefore, the rank of $M^{\star}$ defines the dimension of the statistically optimal subspace reconstruction at sample complexity $\alpha\coloneqq\nicefrac{{n}}{{d}}$. The fact that AMP follows state evolution is proven for such problems in Gerbelot and Berthier [2023]
Appendix B Proofs of the main results
B.1 Linear Approximation
Throughout, we assume that ${\bm{g}}_{\rm out}:\mathbb{R}^{k^{2}+k+1}\rightarrow\mathbb{R}^{k}$ is in $\mathcal{C}^{2}$.
Our analysis relies on the following result:
Lemma B.1.
Let $F({{\bm{M}}})$ be as defined in Lemma 2.1
$F({{\bm{M}}})\approx\alpha\mathcal{F}(\delta{{\bm{M}}})+\mathcal{O}(\alpha\norm{\delta{{\bm{M}}}}_{F}^{2}),$ (28)
where $\norm{}_{F}$ denotes the Frobenius norm and $\mathcal{F}(\delta{{\bm{M}}})$ is a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$:
$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]\,,$ (29)
We proceed through an-entry-wise expansion of each term inside the expection in $F({{\bm{M}}})$ around ${{\bm{M}}}=0$. Since ${{\bm{M}}}_{F}\leq p$ , the first two derivatives of ${\bm{g}}_{\rm out}$ are uniformly bounded in ${\bm{g}}_{\rm out}$ for any fixed $Y,{\bm{\xi}}$. Therefore, applying the multivariate Taylor expansion to ${\bm{g}}_{\rm out}(Y^{t},\Delta_{1}{\bm{\xi}},I+\Delta_{2})$, with $\Delta_{1}=\sqrt{M}$ and $\Delta_{2}=,{\bm{I}}_{p}-{{\bm{M}}}$ yields:
${\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}$
$={\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}+\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)\sqrt{M}{\bm{\xi}}{\bm{\xi}}^{\top}\sqrt{M}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)^{\top}$
$+\langle\partial_{{\bm{V}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)M\rangle{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)^{\top}+{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)\langle M,\partial_{{\bm{V}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)\rangle+\alpha C({\bm{\xi}},y)\mathcal{O}(\norm{M}^{2}_{F}),$
where $C({\bm{\xi}},Y)$ is an integrable function in ${\bm{\xi}},Y$. Since $T^{\star}$ is empty, ${\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)$ vanishes almost surely over $y$. Therefore, using dominated-convergence theorem, we obtain:
$F({{\bm{M}}})$ $=\alpha\mathbb{E}\left[{\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}\right]$
$=\alpha\mathbb{E}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)\sqrt{M}{\bm{\xi}}{\bm{\xi}}^{\top}\sqrt{M}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)^{\top}\right]+\alpha\mathcal{O}(\norm{M}^{2}_{F})$
Since at ${{\bm{M}}}^{t}=0$, ${\bm{\xi}}$ and $Y$ are independent, the above simplifies to:
$F({{\bm{M}}})$ $=\alpha\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]+\alpha\mathcal{O}(\norm{M}^{2}_{F})$
B.2 Proof of Lemma 4.1
Suppose that $M\in\mathcal{S}^{+}$, we have, for any $\mathbf{v}\in\mathbb{R}^{p}$:
$\mathbf{v}^{\top}\mathcal{F}({{\bm{M}}})\mathbf{v}=\alpha\,\mathbb{E}_{y}\left[\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\mathbf{v}\right]\,,$ (30)
since ${{\bm{M}}}\in\mathcal{S}^{+}$, each term inside the expectation is non-negative. Therefore:
$\mathbf{v}^{\top}\mathcal{F}({{\bm{M}}})\mathbf{v}\geq 0$ (31)
Thus, $\mathcal{F}(M)$ is a cone-preserving linear map. The generalised Perron-Frobenius theorem for cone-preserving maps [Krein and Rutman, 1948] then implies that the operator $\mathcal{F}({{\bm{M}}})$ admits a unique eigenvector $M^{*}\in\mathcal{S}^{+}_{p}$ corresponding to the largest eigenvalue $\lambda_{\mathcal{F}}$ such that for any $M\in\mathcal{S}^{+}_{p}\perp M^{*}$:
$\mathcal{F}({{\bm{M}}})<\lambda_{\mathcal{F}}\norm{{{\bm{M}}}}_{F}.$ (32)
Furthermore, all other eigenvalues are strictly smaller than $\lambda_{\mathcal{F}}$. Subsequently, Lemma B.1 implies that $F({{\bm{M}}})$ is stable at ${{\bm{M}}}=0$ if and only if $\alpha\leq\frac{1}{\lambda_{\mathcal{F}}}=\alpha_{c}$.
B.3 Proof of Theorem 4.2
Lemma 2.1 allows us to map the behavior of the variable ${{\bm{M}}}^{t}$ in state-evolution to high-probability statements for the limiting overlaps Suppose that $A\in\mathcal{S}^{+}_{p}$ is a full rank matrix as in Theorem 4.2. Applying Lemmas B.1 and 3.2, we obtain that for ${{\bm{M}}}^{0}=\epsilon A$:
$\norm{{{\bm{M}}}^{t+1}}_{F}\leq\alpha\lambda_{\mathcal{F}}\norm{{{\bm{M}}}^{t}}_{F}+\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F},$ (33)
for some constant $C$. Now, suppose that $\alpha<\frac{1}{\lambda_{F}}$ and $\epsilon<\frac{1}{C}(1-\alpha\lambda_{\mathcal{F}}-\kappa)$, for some $0<\kappa<1-\alpha\lambda_{\mathcal{F}}$ then inductively, we obtain:
$\norm{{{\bm{M}}}^{t+1}}_{F}<(1-\kappa)\norm{{{\bm{M}}}^{t}}_{F},$ (34)
for all $t>0$. Implying that $\norm{{{\bm{M}}}^{t}}_{F}\rightarrow 0$.
Now, suppose that $\alpha>\alpha_{c}$ or equivalently $\alpha\lambda_{F}>1$. Since $A$ is full-rank admits a decomposition:
$\operatorname{tr}(A,{{\bm{M}}}^{\star})=\theta$ (35)
for some $\theta>0$. From Lemma 4.1 and B.1, we obtain:
$\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq\alpha\lambda_{\mathcal{F}}\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (36)
For $\norm{{{\bm{M}}}^{t}}\leq\frac{\theta}{C}(\alpha\lambda_{\mathcal{F}}-1-\kappa)$ for some $\kappa>0$, we inductively obtain:
$\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq(1+\kappa)\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star}),$ (37)
implying that $\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})$ grows as $\omega(e^{\kappa t})$ for sufficiently small $\epsilon$.
Finally, it remains to show that for large enough $\alpha$, ${{\bm{M}}}^{t}$ spans $E^{\star}$.
Notice that for any $\mathbf{v}\in E^{\star}$:
$\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v=\mathbb{E}_{y}\left[(\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p})\mathbf{v})^{2}\right]>0,$ (38)
while for any $A\in\mathcal{S}^{+}_{p}$:
$\mathbf{v}^{\top}\mathcal{F}(A)v\geq 0,$ (39)
since $\mathcal{F}(A)\in\mathcal{S}^{+}_{p}$. Define:
$\nu_{\mathcal{F}}=\inf{\mathbf{v}\in\mathbb{E},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v.$ (40)
Since $\mathbf{v}$ lies in a compact set, $\nu_{\mathcal{F}}>0$. Let $A^{\star}$ denote the projection of $A$ along $E^{\star}$ and consider its eigendecomposition $A^{\star}$:
$A^{\star}=\sum_{i=1}^{p}\lambda_{i}\mathbf{v}_{i}\mathbf{v}^{\top}_{i}$ (41)
We
$\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t+1})\mathbf{v}_{i}\geq\alpha\nu_{\mathcal{F}}\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t})\mathbf{v}_{i}-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (42)
Therefore, for $\alpha\geq\frac{1}{\nu_{\mathcal{F}}}$ and small enough $\epsilon$, ${{\bm{M}}}^{t}$ expands linearly along each $\mathbf{v}_{i}$
B.4 Proof of Theorem 5.1
By assumption, $M_{U}$ is a fixed point of $F(M)$. Therefore, Equation 8 implies that ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}-\sqrt{M}_{U})$ almost surely lies in $U$ and thus $T^{\star}_{U}$ is empty. The proof of Theorem 5.1 then follows that of 4.2 by considering the following linearized operator along $U^{\perp}$
$\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})=\mathrel{\mathop{:}}=\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U}){{\bm{M}}}_{\perp}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U})^{\top}\right]\,,$ (43)
where ${\bm{\omega}}=\sqrt{M}_{U}{\bm{\xi}}$ with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}})$. Similar to Lemma B.1, the linearization follows by noting that $y$ is independent of $\sqrt{{{\bm{M}}}_{\perp}}{\bm{\xi}}$ and $\sqrt{{{\bm{M}}}_{U}+{{\bm{M}}}_{\perp}}=\sqrt{{{\bm{M}}}_{U}}+\sqrt{{{\bm{M}}}_{\perp}}$ since ${{\bm{M}}}_{\perp}\perp{{\bm{M}}}_{U}$.
Appendix C A list of examples of the ${\bm{g}}_{\rm out}$ function
In this section, we list several problems, their corresponding ${\bm{g}}_{\rm out}$, and analysis of their weak recoverability.
C.1 $g(z_{1},z_{2},....,z_{p})=\prod_{j=1}^{p}z_{j}$
For all $p$, ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=0$. Otherwise we have two cases We have two cases. If $p=2$:
$\mathcal{Z}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=\frac{K_{0}(|y|)}{\pi}$ (44)
and
$\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=\begin{bmatrix}|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}-1&y\\ y&|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}-1\\ \end{bmatrix}$ (45)
where $K_{n}(y)$ are the modified Bessel functions of the second kind. If $p>3$ then
$\mathcal{Z}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=\frac{1}{(2\pi)^{p/2}}G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0\end{array}\right)$ (46)
The matrix $\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})$ is diagonal, with identical diagonal elements equal to $f(y)$:
$f(y)=2G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0\end{array}\right)/G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0,1\end{array}\right)-1$ (47)
The alpha critical $\alpha_{c}$ will thus be
$\alpha_{c}=\left[\int_{-\infty}^{\infty}\,{\rm d}yf(y)^{2}\mathcal{Z}_{\rm out}(y)\right]^{-1}$ (48)
C.2 $\frac{1}{p}\sum_{i=1}^{p}z_{i}^{2}$
The function is even, so ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=0$. The trick for the computation is to move to generalised spherical coordinates by choosing $r^{2}=z_{2}^{2}+...z_{p}^{2}$. Recall that the area of the unit sphere in $p-1$ dimensions is
$\frac{2\pi^{\frac{p-1}{2}}}{\Gamma\left(\frac{p-1}{2}\right)}$ (49)
The matrix $\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})$ is diagonal because of parity. All the elements on the diagonal are
$(y-1)^{2}\frac{2^{-\frac{p}{2}}e^{-\frac{py}{2}}(py)^{\frac{p}{2}}}{\Gamma\left(\frac{p}{2}\right)}$ (50)
The critical alpha will be
$\alpha_{c}=\left[\int\,{\rm d}y(y-1)^{2}\frac{2^{-\frac{p}{2}}e^{-\frac{py}{2}}(py)^{\frac{p}{2}}}{\Gamma\left(\frac{p}{2}\right)}\right]^{-1}=\frac{p}{2}$ (51)
Of course this procedure fail at the first step if $p=1$. On the other hand, one can check that the final result is still correct for all positive integer $p$.
C.3 $g(z_{1},\ldots,z_{p})=\operatorname{sign}(\prod_{j=1}^{p}z_{j})$
We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
$d_{ij}({\bm{V}})$ $:=\det\left(\begin{array}[]{cc}v_{i,j}&v_{i,p}\\ v_{j,p}&v_{p,p}\end{array}\right),\quad i,j\in{1,\ldots,p},$ (54)
the function
$E_{s}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}})=1-s\operatorname{erf}\left(\frac{v_{1,p}(\sum_{j=1}^{p-1}v_{j,p}(z_{j}-\omega_{j})-v_{pp}\omega_{p}}{\sqrt{2v_{pp}}}\right),$
the Gaussian probability density
$\rho({\bm{z}}\in\mathbb{R}^{p},{\bm{V}}):=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}\sqrt{\det({\bm{V}})}}e^{-\frac{1}{2}{\bm{z}}^{\top}{\bm{V}}^{-1}{\bm{z}}},$ (55)
and finally
$I_{s}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}}):=E_{s}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\rho\left(\left(\begin{array}[]{ccc}z_{1}-\omega_{2}\\ \ldots\\ z_{p-1}-\omega_{p-1}\end{array}\right)^{\top},v_{pp}\bm{d}({\bm{V}})^{-1}\right).$
Then, defining $\mathbb{R}_{+1}=[0,+\infty)$ and $\mathbb{R}_{-1}=(-\infty,0]$, we have that
$Z_{\rm out}(y,{\bm{\omega}}{\bm{V}})=\sum_{s_{1},\ldots,s_{p}=\pm 1}\frac{\delta_{y,(\prod_{j}s_{j})}}{2}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}}).$ (56)
The components of ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})$ are given by
$({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}}))_{k}$ $=\sum_{s_{1},\ldots,s_{p}=\pm 1}\frac{s_{k}\delta_{y,(\prod_{j}s_{j})}}{2}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\delta(z_{k}-0)$
Stability of the fixed point
We can now focus on studying the stability
$Z_{\rm out}(y,\bm{0},{\bm{I}})=\frac{1}{(2\pi)^{p/2}}\int{\rm d}{\bm{z}}e^{-\nicefrac{{{\bm{z}}^{\top}{\bm{z}}}}{{2}}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}=\frac{1}{2}$ (57)
In order to compute $\nabla_{\omega}{\bm{g}}_{\rm out}(y,\bm{0},{\bm{I}})$, we note at first that
$\frac{1}{(2\pi)^{p/2}}\int{\rm d}{\bm{z}}(z_{k}^{2}-1)e^{-\nicefrac{{{\bm{z}}^{\top}{\bm{z}}}}{{2}}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}=0,\quad\forall k\in\{1,\ldots,p\}.$ (58)
For $p=2$, consider the integral
$\frac{1}{2\pi}\int{\rm d}z_{1}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}$ (59)
$=$ $\frac{2}{2\pi}\left(\delta_{y,1}\int_{0}^{\infty}{\rm d}z_{1}\int_{0}^{\infty}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}+\delta_{y,-1}\int_{0}^{\infty}{\rm d}z_{1}\int_{-\infty}^{0}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}\right)$ (60)
$=\frac{{\rm sign}{(y)}}{\pi}$ (61)
Using (16) we obtain that $\alpha_{c}$ is the inverse of the largest eigenvalues of
$\frac{4}{\pi^{2}}\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix}$ (62)
So $\alpha_{c}=\pi^{2}/4$. For $p\geq 3$, defining (±)$\mathbb{R}^{p}_{(\pm)}=\{{\bm{z}}\in\mathbb{R}^{p}|\operatorname{sign}(z_{1}\ldots z_{p})=\pm 1\}$, we need to compute integrals of the type
$\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int_{\mathbb{R}^{p}_{(+)}}{\rm d}z_{1}\ldots{\rm d}z_{p}z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}}\ldots e^{-\frac{z_{p}^{2}}{2}}$
$=$ $\frac{2}{2^{p-2}(2\pi)^{\nicefrac{{p}}{{2}}}}\left(\sum_{j=0}^{\lfloor\frac{p-2}{2}\rfloor}\left(\begin{array}[]{cc}p-2\\ 2j\end{array}\right)\int_{\mathbb{R}^{2}_{(+)}}{\rm d}z_{1}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}+\sum_{j=1}^{\lfloor\frac{p-1}{2}\rfloor}\left(\begin{array}[]{cc}p-2\\ 2j-1\end{array}\right)\int_{\mathbb{R}^{2}_{(-)}}{\rm d}z_{1}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}\right)$
$=$ $\frac{2^{3-p}}{(2\pi)^{(p+2)/2}}\sum_{j=0}^{p-2}\left(\begin{array}[]{cc}p-2\\ j\end{array}\right)(-1)^{j}=0$
In the same way it is possible to prove that the integral over $\mathbb{R}_{(-)}^{p}$ is also vanishing. This implies that $\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}=0$, so using (16), we obtain that $\alpha_{c}=+\infty$. This model cannot be learned with $n=\mathcal{O}(d)$ samples for $p\geq 3$.
C.4 $g(z_{1},\ldots,z_{p})=z_{1}^{2}+\operatorname{sign}\left(\prod_{j=1}^{p}z_{j}\right)$
We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
$d_{ij}({\bm{V}})$ $:=\det\left(\begin{array}[]{cc}v_{i,j}&v_{i,p}\\ v_{j,p}&v_{p,p}\end{array}\right),\quad i,j\in{1,\ldots,p},$ (65)
the function
$E_{s,s_{1},s_{p}}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}})=1-s_{p}\operatorname{erf}\left(\frac{v_{1,p}(s_{1}\sqrt{y-s}-\omega_{1})+\sum_{j=2}^{p-1}v_{j,p}(z_{j}-\omega_{j})-v_{pp}\omega_{p}}{\sqrt{2v_{pp}}}\right),$
the Gaussian probability density
$\rho({\bm{z}}\in\mathbb{R}^{p},{\bm{V}}):=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}\sqrt{\det({\bm{V}})}}e^{-\frac{1}{2}{\bm{z}}^{\top}{\bm{V}}^{-1}{\bm{z}}},$ (66)
and finally
$I_{s,s_{1},s_{p-1}}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}}):=\frac{E_{s,s_{1},s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})}{4\sqrt{y-s}}\rho\left(\left(\begin{array}[]{cccc}s_{1}\sqrt{y-s}-\omega_{1}\\ z_{2}-\omega_{2}\ldots\\ z_{p-1}-\omega_{p-1}\end{array}\right),v_{pp}\bm{d}({\bm{V}})^{-1}\right).$
Then, defining $\mathbb{R}_{+1}=[0,+\infty)$ and $\mathbb{R}_{-1}=(-\infty,0]$, we have that
$Z_{\rm out}(y,{\bm{\omega}}{\bm{V}})=\sum_{s,s_{1},\ldots,s_{p}=\pm 1}\delta_{s,(\prod_{j}s_{j})}\bm{1}_{y>s}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s,s_{1},s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}}),$ (67)
where $\bm{1}_{x>0}$ is the Heaviside step function. The components of ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})$ are given by
$({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}}))_{1}$ $=\sum_{s,s_{1},\ldots,s_{p}=\pm 1}\frac{\delta_{s,(\prod_{j}s_{j})}\bm{1}_{y>s}}{Z_{\rm out}}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}\frac{\partial}{\partial\omega_{1}}I_{s,s_{1},s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})$
$({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}}))_{k\neq 1}$ $=\sum_{s,s_{1},\ldots,s_{p}=\pm 1}\frac{s_{k}\delta_{s,(\prod_{j}s_{j})}\bm{1}_{y>s}}{Z_{\rm out}}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s,s_{1},s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\delta(z_{k}-0)$
Stability of the fixed point
The result is dependent on a number of integrals which we can compute analytically. First we have
$Z_{\rm out}(y,\bm{0},{\bm{I}})$ $=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (68)
$=\frac{1}{2(2\pi)^{\nicefrac{{p}}{{2}}}}\sum_{s_{1},\ldots,s_{p}=\pm 1}\frac{\bm{1}_{y>s}}{\sqrt{y-\prod_{j}s_{j}}}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}e^{-\frac{y-s_{1}\ldots s_{p}}{2}-\frac{z_{2}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (69)
$=\frac{1}{2\sqrt{2\pi}}\sum_{s=\pm 1}\bm{1}_{y>s}\frac{e^{-\frac{y-s}{2}}}{\sqrt{y-s}}$ (70)
$=\frac{e^{-\nicefrac{{y}}{{2}}}}{2\sqrt{2\pi e}}\left(\frac{1}{\sqrt{y+1}}+\mathbf{1}_{y>1}\frac{e}{\sqrt{y-1}}\right)$ (71)
From this it is straightforward to see that ${\bm{g}}_{\rm out}^{0}(y)=0$. In order to compute $\nabla_{\omega}{\bm{g}}_{\rm out}(y,0,1)$ we need to consider the additional integrals
1 1)\;1 ) $\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}(z_{1}^{2}-1)e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))$ (72)
$=\frac{1}{2\sqrt{2\pi}}\sum_{s=\pm 1}\bm{1}_{y>s}\frac{e^{-\frac{y-s}{2}}(y-s-1)}{\sqrt{y-s}}$ (73)
$=\frac{e^{-y}}{2\sqrt{2\pi e}}\left(\frac{y}{\sqrt{y+1}}+\mathbf{1}_{y>1}\frac{e(y-2)}{\sqrt{y-1}}\right)$ (74)
2 2)\;2 ) $\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}(z_{k\neq 1}^{2}-1)e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))=0$ (75)
3 3)\;3 ) $\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}z_{j}z_{k}e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))$ (76)
$\propto\sum_{s,s_{j},s_{k}=\pm 1}\bm{1}_{y>s}s_{j}s_{k}e^{-\frac{y-s}{2}}=0$ (77)
This shows that
$\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y)=\begin{bmatrix}C(y)&0&0\\ 0&0&0\\ 0&0&0\\ \end{bmatrix}$ (79)
Where $C(y)$ is
$C(y)=\begin{cases}y&-1<y<1\\ y-2e\frac{\sqrt{y+1}}{e\sqrt{y+1}+\sqrt{y-1}}-1&y>1\\ \end{cases}$ (80)
The critical alpha is simply found by one last integral
$\alpha_{c}=\left[\int\,{\rm d}yC(y)^{2}Z_{\rm out}(y)\right]^{-1}\approx 0.575166$ (81)
C.5 $g(z_{1},\ldots,z_{p})=\sum_{j=1}^{p}\operatorname{sign}(z_{p})$
We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
$d_{ij}({\bm{V}})$ $:=\det\left(\begin{array}[]{cc}v_{i,j}&v_{i,p}\\ v_{j,p}&v_{p,p}\end{array}\right),\quad i,j\in{1,\ldots,p},$ (84)
the function
$E_{s}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}})=1-s\operatorname{erf}\left(\frac{\sum_{j=1}^{p-1}v_{j,p}(z_{j}-\omega_{j})-v_{pp}\omega_{p}}{\sqrt{2v_{pp}}}\right),$ (85)
the Gaussian probability density
$\rho({\bm{z}}\in\mathbb{R}^{p},{\bm{V}}):=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}\sqrt{\det({\bm{V}})}}e^{-\frac{1}{2}{\bm{z}}^{\top}{\bm{V}}^{-1}{\bm{z}}},$ (86)
and finally
$I_{s}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}}):=E_{s}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\rho\left(\left(\begin{array}[]{ccc}z_{1}-\omega_{2}\\ \ldots\\ z_{p-1}-\omega_{p-1}\end{array}\right)^{\top},v_{pp}\bm{d}({\bm{V}})^{-1}\right).$
Then, defining $\mathbb{R}_{+1}=[0,+\infty)$, $\mathbb{R}_{-1}=(-\infty,0]$ and $S_{a}=\{\bm{s}=(s_{1},...,s_{p})|\exists\bm{j}=\{j_{1},...,j_{a}\}\subseteq\{1,...,n\}\text{ s.t }s_{j}=1\text{ for }j\in\bm{j}\text{ and }s_{j}=-1\text{ otherwise}\}$, we have that
$Z_{\rm out}=\sum_{a=1}^{p}\frac{\delta_{y,2a-p}}{2}\sum_{\bm{s}\in S_{a}}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}}),$ (87)
and the $k^{\text{th}}$ component of ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})$ is
$\left({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\right)_{k}=\frac{1}{Z_{\rm out}}\sum_{a=1}^{p}\frac{\delta_{y,2a-p}}{2}\sum_{\bm{s}\in S_{a}}s_{k}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\delta(z_{k}-0)$
Appendix D Further numerical observations and details
Here we give more details about the numerical implementation of State Evolution and AMP. Both approaches require to compute ${\bm{g}}_{\rm out}$. All the examples we implemented are detailed in Appendix C.
The integrals are doing with the quadrature package in Scipy. In order to avoid instabilities we regularize the interval of integration by replacing $\infty$ with $\Lambda$. We typically choose $\Lambda\approx 10$. Similarly, we add $\epsilon\approx 10^{-4}$ to the diagonal of $V$. In State Evolution we need to integrate over functions of ${\bm{g}}_{\rm out}$. We do these integrals using a simple Monte Carlo approach. For Figures 1 and 2 we used a total $72000$ samples, for Figure 4 $7200$. Computing such integrals is the numerical bottleneck. In order to make this part faster we parallelised the MCMC: for each iteration of State Evolution we make every worker in our pool estimate the integral, and then average the estimation and then average among workers. In the cases in which ${{\bm{M}}}=0$ is a fixed point, we initialize ${{\bm{M}}}$ with the empirical overlap of AMP at the beginning of the iteration.
In both the AMP and State Evolution implementation we used some damping: the overlap ${{\bm{M}}}$ or the $\hat{{{\bm{{W}}}}}$ at the new iterations are averaged with the current value, with a weight $\delta$ for the new one, where typically $0.90.6<\delta<$ 0.9. We display the evolution of the overlaps in a typical run of AMP for the model $z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ in Figure 3. We already displayed the values of the overlap and generalisation error at convergence in Figure 2. We can see how AMP has a saddle-to-saddle dynamic, where the algorithm alternates plateaus for $\mathcal{O}(\log{d})$ ) iterations to fast drops in generalisation error, which are associated with new directions being learned.
As stated in the main text, models non-trivial subspaces are associated with symmetries. This also implies that the associated overlaps are have the same invariances. In order to make the plots readable we remove all such symmetries by hand. For this reason we take the absolute value of the overlap if the model is even, and we impose a specific inequality in the overlap if there is invariance under permutation. As a general idea we want to always have the ”best” possible overlap. Meaning we want ${{\bm{M}}}$ to be as diagonal as possible. We list what this mean for the examples in the figures:
  • ${\rm sign}(z_{1}z_{2})$: Because of invariance under exchange of $z_{1}$ and $z_{2}$ we always have ${{\bm{M}}}_{11}={{\bm{M}}}_{22}$. Here AMP can reach $2$ equivalent configurations: either the diagonal or the anti-diagonal is zero. We choose the configuration where the anti-diagonal is zero.
  • $z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$: This model will first learn just $z_{1}$. We fix ${{\bm{M}}}_{11}>0$. For the rest of the components we are reduced to the case above.
  • ${\rm sign}(z_{1})+{\rm sign}(z_{2})+{\rm sign}(z_{3})$: Because of the invariance under permutation each row of ${{\bm{M}}}$ will have either the same element in all the entries (in which case we don’t need to do anything) or two are the same and one is bigger. We permute the matrix such that the largest entry is on the diagonal.
The code to run AMP and State Evolution on the examples in the is available on GitHub

### v2

@@
-Fundamental limits of weak learnability in high-dimensional multi-index models
+Fundamental computational limits of weak learnability in high-dimensional multi-index models
 Abstract
 Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural networks.
-This paper examines the theoretical boundaries of learnability in this hypothesis class, focusing particularly on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples is $n=\alpha d$ is proportional to the covariate dimension $d$.
+This paper examines the theoretical boundaries of efficient learnability in this hypothesis class, focusing particularly on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples is $n=\alpha d$ is proportional to the covariate dimension $d$.
 Our findings unfold in three parts: (i) first, we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!>\!0$; (ii) second, in the case where the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace consisting of directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$.
-The critical threshold $\alpha_{c}$ marks the presence of a computational phase transition, in the sense that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$.
+The critical threshold $\alpha_{c}$ marks the presence of a computational phase transition, in the sense that it is conjectured that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$.
 In a limited but interesting set of really hard directions —akin to the parity problem— $\alpha_{c}$ is found to diverge.
@@
 Multi-index functions are a class of statistical models encoding the inductive bias that the relevant directions for prediction depend only on a low-dimensional subspace of the covariates ${\bf x}\in\mathbb{R}^{d}$:
-$y=g({{\bm{{W}}}}{\bm{x}}),\qquad{{\bm{{W}}}}\in\mathbb{R}^{p\times d},\quad k\in[p].$ (1)
+$y=g({{\bm{{W}}}}{\bm{x}}),\qquad{{\bm{{W}}}}\in\mathbb{R}^{p\times d},$ (1)
+where the mapping $g:\mathbb{R}^{p}\rightarrow\mathbb{R}$ can further be stochastic i.e.
+depend on independent noise $\xi\in\mathbb{R}^{p^{\prime}}$ for some finte $p^{\prime}$.
 They define a rich class of hypotheses (Li, 1991), containing many widely studied functions in the statistical learning and theoretical computer science literature.
-The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z)\!=\!z\!+\!\sqrt{\Delta}Z$, with $Z\!\sim\!{\cal N}(0,1)$.
+The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z,\xi)\!=\!z\!+\!\sqrt{\Delta}\xi$, with $\xi\!\sim\!{\cal N}(0,1)$.
 In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, the terminology single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is employed when the link function is known and only the weights are learned.
@@
 Several authors have thus used multi-index models as a test-bed for understanding the behavior of neural nets and gradient-descent in non-convex, high-dimensional contexts, e.g.
-(Saad and Solla, 1995; 1996; Arous et al., 2021; Abbe et al., 2022; 2023; Veiga et al., 2022; Ba et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2023; Damian et al., 2023; Bietti et al., 2023; Moniri et al., 2023; Berthier et al., 2024).
+(Saad and Solla, 1995; 1996; Arous et al., 2021; Abbe et al., 2022; 2023; Veiga et al., 2022; Ba et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2023; Damian et al., 2023; Bietti et al., 2023; Moniri et al., 2023; Berthier et al., 2023).
 To serve as a useful benchmark, it is necessary to have an idea of the fundamental limit of learnability in such models.
 This translates to the question of how many observations from model Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability.
-This can be studied both statistically (within the class of all, including exponential, algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms).
+This can be studied both statistically (within the class of all, including exponential time algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms).
 In the single-index case ($p=1$), weak learnability has been heavily studied under probabilistic assumptions for the weights and data distribution (e.g.
@@
 (2019) when the covariate dimension $d$ is large.
-Optimal computational thresholds for the class of first-order iterative algorithms were derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020a) in the same regime.
+Optimal computational thresholds for the class of first-order iterative algorithms were derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020) in the same regime.
 Similar results were also proven for other computational models: Damian et al.
@@
 Damian et al.
-(2024b) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
+(2024) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
 Aside from the particular case of committee machines (Aubin et al., 2019; Diakonikolas et al., 2020; Goel et al., 2020; Chen et al., 2022), results for general multi-index models $p>1$ are scarce, as they crucially depend on the way different directions are coupled by the link function.
@@
   • We analyse the computational limits of weak learnability (as defined in Def.
-2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\to\infty$ at fixed ratio $\alpha=\nicefrac{{n}}{{d}}$.
+2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\to\infty$ at fixed ratio $\alpha=\nicefrac{{n}}{{d}}$ and constant number of indices $p=\Theta(1)$.
 Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural networks trained with gradient descent.
   • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples in multi-index models: trivial directions can be learned with a single AMP iteration for any $\alpha>0$.
-When no such direction exists, we show there might be easy directions requiring $O(\log d)$ iterations and at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition.
+When no such direction exists, we show there might be easy directions allowing recovery from arbitrarily small initial overlap $\epsilon>0$ in $\mathcal{O}(\log\frac{1}{\epsilon})$ steps with at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition.
+Heuristically, we expect recovery along easy directions from random initialization with $\alpha>\alpha_{c}$ in $\mathcal{O}(\log d)$ iterations.
+We further confirm this numerically through extensive simulations.
 [^2] Note that $\alpha_{c}$ may or may not coincide with the statistically optimal threshold.
@@
 We provide a sharp formula for the critical sample complexity $\alpha_{c}$ above which easy directions are weakly learnable.
-No efficient iterative algorithm can succeed for $\alpha<\alpha_{c}$.
-The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!O(d)$ samples.
+It is conjectured that no efficient iterative algorithm can succeed for $\alpha<\alpha_{c}$.
+The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!\mathcal{O}(d)$ samples.
   • We discuss how interactions among directions can lead to hierarchical learning phenomena.
@@
 However, the class of functions that are weakly learnable hierarchically with AMP differs from SGD and is, in particular, considerably larger.
-We fully characterise this class of Hierarchical Iterative Denoising Staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
+We characterise this class of Hierarchical Iterative Denoising Staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
 Further related work —
@@
 (2023) has shown that a large-width two-layer neural network trained under one-pass SGD can learn an ”easy” single-index target ($\ell=1$) in $n=\Theta(d)$ steps.
-This is to be contrasted with full-batch GD, which can achieve weak recovery in $\Theta(1)$ steps with sample complexity $n=\Theta(d)$ even for particular problems with $\ell>1$ (Dandi et al., 2024b) (which is, indeed, the statistical optimal rate (Barbier et al., 2019; Mondelli and Montanari, 2019)).
+This is to be contrasted with full-batch GD, which can achieve weak recovery in $\Theta(1)$ steps with sample complexity $n=\Theta(d)$ even for particular problems with $\ell>1$ (Dandi et al., 2024) (which is, indeed, the statistical optimal rate (Barbier et al., 2019; Mondelli and Montanari, 2019)).
 Similar results have also been proven for abstract computational models.
@@
 More recently, Damian et al.
-(2024b) has proven a similar result for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$ and hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
-This turns out to be equivalent to the optimal computational weak recovery threshold of (Barbier et al., 2019; Mondelli and Montanari, 2019; Luo et al., 2019; Maillard et al., 2020b).
+(2024) has proven a similar result for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$ and hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
+This turns out to be equivalent to the optimal computational weak recovery threshold of (Barbier et al., 2019; Mondelli and Montanari, 2019; Luo et al., 2019; Maillard et al., 2020).
 While the situation is less understood, the multicase has also whitenessed a surge of recent interest.
@@
 Interestingly, it was then shown that the limit discovered in these set of works could be bypassed by slightly different algorithm, for instance by smoothing the landscapes Damian et al.
-(2024a) or reusing batches multi-time Dandi et al.
-(2024b).
+(2023) or reusing batches multi-time Dandi et al.
+(2024).
 The latter paper, in particular, showing that gradient descent could learn efficiently a larger class of multi-index models that previously believed to be possible.
@@
 (2020b; a), and are at the roots of the spectral method that underline the learnability phase transition in this case Mondelli and Montanari (2019); Maillard et al.
-(2020b), much less is known in the multi-index case, with the exception of Aubin et al.
-(2019), who showed that particular instances of multi-index models known as committee machines can be learned with $n=\Theta(d)$ samples.
+(2020), much less is known in the multi-index case, with the exception of Aubin et al.
+(2019), who showed that particular instances of multi-index models known as committee machines can be learned with $n=\mathcal{O}(d)$ samples.
 Different from our approach are (Diakonikolas et al., 2020; Goel et al., 2020; Chen et al., 2022), who focused also on peculiar form of the multi-index models (using combination of Relu) and derived worst case bound (in terms of the hardest possible function).
@@
 Notation
-We denote by $\mathcal{S}^{+}_{p}$, the cone of positive-semi-definite matrices in $\mathbb{R}^{p}$ and by $\succ$ the associated partial ordering.
-For ${{\bm{M}}}\in\mathcal{S}^{+}_{p}$, we denote by $\sqrt{M}$ the matrix-square root.
-|||||$\cdot$||| | ⋅ | | denotes the operator norm and $||\cdot||_{F}$ the Frobenius norm.
+[…]
 2 Settings and definitions
@@
 Definition 1 (Gaussian multi-index models).
-Given a covariate ${\bm{x}}\sim\mathcal{N}(0,\nicefrac{{1}}{{d}}{\bm{I}}_{d})$, we define the class of Gaussian multi-index models as likelihoods of the type:
-$\mathbb{E}[y|{\bm{x}}]=g({{\bm{{W}}}}^{\star}{\bm{x}})$ (2)
-where $g:\mathbb{R}^{p}\to\mathbb{R}$ is the link function and ${{\bm{{W}}}}^{\star}\in\mathbb{R}^{p\times d}$ is a weight matrix with i.i.d.
+Given a covariate ${\bm{x}}\sim\mathcal{N}(0,\nicefrac{{1}}{{d}}{\bm{I}}_{d})$, we define the class of Gaussian multi-index models as output mappings of the type:
+$y=g({{\bm{{W}}}}^{\star}{\bm{x}})$ (2)
+where $g:\mathbb{R}^{p}\to\mathbb{R}$ denotes the link function and ${{\bm{{W}}}}^{\star}\in\mathbb{R}^{p\times d}$ is a weight matrix with i.i.d.
 rows ${\bm{w}}^{\star}_{k}\sim\mathcal{N}(0,{\bm{I}}_{d})$.
-Note that this uniquely defines a joint distribution $p({\bm{x}},y)$ over $\mathbb{R}^{d+1}$.
+Note we allow the map $g:\mathbb{R}^{p}\to\mathbb{R}$ to be stochastic i.e $g(\cdot)=\tilde{g}(\cdot,{\bm{\xi}})$ for random variable ${\bm{\xi}}\in\mathbb{R}^{p}$ independent of $\mathbf{x}$ having a fixed dimension-independent law such that $\mathbb{E}\left[\norm{\xi}_{2}^{2}\right]<\infty$.
+[^3] Technically, our results hold for ${\bm{\xi}}\in\mathbb{R}^{p^{\prime}}$ with $p^{\prime}=\mathcal{O}(1)$, for example the case of a degenerated noise.
+But for simplicity of exposition we focus on the $p^{\prime}=p$ case.
 Given $n$ i.i.d.
@@
 Let $V^{\star}\subset\mathbb{R}^{p}$ denote a subspace spanned by vectors representing components along ${{\bm{{W}}}}^{\star}$ such that each $\mathbf{v}\in V^{\star}$ maps to a vector $\mathbf{v}_{d}$ in $\mathbb{R}^{d}$ through the map $\mathbf{v}_{d}=({{\bm{{W}}}}^{\star})^{\top}\mathbf{v}$.
-Given an estimator $\hat{{{\bm{{W}}}}}\in\mathbb{R}^{p\times d}$ of ${{\bm{{W}}}}^{\star}$, we have weak recovery of a $V^{\star}$ if:
-$\underset{{\bm{v}}\in V^{\star},\norm{{\bm{v}}}=1}{\inf}\norm{\frac{\hat{{{\bm{{W}}}}}({{\bm{{W}}}}^{\star})^{\top}{\bm{v}}}{d}}=\Theta_{d}(1).$ (3)
+Given an estimator $\hat{{{\bm{{W}}}}}\in\mathbb{R}^{p\times d}$ of ${{\bm{{W}}}}^{\star}$ with $\norm{\hat{{{\bm{{W}}}}}_{i}}^{2}_{F}=\Theta(d)$, we have weak recovery of a $V^{\star}$ if:
+$\underset{{\bm{v}}\in\mathbb{R}^{p},\norm{{\bm{v}}}=1}{\inf}\norm{\frac{\hat{{{\bm{{W}}}}}({{\bm{{W}}}}^{\star})^{\top}{\bm{v}}}{d}}=\Theta(1).$ (3)
 with high probability as $d\to\infty$.
 Our main tool for characterising the computational bottlenecks in the Gaussian multi-index problem is an Approximate Message Passing (AMP) Algorithm 1 tailored to our problem.
-Note that Algorithm 1 only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and the application of a problem dependent function ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$.
-Therefore, it belongs to the class of first-order algorithms with linear running time on $O(nd+p^{3})$.
+Note that Algorithm 1 only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and the application of a problem dependent function ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$ (defined later in Equation 4).
+Therefore, it belongs to the class of first-order algorithms with linear running time of $\mathcal{O}(nd)$.
+Intuitively, AMP estimates the mean ${\bm{\omega}}$ and variance ${\bm{V}}$ of ${{\bm{{W}}}}^{*}{\bm{x}}$ and the mean $\hat{{{\bm{{W}}}}}$ and covariance $\hat{{\bm{C}}}$ of ${{\bm{{W}}}}^{*}$ through a fixed point iteration described in Algorithm 1.
 The key property of AMP is that for a well-chosen ${\bm{g}}_{\rm out}$, it is provably optimal within the class of first-order methods (Celentano et al., 2020).
@@
 where the conditional expectation is defined through the output channel $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$.
+For stochastic mappings $g(\cdot)=g(\cdot,\xi)$, the expectation is w.r.t both $Z,\xi$ conditioned on $Y$.
 Here ${\bm{\omega}}$ can be interpreted as an estimate of the pre-activations.
+Our results rely on the following assumption on the above denoiser, translating to an indirect assumption on $g,\xi$:
+Assumption 1.
+${\bm{g}}_{\rm out}:\mathbb{R}^{p^{2}+p+1}\rightarrow\mathbb{R}^{p}\in\mathcal{C}^{2}$.
+Furthermore, the mapping $\hat{g}_{\rm out}:\mathbb{R}^{p+{p^{\prime}}}\rightarrow\mathbb{R}^{p}$ defined by ${\hat{g}}_{\rm out}(\mathbf{h},\xi)={\bm{g}}_{\rm out}(\tilde{g}((\mathbf{h},\xi),{\bm{\omega}},{\bm{V}})$ is pseudo-Lipschitz of finite-order for all ${\bm{\omega}},{\bm{V}}\in\mathbb{R}^{p},\mathbb{R}^{p\times p}$.
+Here, $\xi\in\mathbb{R}^{p^{\prime}}$ denotes independent noise satisfying $\mathbb{E}\left[\norm{\xi}^{2}\right]<\infty$ as per Definition 1.
 The above denoiser can also be derived as an exact high-dimensional approximation of the Belief Propagation (BP) algorithm for the estimation of the marginals of the Bayesian posterior distribution:
@@
 Our analysis is based on two remarkable properties that make AMP a particularly useful tool for studying high-dimensional estimation.
-The first property is that for any $t<O(\log{d})$ ), in the high-dimensional limit $d\to\infty$ the performance of AMP can be tracked without actually running the algorithm.
+The first property is that for any $t=\Theta(1)$, in the high-dimensional limit $d\to\infty$ the performance of AMP can be tracked without actually running the algorithm.
 This result, known as state evolution, makes AMP mathematically tractable in high-dimensions (Bayati and Montanari, 2011).
@@
 We discuss this further in Appendix A.
-For multi-index models, the state evolution equations were derived by Aubin et al. (2020b) and rigorously proven by Gerbelot and Berthier (2023):
+For multi-index models, the state evolution equations were derived by Aubin et al.
+(2020b) and rigorously proven by Gerbelot and Berthier (2023).
+The state evolution equations, in particular, provide an exact characterization of the asymptotic overlaps and prediction error.
 Lemma 2.1 (State evolution (Aubin et al., 2019; Gerbelot and Berthier, 2023)).
@@
 Denote by $\hat{{{\bm{{W}}}}}^{t}$ the resulting estimator at time $0\leq t\leq T$.
-Then, in the high-dimensional limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$ and $T<O(\log{d})$ ), the asymptotic mean-squared error on the label prediction is given by:
-$\lim\limits_{n,d\to\infty}\mathbb{E}[(y-g(\hat{{{\bm{{W}}}}}^{t}{\bm{x}}))^{2}]=\mathbb{E}[(Y^{t}-g({\bm{Z}}))^{2}],$ $0\leq t\leq T$ (6)
-where the expectation is taken over the following effective estimation process:
-$Y^{t}=g\left(({\bm{I}}_{p}-{{\bm{M}}}^{t})^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{{{\bm{M}}}^{t}}^{\nicefrac{{1}}{{2}}}{\bm{\xi}}\right),\qquad{\bm{Z}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ (7)
-with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ independently from ${\bm{Z}}$, and ${{\bm{M}}}^{t}$ is given by iterating the following state evolution equations from initial condition ${{\bm{M}}}^{0}$:
+Then, in the high-dimensional limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, constant $p$ and any finite time $T$, the limiting overlaps satisfy:
+[…]
+with ${{\bm{M}}}^{t}$ satisfying the following state evolution equations from initial condition ${{\bm{M}}}^{0}$:
 ${{\bm{M}}}^{t+1}\!$ $=\!F({{\bm{M}}}^{t})\!\coloneqq{G}\!\left(\!\alpha\,\mathbb{E}\Big{[}{\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}\!\Big{]}\right).$ (7)
-where $G({\bm{X}})=({\bm{I}}_{p}+{\bm{X}})^{-1}{\bm{X}}$ and the expectation is taken over the effective process Equation 7.
+where $G({\bm{X}})=({\bm{I}}_{p}+{\bm{X}})^{-1}{\bm{X}}$ and the expectation is taken over the following effective process
+$Y^{t}=g\left(({\bm{I}}_{p}-{{\bm{M}}}^{t})^{\nicefrac{{1}}{{2}}}{\bm{\xi}}+{{{\bm{M}}}^{t}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}\right),\qquad{\bm{Z}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ (8)
+with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ independently from ${\bm{Z}}$.
 Furthermore, ${{\bm{M}}}^{t}$ is constrained to lie in $\mathcal{S}^{+}_{p}$ due to Nishimori-identity (section G.3 in (Aubin et al., 2019)).
-The limiting overlap $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ converge in probability to ${{\bm{M}}}^{t}$.
+As a consequence, the asymptotic mean-squared error on the label prediction is given by:
+${\mathbb{E}_{\bm{X},\bm{y}}}\mathbb{E}_{\bm{x},y}\left[\left(y-g\left(\hat{\bm{W}}^{t}(\bm{X},\bm{y})\bm{x}\right)\right)^{2}\right]=\mathbb{E}[(Y^{t}-g({\bm{Z}}))^{2}],$ $0\leq t\leq T$ (9)
+where the expectation is taken over the effective estimation process Equation 8.
 3 The trivial subspace ($\alpha_{c}=0$)
@@
 In the absence of any prior information on ${{\bm{{W}}}}^{\star}$ aside from its distribution, one cannot do better than taking $\hat{{\bm{w}}}_{k}^{t=0}\sim\mathcal{N}(0,{\bm{I}}_{d})$ with $k\in[p]$ independently at random from the prior.
-With high-probability, at initialisation, the elements of the overlap matrix $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ are $\Theta_{d}(d^{-\nicefrac{{1}}{{2}}})$ numbers.
+With high-probability, at initialisation, the elements of the overlap matrix $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ are $\Theta(d^{-\nicefrac{{1}}{{2}}})$ element-wise.
 The asymptotic overlap for an uninformed initial condition is thus identically zero ${{\bm{M}}}^{0}=0$, a null-rank matrix.
@@
 ${{\bm{M}}}=0\in\mathbb{R}^{p\times p}$ is a fixed point of Equation 7 if and only if the following condition holds almost surely over $Y$:
-${\bm{g}}_{\rm out}(y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})=\mathbb{E}[{\bm{Z}}|Y=y]=\mathbf{0},$ (9)
+${\bm{g}}_{\rm out}(Y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})=\mathbb{E}[{\bm{Z}}|Y]=\mathbf{0},$ (10)
 This implies that as long as the conditional expectation above is not zero almost surely, AMP weakly learns a non-empty subspace immediately in the first iteration for any number of samples $n=\Theta(d)$.
@@
 Definition 3 (Trivial subspace).
-Let $H^{\star}_{T}\subseteq\mathbb{R}^{p}$ denote the subspace spanned by $\mathbf{v}\in\mathbb{R}^{p}$ such that:
-${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\mathbb{E}[\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle|Y=y]=0,$ (10)
-almost surely over $Y$ Equation 7. The trivial subspace $T^{\star}$ is defined as the orthogonal complement of $H^{\star}$.
-Note that for single-index models ($p=1$), the condition in Equation 9 reduces to the one derived by (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020b).
+We define $H^{\star}_{T}\subseteq\mathbb{R}^{p}$ as the subspace spanned by $\mathbf{v}\in\mathbb{R}^{p}$ such that the following holds almost surely over $Y=g({\bm{Z}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
+${\bm{g}}_{\rm out}(Y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}[\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle|y=Y]=0,$ (11)
+where the expectation is w.r.t the joint measure $p(\mathbf{x},y)$ defined in Definition 1. Equation 8. The trivial subspace $T^{\star}$ is defined as the orthogonal complement of $H^{\star}$.
+Note that for single-index models ($p=1$), the condition in Equation 10 reduces to the one derived by (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020).
 Interestingly, this is exactly the same condition appearing in Damian et al.
-(2024b) for weak learnability of single-index models in the SQ model; see eq.
+(2024) for weak learnability of single-index models in the SQ model; see eq.
 (3) therein.
@@
 From this perspective, the statement on the condition for the existence of a non-empty trivial subspace Equation 11 translates to the following condition:
-$\mathbb{E}\left[{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=\mathbb{E}\left[\mathbb{E}[(\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|Y=y])^{2}\right]\neq 0,$ (12)
+$\mathbb{E}\left[{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=\mathbb{E}\left[\mathbb{E}[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|Y=y]^{2}\right]\neq 0,$ (13)
 where ${\bm{v}}\in T^{\star}$.
-Indeed, this can be interpreted as a statistical query of the type $\mathbb{E}[\varphi(y)\psi(x)]$ with label pre-processing $\varphi={\bm{g}}_{\rm out}$.
+Indeed, the left-hand side can be interpreted as a statistical query of the type $\mathbb{E}[\varphi(y)\psi(x)]$ with label pre-processing $\varphi={\bm{g}}_{\rm out}$.
 The fact that this linear correlation in the transformed labels is non-vanishing implies that one can weakly recover $\mathbf{v}$ through a tailored spectral method (Mondelli and Montanari, 2019).
@@
 Lemma 3.2.
-For any $\mathbf{u}\in H^{\star}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
-$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{u}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=0$ (13)
-Note that this is a non-asymptotic statement about the denoising function ${\bm{g}}_{\rm out}$, and the expectation is with respect to the distribution of the labels $y$ (not the effective problem).
+For any $\mathbf{v}\in H^{\star}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
+$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=0$ (14)
+Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem).
 Proof.
 The above is a direct consequence of the Tower law of expectation. Specifically, we have:
-$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right].$
-The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}W^{\star}\mathbf{v}=\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$.
+$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right].$
+The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$.
 ∎
-This implies that when $T^{\star}$ is empty, running one-pass SGD after applying any transformation $\mathcal{T}$ fails to obtain weak recovery with $O(d)$ samples.
+This implies that when $T^{\star}$ is empty, running one-pass SGD after applying any transformation $\mathcal{T}$ fails to obtain weak recovery with $\mathcal{O}(d)$ samples, as the absence of a first-order Hermite coefficient implies a sample complexity of $\mathcal{O}(d\log d)$ as shown in Theorems 1.3 and 1.4 in Arous et al.
+(2021).
 The situation can be very different when reusing batches: Dandi et al.
-(2024b) showed that full batch GD implicitly applies a transformation $\mathcal{T}$ to the labels, in a mechanism similar to the one discussed above.
+(2024) showed that full batch GD implicitly applies a transformation $\mathcal{T}$ to the labels, in a mechanism similar to the one discussed above.
 However, this transformation may not be optimal.
-For instance, while AMP learns the trivial subspaces of both $g(z)={\rm He}_{3}(z)$ and staircase functions (11) in a single-step, GD requires 2-steps for the first and $p$ steps for the latter (Dandi et al., 2024b).
+For instance, while AMP learns the trivial subspaces of both $g(z)={\rm He}_{3}(z)$ and staircase functions (12) in a single-step, GD requires 2-steps for the first and $p$ steps for the latter (Dandi et al., 2024).
 4 Computational phase transitions ($\alpha_{c}>0$)
@@
 with $\|\cdot\|_{F}$ the Frobenius norm, and the extremum being achieved at a unique extremizer ${{\bm{M}}}^{\star}\in\mathcal{S}^{+}_{p}$.
-Moreover, if $\norm{\mathcal{F}({{\bm{M}}})}_{F}=0$, then ${{\bm{M}}}=\mathbf{0}$ is stable a fixed point for any $n=\Theta(d)$.
+Moreover, if $\norm{\mathcal{F}({{\bm{M}}})}_{F}=0$, then ${{\bm{M}}}=\mathbf{0}$ is a stable fixed point for any $n=\Theta(d)$.
 The above result follows from a generalization of the Perron-Frobenius theorem to the cone $\mathcal{S}^{+}_{p}$; see Section B.2 for a proof.
@@
 This marks the onset of a computational phase transition for weak learnability in the first-order methods class.
-Note that the computational weak learnability threshold $\alpha_{c}$ can be equal to the information-theoretical one (as, for instance, in ”real” phase retrieval (Mondelli and Montanari, 2019; Maillard et al., 2020a)), and by no means we imply the existence (or lack of thereof) of a computational-to-statistical gap (Zdeborová and Krzakala, 2016).
+Note that the computational weak learnability threshold $\alpha_{c}$ can be equal to the information-theoretical one (as, for instance, in ”real” phase retrieval (Mondelli and Montanari, 2019; Maillard et al., 2020)), and by no means we imply the existence (or lack of thereof) of a computational-to-statistical gap (Zdeborová and Krzakala, 2016).
 The easy subspace spanned by the directions that can be learned with $n>\alpha_{c}d$ samples can be characterised as follows:
@@
 Theorem 4.2.
-Suppose that $T^{\star}={0}$ and the AMP algorithm 1 is initialised at $\hat{{{\bm{{W}}}}}^{0}$ such that ${{\bm{M}}}_{d}^{0}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}=\epsilon{\bm{A}}$, with $\norm{{\bm{A}}}_{F}=1$ and ${\bm{A}}$ being full-rank.
-Then, with high probability as $d\rightarrow\infty$, for $\alpha\geq\alpha_{c}$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}^{\star}$ for $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{\star}$ is as defined in Lemma 4.1.
+Let ${{\bm{M}}}_{d}^{t}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ denote the model-target overlap matrix at any finite time $t$.
+Suppose that $T^{\star}={0}$ and the AMP algorithm 1 is initialised at $\hat{{{\bm{{W}}}}}^{0}$ such that ${{\bm{M}}}_{d}^{0}=\epsilon{\bm{A}}$, with $\norm{{\bm{A}}}_{F}=1$ and ${\bm{A}}$ being full-rank.
+Then,under Assumption 1, with high probability as $d\rightarrow\infty$, for $\alpha\geq\alpha_{c}$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}^{\star}$ for $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{\star}$ is as defined in Lemma 4.1.
 For $\alpha<\alpha_{c}$ however, ${{\bm{M}}}^{t}_{d}=0$ is asympotically stable i.e.
-there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}}\rightarrow 0$ as $t\rightarrow\infty$.
-Furthermore, there exists an $\alpha>0$ and a $\delta>0$ such that is ${{\bm{M}}}_{d}^{t}\succ\delta M_{E^{*}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$ iterations, where $M_{E^{\star}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}$.
-Heuristically, random initialization is equivalent to setting $\epsilon=O(\nicefrac{{1}}{{\sqrt{d}}})$ in the above theorem.
-Therefore, we expect that weak-recovery along $E^{\star}$ can be achieved in $O(\log d)$ iterations.
-As for Lemma 3.1, the expression for the weak recovery threshold Equation 16 generalises the single-index expression from (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020a).
+there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{d}}\rightarrow 0$ as $t\rightarrow\infty$.
+Furthermore, there exists an $\alpha>0$ and a $\delta>0$ such that ${{\bm{M}}}_{d}^{t}\succ\delta M_{E^{*}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$ iterations, where $M_{E^{\star}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}$.
+Note that in the above the initial overlap scale $\epsilon>0$ can be arbitrarily small but constant with respect to $d$.
+Heuristically, $\epsilon$ is often identified with $\mathcal{O}(\nicefrac{{1}}{{\sqrt{d}}})$, the scale of a random initial condition.
+Although this heuristic is not covered by Theorem 4.2, we have numerically checked it is consistent with the extrapolation of the theory, which suggests that for $\alpha>\alpha_{c}$ easy subspaces can be weakly recovered with $\mathcal{O}(\log d)$ steps, see Figure 2.
+As for Lemma 3.1, the expression for the weak recovery threshold Equation 17 generalises the single-index expression from (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020).
 Indeed, for $p=1$, we have $\partial_{\omega}g_{\rm out}(y,0,1)=\mathbb{E}[{\rm He}_{2}(Z)|Y=y]$, and therefore
@@
 (2019).
-In other words, as long as the link function $g$ has a non-zero second-order coefficient in the Hermite basis, it is learnable by Algorithm 1 with $n>\alpha_{c}d$ samples.
 Optimality of AMP then implies that no first-order method can achieve non-vanishing correlation with ${\bm{w}}^{\star}$ when $n<\alpha_{c}d$ as $d\to\infty$.
@@
 Details on the numerical implementation are discussed in Appendix D.
-The sparsity parity function is a classic example of a computationally hard function in theoretical computer science (Blum et al., 1994; 2003).
+The sparse-parity function is a classic example of a computationally hard function in theoretical computer science (Blum et al., 1994; 2003).
 [^4] Strictly speaking, the noiseless parity is efficiently learnable by Gaussian elimination (Blum et al., 1994).
@@
 Lemma 4.3.
-For any $\mathbf{v}^{\star}\in H^{\star}_{E}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
-$\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\star},\mathbf{x}\rangle)\right]=0$ (20)
-Note that this is a non-asymptotic statement about the denoising function ${\bm{g}}_{\rm out}$, and the expectation is with respect to the distribution of the labels $y$ (not the effective problem from Equation 7).
-For an AMP-hard function (that is, when $\alpha_{c}=\infty$), this implies that, even after applying any transformation to the output (as allowed in SQ) an algorithm like SGD would require at least $O(d^{2})$ data, or $O(d^{3/2})$ if we allow smoothing as in Damian et al.
-(2024a).
+For any $\mathbf{v}\in H^{\star}_{E}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
+$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0$ (21)
+Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem from Equation 8).
+For an AMP-hard function (that is, when $\alpha_{c}=\infty$), the above Lemma implies that, even after applying any transformation to the output (as allowed in SQ) online-SGD would require at least $\Omega(d^{2})$ data, since online-SGD with squared loss requires $\Omega(d^{\kappa-1})$ samples/iterations for weak-recovery of a target function when the first $\kappa$ Hermite-coefficients are 00 (Theorems 1.3 and 1.4 in Arous et al.
+(2021)).
+Furthermore, for single-index models satisfying the condition in Lemma 4.3 (i.e $\alpha_{c}=\infty$), the SQ lower-bound predicts a sample complexity requirement of at-least $\Omega(d^{3/2})$ (Damian et al., 2024).
+Therefore, learning such functions with $\mathcal{O}(d)$ samples is conjectured to be algorithmically-hard.
 5 Hierarchical Iterative Denoising
@@
 almost surely over ${\bm{\xi}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
-We define the “trivially-denoising-coupled” subspace $T^{\star}_{U}$ for $U$ as the orthogonal complement of $H_{T}^{\star}(U)$.
+We define the “trivially-denoising-coupled” subspace $T^{\star}_{U}$ for $U$ as the orthogonal complement of $H_{E}^{\star}(U)$.
 Analogously, let $H^{\star}_{E}(U)$ be the subspace spanned by directions ${\bm{v}}\in U^{\perp}$ such that:
 ${\bm{v}}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U}){\bm{v}}=0,$ (23)
-almost surely over $bxi$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
+almost surely over ${\bm{\xi}}$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
 When ${{\bm{M}}}_{U}$ is additionally a fixed point of $\mathcal{F}_{{{\bm{M}}}}$, we can linearise $\mathcal{F}_{{{\bm{M}}}}$ along the orthogonal complement of $U$.
@@
 Let $U\in\mathbb{R}^{p}$ be a fixed subspace.
-Suppose the AMP algorithm 1 is initialised such that ${{\bm{M}}}_{d}^{0}={{\bm{M}}}_{U}+\epsilon{\bm{A}}$, where $\norm{{\bm{A}}}=1$ and ${{\bm{M}}}_{U}$ is a fixed point of $F({{\bm{M}}})$ in Equation 8 with $\operatorname{span}({{\bm{M}}}_{U})=U$.
-If $\alpha\geq\alpha_{\text{hid}}({{\bm{M}}}_{U})$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}^{(t)}\succ\delta{{\bm{M}}}^{*}_{U}$ for some $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{*}_{U}$ is defined as in Definition 5.
+Suppose the AMP algorithm 1 is initialised such that ${{\bm{M}}}_{d}^{0}={{\bm{M}}}_{U}+\epsilon{\bm{A}}$, where $\norm{{\bm{A}}}_{F}=1$ and ${{\bm{M}}}_{U}$ is a fixed point of $F({{\bm{M}}})$ in Equation 7 with $\operatorname{span}({{\bm{M}}}_{U})=U$.
+Then, under Assumption 1, the following holds with high probability as $d\rightarrow\infty$: If $\alpha\geq\alpha_{\text{hid}}({{\bm{M}}}_{U})$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}^{(t)}\succ\delta{{\bm{M}}}^{*}_{U}$ for some $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{*}_{U}$ is defined as in Definition 5.
 For $\alpha<\alpha_{\text{hid}}({{\bm{M}}}_{U})$ however, there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{\perp}}=0$ as $t\rightarrow\infty$, where ${{\bm{M}}}^{t}_{\perp}$ denotes the projection of ${{\bm{M}}}^{t}$ orthogonal to $U$.
@@
 It is easy to construct a multi-index model where AMP will iterate over any number of such plateaus. For instance, consider the model:
-$g({\bm{z}})=z_{1}^{4}+{\rm sign}(z_{1}z_{2}z_{3})+{\rm sign}(z_{3}z_{4}z_{5})+{\rm sign}(z_{5}z_{6}z_{7})+\ldots$ (25)
+$g({\bm{z}})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})+{\rm sign}(z_{3}z_{4}z_{5})+{\rm sign}(z_{5}z_{6}z_{7})+\ldots$ (26)
 After the first plateau to learn $z_{1}$, there will be one for $z_{2}$, then for $z_{3}$, for $z_{4}$ and so on.
@@
 (2020a), and provide a very early example of iterative sequential learning.
-Finally, it is important to stress that hierarchical learnable functions are not equivalent to staircase functions (Abbe et al., 2022; 2023).
-For instance, (25) is learned with $\Theta(d)$, while it requires $O(d^{3})$ samples in the CSQ staircase characterisation.
+Finally, it is important to stress that hierarchical learnable functions are not equivalent to staircase functions (Abbe et al., 2022; 2023). For instance, consider the target:
+$g({\bm{z}})=\rm He_{4}(z_{1})+{\rm sign}(z_{1}z_{2}z_{3}).$ (27)
+Similar to Example 26, the above target is learned sequentially with $\mathcal{O}(d)$ samples by first recovering the direction corresponding to $z_{1}$.
+However, under the CSQ staircase classification, $g({\bm{z}})$ corresponds to leap $3$.
+The CSQ lower-bound therefore implies a sample complexity-requirement of $\Omega(d^{1.5})$ with online-SGD learning it in $\mathcal{O}(d^{2})$ samples (Abbe et al., 2023).
 As we argued in Lemmas 3.2 and 4.3, we believe Hierarchical Iterative Denoising provides the right notion of computational complexity for these models, since it encompasses a larger class of natural algorithms such as batch gradient descent.
@@
 For the former, numerical experiments support that the predictions are already accurate at fairly moderate sizes.
-For the latter, a number of universality results justify the scope of high-dimensional Gaussian covariates and extend AMP to more structured designs (Loureiro et al., 2021; Wang et al., 2022; Dudeja et al., 2023; Dandi et al., 2024a).
+For the latter, a number of universality results justify the scope of high-dimensional Gaussian covariates and extend AMP to more structured designs (Loureiro et al., 2021; Wang et al., 2022; Dudeja et al., 2023; Dandi et al., 2023).
 Acknowledgements
 The authors would like to thank Joan Bruna, Theodor Misiakiewicz, Luca Pesce, and Nati Srebro for stimulating discussions.
-BL acknowledges funding from the Choose France - CNRS AI Rising Talents program.
-This work was supported by the Swiss National Science Foundation under grants SNSF SMArtNet (grant number 212049) and SNSF OperaGOST (grant number 200390).
+This work was supported by the Swiss National Science Foundation under grants SNSF SMArtNet (grant number 212049) and SNSF OperaGOST (grant number 200390) and the Choose France - CNRS AI Rising Talents program.
 Appendix / supplemental material
@@
 Denote by $\hat{{{\bm{{W}}}}}_{\rm bo}=\mathbb{E}[W|X,y]\in\mathbb{R}^{p\times d}$ the mean of the posterior marginals Equation 5.
-Then, in the high-dimensional asymptotic limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, the asymptotic correlation between the posterior mean and ${{\bm{{W}}}}^{\star}$:
+Then, under Assumption 1 in the high-dimensional asymptotic limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, the asymptotic correlation between the posterior mean and ${{\bm{{W}}}}^{\star}$:
 ${{\bm{M}}}^{\star}=\lim\limits_{d\to\infty}\mathbb{E}\left[\frac{1}{d}\hat{{{\bm{{W}}}}}_{\rm bo}{{{\bm{{W}}}}^{\star}}^{\top}\right]$ (28)
@@
 Therefore, the rank of $M^{\star}$ defines the dimension of the statistically optimal subspace reconstruction at sample complexity $\alpha\coloneqq\nicefrac{{n}}{{d}}$.
-The fact that AMP follows state evolution is proven for such problems in Gerbelot and Berthier [2023]
+The fact that AMP follows state evolution is proven for such problems in [Javanmard and Montanari, 2013, Gerbelot and Berthier, 2023].
 Appendix B Proofs of the main results
 B.1 Linear Approximation
-Throughout, we assume that ${\bm{g}}_{\rm out}:\mathbb{R}^{k^{2}+k+1}\rightarrow\mathbb{R}^{k}$ is in $\mathcal{C}^{2}$.
 Our analysis relies on the following result:
@@
 We proceed through an-entry-wise expansion of each term inside the expection in $F({{\bm{M}}})$ around ${{\bm{M}}}=0$.
+Recall that ${\bm{g}}_{\rm out}\in\mathcal{C}^{2}$ by Assumption 1.
 Since ${{\bm{M}}}_{F}\leq p$ , the first two derivatives of ${\bm{g}}_{\rm out}$ are uniformly bounded in ${\bm{g}}_{\rm out}$ for any fixed $Y,{\bm{\xi}}$.
@@
 since $\mathcal{F}(A)\in\mathcal{S}^{+}_{p}$. Define:
-$\nu_{\mathcal{F}}=\inf{\mathbf{v}\in\mathbb{E},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v.$ (40)
+$\nu_{\mathcal{F}}=\inf_{\mathbf{v}\in\mathbb{E},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v.$ (42)
 Since $\mathbf{v}$ lies in a compact set, $\nu_{\mathcal{F}}>0$. Let $A^{\star}$ denote the projection of $A$ along $E^{\star}$ and consider its eigendecomposition $A^{\star}$:
@@
 $\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t+1})\mathbf{v}_{i}\geq\alpha\nu_{\mathcal{F}}\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t})\mathbf{v}_{i}-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (44)
-Therefore, for $\alpha\geq\frac{1}{\nu_{\mathcal{F}}}$ and small enough $\epsilon$, ${{\bm{M}}}^{t}$ expands linearly along each $\mathbf{v}_{i}$
-B.4 Proof of Theorem 5.1
+Therefore, for $\alpha\geq\frac{1}{\nu_{\mathcal{F}}}$ and small enough $\epsilon$, ${{\bm{M}}}^{t}$ expands linearly along each $\mathbf{v}_{i}$.
+B.4 Proof of Lemma 4.3
+From the definition of $g_{out}$ in Equation 4 and the dominated-convergence theorem, it is straightforward to see that:
+$\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}}=\mathbf{0},{\bm{I}})=\mathbb{E}[{\bm{Z}}{\bm{Z}}^{\top}-{\bm{I}}|Y].$ (45)
+Therefore for any $\mathbf{v}\in\mathbb{R}^{p}$ with $\norm{\mathbf{v}}=1$:
+$\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}},{\bm{I}})_{{\bm{\omega}}=\mathbf{0}}\mathbf{v}$ $=\mathbb{E}[(\mathbf{v}^{\top}{\bm{Z}})^{2}-1|Y]$
+$=\lim_{d\rightarrow\infty}\mathbb{E}[\rm He_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star}\mathbf{x}\rangle)|y=Y],$
+where the last expectation is w.r.t the joint measure $p(\mathbf{x},y)$ defined in Definition 1.
+Now, for any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$, we have, by the tower law of expectation:
+$\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\rm He_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)|y\right]\right].$ (46)
+For any $\mathbf{v}\in H^{\star}_{E}$, the integrand vanishes asymptotically almost surely over $y$. Therefore, the dominated-convergence theorem implies that for $\mathbf{v}\in H^{\star}_{E}$:
+$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0.$ (47)
+B.5 Proof of Theorem 5.1
 By assumption, $M_{U}$ is a fixed point of $F(M)$.

### v3

@@
 Abstract
-Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural networks.
-This paper examines the theoretical boundaries of efficient learnability in this hypothesis class, focusing particularly on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples is $n=\alpha d$ is proportional to the covariate dimension $d$.
-Our findings unfold in three parts: (i) first, we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!>\!0$; (ii) second, in the case where the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace consisting of directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$.
-The critical threshold $\alpha_{c}$ marks the presence of a computational phase transition, in the sense that it is conjectured that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$.
-In a limited but interesting set of really hard directions —akin to the parity problem— $\alpha_{c}$ is found to diverge.
-Finally, (iii) we demonstrate that interactions between different directions can result in an intricate hierarchical learning phenomenon, where some directions can be learned sequentially when coupled to easier ones.
-Our analytical approach is built on the optimality of approximate message-passing algorithms among first-order iterative methods, delineating the fundamental learnability limit across a broad spectrum of algorithms, including neural networks trained with gradient descent.
+Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural nets.
+This paper examines the theoretical boundaries of efficient learnability in this hypothesis class, focusing on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples $n\!=\!\alpha d$ is proportional to the covariate dimension $d$.
+Our findings unfold in three parts: (i) we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!\!>\!\!0$; (ii) if the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace where directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$, where $\alpha_{c}$ marks a computational phase transition.
+In a limited but interesting set of really hard directions –akin to the parity problem– $\alpha_{c}$ is found to diverge.
+Finally, (iii) we show that interactions between different directions can result in an intricate hierarchical learning phenomenon, where directions can be learned sequentially when coupled to easier ones.
+We discuss in detail the grand staircase picture associated to these functions (and contrast it with the original staircase one).
+Our theory builds on the optimality of approximate message-passing among first-order iterative methods, delineating the fundamental learnability limit across a broad spectrum of algorithms, including neural networks trained with gradient descent, which we discuss in this context.
 1 Introduction
-A fundamental property of neural networks is their ability to learn features from data and adapt to relevant structures in high-dimensional noisy data.
+A fundamental property of neural networks is their ability to learn features and adapt to relevant structures in high-dimensional noisy data.
 However, our mathematical understanding of this mechanism remains limited.
@@
 The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z,\xi)\!=\!z\!+\!\sqrt{\Delta}\xi$, with $\xi\!\sim\!{\cal N}(0,1)$.
-In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, the terminology single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is employed when the link function is known and only the weights are learned.
+In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is used when the link function is known and only the weights are learned.
 In this work, we use both interchangeability., $g(z)$ can be an arbitrary, possibility stochastic, link function, including some widely studied problems such as phase retrieval $g(z)\!=\!z^{2}$ and the perceptron $g(z)={\rm sign}(z)$.
@@
 To serve as a useful benchmark, it is necessary to have an idea of the fundamental limit of learnability in such models.
-This translates to the question of how many observations from model Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability.
+This translates to the question of how many observations from the model in Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability.
 This can be studied both statistically (within the class of all, including exponential time algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms).
@@
 (2019) when the covariate dimension $d$ is large.
-Optimal computational thresholds for the class of first-order iterative algorithms were derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020) in the same regime.
-Similar results were also proven for other computational models: Damian et al.
+Optimal computational thresholds for the class of first-order iterative algorithms were also derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020).
+Algorithmic results were also proven for other computational models: Damian et al.
 (2022) provided a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$.
-Damian et al.
-(2024) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
+Here $\ell$ denotes the information exponent (Arous et al., 2021) of the target defined as the first non-zero integer $\ell\!>\!0$ in the Hermite expansion of $g$: $\ell\!=\!\min\{j\!\in\!\mathbb{N}:\langle g,\mathcal{H}_{j}\rangle_{\gamma}\neq 0\}.$
+A related notion is that of staircase functions (Abbe et al., 2022; 2023), which characterizes the sample complexity for sequential learning of directions under the under the CSQ model or online SGD.
+For staircase functions, the CSQ sample complexity is governed by the “leap”, which informally equals the maximum jump in degree (in the Hermite expansion) conditioned on previously learned directions.
+For instance the target $g(z_{1}z_{2}z_{3})=z_{1}+z_{1}z_{2}+z_{1}z_{2}z_{3}$ has leap $1$, since the term $z_{1}z_{2}$ is linear conditioned on $z_{1}$ while $z_{1}z_{2}z_{3}$ is linear conditioned on $z_{1},z_{2}$.
+Additionally Damian et al.
+(2024) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, and a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
+As we shall see, our results allows to reconcile this picture with an extended version of staircase functions that we refer to as grand staircase functions.
 Aside from the particular case of committee machines (Aubin et al., 2019; Diakonikolas et al., 2020; Goel et al., 2020; Chen et al., 2022), results for general multi-index models $p>1$ are scarce, as they crucially depend on the way different directions are coupled by the link function.
@@
 More precisely, our main contributions are:
-  • We analyse the computational limits of weak learnability (as defined in Def.
+  • We analyse the computational limits of weak learnability (Def.
 2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\!\to\!\infty$ at fixed ratio $\alpha\!=\!\nicefrac{{n}}{{d}}$ and constant number of indices $p\!=\!\Theta(1)$.
-Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural networks trained with gradient descent.
-  • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples in multi-index models: trivial directions can be learned with a single AMP iteration for any $\alpha>0$.
-When no such direction exists, we show there might be easy directions allowing recovery from arbitrarily small initial overlap $\epsilon>0$ in $\mathcal{O}(\log\frac{1}{\epsilon})$ steps with at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition.
-Heuristically, we expect recovery along easy directions from random initialization with $\alpha>\alpha_{c}$ in $\mathcal{O}(\log d)$ iterations.
-We further confirm this numerically through extensive simulations.
+Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural nets trained with gradient descent.
+  • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples: trivial directions can be learned with a single AMP iteration for any $\alpha\!>\!0$.
+When no such direction exists, we show there might be easy directions allowing recovery from arbitrarily small side-information $\sqrt{\lambda}\!>\!0$ in $\mathcal{O}(\log\nicefrac{{1}}{{\lambda}})$ steps with at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition.
+Heuristically, we expect recovery along easy directions from random initialization with $\alpha\!>\!\alpha_{c}$ in $\mathcal{O}(\log d)$ iterations and confirm this numerically through extensive simulations.
 [^2] Note that $\alpha_{c}$ may or may not coincide with the statistically optimal threshold.
@@
 The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!\mathcal{O}(d)$ samples.
-  • We discuss how interactions among directions can lead to hierarchical learning phenomena.
+  • We discuss how interactions among directions can lead to hierarchical learning phenomena, with the hierarchy referring to learning a sequence of growing subspaces.
 Interestingly, even hard directions can be learned when coupled to easier ones through a hierarchical process, which is reminiscent of the staircase phenomenon for one-pass SGD (Abbe et al., 2023).
 However, the class of functions that are weakly learnable hierarchically with AMP differs from SGD and is, in particular, considerably larger.
-We characterise this class of Hierarchical Iterative Denoising Staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
+We characterise this class of grand staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
+Finally, the code to reproduce our plots, to run the AMP algorithm, and to deploy state evolution is available on GitHub https://github.com/SPOC-group/FundamentalLimitsMultiIndex
 Further related work —
@@
 Instead, we focus on the typical-case learnability of a given, explicit, multi-index model.
-Notation
+2 Settings and definitions
 […]
-2 Settings and definitions
-As motivated in the introduction, our main focus in this work will be to study subspace identifiability in the class of Gaussian multi-index models.
 Definition 1 (Gaussian multi-index models).
@@
 rows ${\bm{w}}^{\star}_{k}\sim\mathcal{N}(0,{\bm{I}}_{d})$.
-Note we allow the map $g:\mathbb{R}^{p}\to\mathbb{R}$ to be stochastic i.e $g(\cdot)=\tilde{g}(\cdot,{\bm{\xi}})$ for random variable ${\bm{\xi}}\in\mathbb{R}^{p}$ independent of $\mathbf{x}$ having a fixed dimension-independent law such that $\mathbb{E}\left[\norm{\xi}_{2}^{2}\right]<\infty$.
+Note we allow the map $g:\mathbb{R}^{p}\to\mathbb{R}$ to be stochastic i.e $g(\cdot)=\tilde{g}(\cdot,{\bm{\xi}})$ for random variable ${\bm{\xi}}\in\mathbb{R}^{p}$ independent of $\mathbf{x}$ having a fixed dimension-independent law such that $\mathbb{E}\left[\norm{\bm{\xi}}_{2}^{2}\right]<\infty$.
 [^4] Technically, our results hold for ${\bm{\xi}}\in\mathbb{R}^{p^{\prime}}$ with $p^{\prime}=\mathcal{O}(1)$, for example the case of a degenerated noise.
@@
 Therefore, in this work, we will be interested in weak subspace learnability, which corresponds to obtaining an estimation of the subspace spanned by ${{\bm{{W}}}}^{\star}$ better than a random estimator.
-This can be defined in an invariant way as follows
+This can be defined in an invariant way:
 Definition 2 (Weak subspace recovery).
@@
 with high probability as $d\to\infty$.
-Our main tool for characterising the computational bottlenecks in the Gaussian multi-index problem is an Approximate Message Passing (AMP) Algorithm 1 tailored to our problem.
-Note that Algorithm 1 only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and the application of a problem dependent function ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$ (defined later in Equation 4).
-Therefore, it belongs to the class of first-order algorithms with linear running time of $\mathcal{O}(nd)$.
-Intuitively, AMP estimates the mean ${\bm{\omega}}$ and variance ${\bm{V}}$ of ${{\bm{{W}}}}^{*}{\bm{x}}$ and the mean $\hat{{{\bm{{W}}}}}$ and covariance $\hat{{\bm{C}}}$ of ${{\bm{{W}}}}^{*}$ through a fixed point iteration described in Algorithm 1.
-The key property of AMP is that for a well-chosen ${\bm{g}}_{\rm out}$, it is provably optimal within the class of first-order methods (Celentano et al., 2020).
-This means that establishing learnability for the optimal AMP Algorithm 1 implies a computational lower bound in the class of first-order methods, which includes popular machine learning algorithms such as gradient descent (SG).
-For the Gaussian multi-index estimation problem 1, the optimal ${\bm{g}}_{\rm out}$ is simply given by the optimal denoiser of an effective $p$-dimensional problem $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
+Our main tool for characterising the computational bottlenecks in the multi-index problem is an Approximate Message Passing (AMP, Algorithm 1) acting on ${\bf B}\in\mathbb{R}^{d\times p}$ and ${\bf\Omega}\in\mathbb{R}^{n\times p}$ tailored to our needs:
+${\bm{\Omega}}^{t}$ $={\bm{X}}f_{t}({\bf B}^{t})-g_{t-1}({\bm{\Omega}}^{t-1},{\bf y}){\bm{V}}_{t}$ (4)
+${\bf B}^{t+1}$ $={\bm{X}}^{T}g_{t}({\bm{\Omega}}^{t},{\bf y})+f_{t}({\bf B}^{t}){\bm{A}}_{t}$ (5)
+where the denoisers $g_{t}(.,y_{i})$ and $f_{t}(.)$ are vector-valued mapping applied row-wise on ${\bf\omega}_{i}\in\mathbb{R}^{p}={\bf\Omega}_{.i}$ and ${\bf b}_{i}\in\mathbb{R}^{p}={\bf B}_{.i}$.
+Here $\hat{{\bm{{W}}}}^{t}\!\in\!\mathbb{R}^{p\times d}=f_{t}({\bf B}^{t})^{\top}$ are the estimator of the weights matrix ${{\bm{{W}}}}^{\star}$, $g_{t}({\bm{\Omega}}^{t})\in\mathbb{R}^{n\times p}$ are an estimate of the pre-activations, while ${\bm{A}}_{t}$ and ${\bm{V}}_{t}$ are the so-called Onsager terms (see Sec.A).
+AMP only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and thus belongs to the class of first-order algorithms with linear running time of $\mathcal{O}(nd)$.
+The key property of AMP is that for well-chosen $g_{t}$ and $f_{t}$ (using Bayesian denoisers) it is provably optimal within the class of first-order methods (Celentano et al., 2020).
+In our case, this require $g_{t}({\bm{\omega}},y)={\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$ as in eq.
+(6), while $f_{t}$ is just a Gaussian Bayesian denoiser (see Sec.A).
+Learnability for the optimal AMP thus implies a computational lower bound in the class of first-order methods, including in particular popular machine learning algorithms such as gradient descent (SG).
+For the Gaussian multi-index estimation problem 1, ${\bm{g}}_{\rm out}$ is given by the optimal denoiser of an effective $p$-dimensional problem $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$, which reads
 […]
@@
 For stochastic mappings $g(\cdot)=g(\cdot,\xi)$, the expectation is w.r.t both $Z,\xi$ conditioned on $Y$.
-Here ${\bm{\omega}}$ can be interpreted as an estimate of the pre-activations.
-Our results rely on the following assumption on the above denoiser, translating to an indirect assumption on $g,\xi$:
+Our results rely on the following assumption on the above denoiser, translating to weak assumptions on $g,\xi$:
 Assumption 1.
-${\bm{g}}_{\rm out}:\mathbb{R}^{p^{2}+p+1}\rightarrow\mathbb{R}^{p}\in\mathcal{C}^{2}$.
-Furthermore, the mapping $\hat{g}_{\rm out}:\mathbb{R}^{p+{p^{\prime}}}\rightarrow\mathbb{R}^{p}$ defined by ${\hat{g}}_{\rm out}(\mathbf{h},\xi)={\bm{g}}_{\rm out}(\tilde{g}((\mathbf{h},\xi),{\bm{\omega}},{\bm{V}})$ is pseudo-Lipschitz of finite-order for all ${\bm{\omega}},{\bm{V}}\in\mathbb{R}^{p},\mathbb{R}^{p\times p}$.
-Here, $\xi\in\mathbb{R}^{p^{\prime}}$ denotes independent noise satisfying $\mathbb{E}\left[\norm{\xi}^{2}\right]<\infty$ as per Definition 1.
+(a) ${\bm{g}}_{\rm out}\!:\!\mathbb{R}^{p^{2}+p+1}\rightarrow\mathbb{R}^{p}\!\in\!\mathcal{C}^{2}$.
+(b) the mapping $\hat{g}_{\rm out}:\mathbb{R}^{p+{p^{\prime}}}\rightarrow\mathbb{R}^{p}$ defined by ${\hat{g}}_{\rm out}(\mathbf{h},{\bm{\xi}})={\bm{g}}_{\rm out}(\tilde{g}((\mathbf{h},{\bm{\xi}}),{\bm{\omega}},{\bm{V}})$ is pseudo-Lipschitz of finite-order for all ${\bm{\omega}},{\bm{V}}\in\mathbb{R}^{p},\mathbb{R}^{p\times p}$.
+Here, ${\bm{\xi}}\!\in\!\mathbb{R}^{p^{\prime}}$ denotes independent noise satisfying $\mathbb{E}\left[\norm{\bm{\xi}}^{2}\right]\!<\!\infty$ as per Def.
+1.
 The above denoiser can also be derived as an exact high-dimensional approximation of the Belief Propagation (BP) algorithm for the estimation of the marginals of the Bayesian posterior distribution:
-$p({{\bm{{W}}}}|{\bm{X}},{\bm{y}})\propto\prod\limits_{i=1}^{n}\delta(y_{i}-g({{\bm{{W}}}}{\bm{x}}_{i}))\prod_{k=1}^{p}\mathcal{N}({\bm{w}}_{k}|0,{\bm{I}}_{d}).$ (5)
-Algorithm 1 was derived for Gaussian multi-index models by Aubin et al.
+$p({{\bm{{W}}}}|{\bm{X}},{\bm{y}})\propto\prod\limits_{i=1}^{n}\delta(y_{i}-g({{\bm{{W}}}}{\bm{x}}_{i}))\prod_{k=1}^{p}\mathcal{N}({\bm{w}}_{k}|\mathbf{0},{\bm{I}}_{d}).$ (7)
+The Bayes-AMP algorithm was derived for Gaussian multi-index models by Aubin et al.
 (2019), although it has been analysed only for the particular link function $g({\bm{z}})={\rm sign}(\sum_{k\in[p]}{\rm sign}(z_{k}))$, known as the committee machine.
 Our goal in this work is to leverage Algorithm 1 in order to provide a sharp classification of which link functions $g$ are computationally challenging to learn.
+Our analysis is based on two remarkable properties that make AMP a particularly useful tool for studying high-dimensional estimation.
+The first property is that for any $t=\Theta(1)$, in the high-dimensional limit $d\to\infty$ the performance of AMP can be tracked without actually running the algorithm.
+This result, known as state evolution, makes AMP mathematically tractable in high-dimensions (Bayati and Montanari, 2011).
+The second property is its optimality with respect to Bayesian estimation (see SI, A).
+For multi-index models, the state evolution equations were derived by Aubin et al.
+(2020b) and rigorously proven by Gerbelot and Berthier (2023).
+It provides, in particular, an exact characterization of the asymptotic overlaps and prediction error:
+Lemma 2.1 (State evolution (Aubin et al., 2019; Gerbelot and Berthier, 2023)).
+[…]
+[…]
+with ${{\bm{M}}}^{t}$ satisfying the state evolution equations from initial condition ${{\bm{M}}}^{0}$ iterated with ${{\bm{M}}}^{t+1}\!=\!F({{\bm{M}}}^{t})$:
+$F({{\bm{M}}}^{t})={G}\!\left(\!\alpha\,\mathbb{E}\Big{[}{\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\otimes 2}\Big{]}\right).$ (9)
+where $G({\bm{X}})=({\bm{I}}_{p}+{\bm{X}})^{-1}{\bm{X}}$ and the expectation is taken over the following effective process
+$Y^{t}=g\left(({\bm{I}}_{p}-{{\bm{M}}}^{t})^{\nicefrac{{1}}{{2}}}{\bm{\xi}}+{{{\bm{M}}}^{t}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}\right),$ (10)
+with ${\bm{Z}},{\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ independently from ${\bm{Z}}$. The asymptotic mean-squared error on the label prediction is then given by:
+[…]
+[…]
+3 The trivial subspace ($\alpha_{c}=0$)
+The state evolution Equation 9 maps the problem of characterising the computational bottlenecks of first-order methods for high-dimensional Gaussian multi-index models to the study of the deterministic, $p$-dimensional dynamical system ${{\bm{M}}}^{t+1}=F({{\bm{M}}}^{t})$.
+A starting point is identifying its fixed points and their basins of attraction.
+In the absence of any prior information on ${{\bm{{W}}}}^{\star}$ aside from its distribution, one cannot do better than taking $\hat{{\bm{w}}}_{k}^{t=0}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{d})$ with $k\in[p]$ independently at random from the prior.
+With high-probability, at initialization, the elements of the overlap matrix $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ are $\Theta(d^{-\nicefrac{{1}}{{2}}})$ element-wise.
+The asymptotic overlap for an uninformed initial condition is thus ${{\bm{M}}}^{0}=0$, a null-rank matrix.
+If ${{\bm{M}}}^{0}=0$ is not a fixed point, then ${{\bm{M}}}^{1}\succ 0$, implying the weak recovery of a subspace of dimension $k={\rm rank}({{\bm{M}}}^{1})>0$ with just a single step of AMP.
+Lemma 3.1 (Existence of uninformed fixed point).
+${{\bm{M}}}=\mathbf{0}\in\mathbb{R}^{p\times p}$ is a fixed point of Equation 9 if and only if the following condition holds almost surely over $Y$:
+${\bm{g}}_{\rm out}(Y,{\bm{\omega}}=\mathbf{0},{\bm{V}}={\bm{I}}_{p})=\mathbb{E}[{\bm{Z}}|Y]=\mathbf{0},$ (11)
+As long as the conditional expectation above is not zero almost surely, then AMP weakly learns a non-empty subspace immediately in the first iteration for any number of samples $n=\Theta(d)$.
+For this reason, we refer to this subspace as a trivial subspace:
+Definition 3 (Trivial subspace).
+We define $H^{\star}_{T}\subseteq\mathbb{R}^{p}$ as the subspace spanned by $\mathbf{v}\in\mathbb{R}^{p}$ such that the following holds almost surely over $Y=g({\bm{Z}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
+${\bm{g}}_{\rm out}(Y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}[\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle|y=Y]=0,$ (12)
+where the expectation is w.r.t the joint measure $p(\mathbf{x},y)$ defined in Def. 1 eq.(10). The trivial subspace $T^{\star}$ is the orthogonal complement of $H^{\star}_{T}$.
+Theorem 3.2.
+For any $\alpha>0$, with high-probability as $d\rightarrow\infty$, the AMP algorithm with the Bayes-optimal choice of $f_{t},g_{t}$ recovers $T^{\star}$ as per Definition 2 in a single iteration.
+Note that for single-index models ($p=1$), the condition in Equation 11 reduces to the one derived by (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020).
+Interestingly, this is exactly the same condition appearing in Damian et al.
+(2024) for weak learnability of single-index models in the SQ model; see eq.
+(3) therein.
+To make Lemma 3.1 concrete, let us look at a few examples (A detailed derivation of these examples is presented in Appendix F):
+  1. 1.
+    For single-index models ($p=1$), $T^{*}$ is one dimensional if and only if $g$ is non-even, e.g.
+$g(z)={\rm He}_{3}(z)$.
+This follows from requiring that $g_{\rm out}(y,0,1)\neq 0$ for at least one value of $y$.
+In particular, on any open interval where $g_{\rm out}$ is invertible we have $g_{\rm out}=g^{-1}$.
+  2. 2.
+    For a linear multi-index model, $g({\bm{z}})=\sum_{i=1}^{p}z_{i}$, $T^{\star}$ is spanned by $\mathbf{1}_{p}\in\mathbb{R}^{p}$ (all-one vector).
+  3. 3.
+    For a committee $g({\bm{z}})\!=\!\sum_{i=1}^{p}{\rm sign}(z_{i})$, the trivial subspace $T^{\star}$ is again 1-d, spanned by $\mathbf{1}\!\in\!\mathbb{R}^{p}$.
+  4. 4.
+    For monomials $g({\bm{z}})=z_{1}\dots z_{p}$, the trivial subspace $T^{\star}$ is non-empty if and only if $p=1$.
+  5. 5.
+    For leap one staircase functions (Abbe et al., 2023):
+    $g({\bm{z}})=z_{1}+z_{1}z_{2}+z_{1}z_{2}z_{3}+\dots$ (13)
+    The trivial subspace is $T^{\star}=\mathbb{R}^{p}$ and is spanned by the canonical basis. In other words, AMP learns all the directions with a single step for any $\alpha>0$.
+Lemma 3.1 can also be related to computational models based on queries, such as SQ learning (Kearns, 1998): the denoiser ${\bm{g}}_{\rm out}$ can indeed be interpreted as a non-linear transformation on the labels $y\!\mapsto\!{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$.
+From this perspective, the statement on the condition for the existence of a non-empty trivial subspace (12) translates to the condition $\mathbb{E}\left[{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=\mathbb{E}\left[\mathbb{E}[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|Y=y]^{2}\right]\neq 0$ where ${\bm{v}}\in T^{\star}$.
+The left-hand side can be seen as a statistical query of the type $\mathbb{E}[\varphi(y)\psi(x)]$ with label pre-processing $\varphi={\bm{g}}_{\rm out}$.
+The fact that this linear correlation in the transformed labels is non-vanishing implies that one can weakly recover $\mathbf{v}$ through a tailored spectral method (Mondelli and Montanari, 2019; Luo et al., 2019).
+In fact the denoiser ${\bm{g}}_{\rm out}$ is the optimal such transformation in the sense that when ${\bm{g}}_{\rm out}$ fails to obtain a linear correlation along $\mathbf{v}$, i.e when $\mathbf{v}\in H^{\star}$, then no transformation can:
+Lemma 3.3.
+$\mathbf{v}\in H^{\star}$ if and only if for any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
+$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=0.$ (14)
+Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem).
+Equivalently, $\mathbf{v}\in T^{\star}$ if and only if there exists a transformation such that $\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]\neq 0$.
+Remarks on GD —
+The characterization of $H^{\star}$ and $T^{\star}$ allows us to contrast the (optimal) performance of AMP with that of gradient-descent based methods on neural nets: First, note that when $T^{\star}$ is empty, then $\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]\!=\!0$ for all $\bm{v}\in\mathbb{R}^{p}$.
+This is precisely the condition for the generative exponent (Damian et al., 2024) being strictly larger than $1$.
+This implies that one-pass SGD —even after applying any transformation $\mathcal{T}$— fails to obtain weak recovery with $\mathcal{O}(d)$ samples (see Thms.
+1.3 and 1.4 in Arous et al.
+(2021)).
+The converse question is interesting: For non-emtpy $T^{\star}$, can GD recover $T^{\star}$ with just $\mathcal{O}(d)$ steps?
+In the absence of any transformations of the labels, this remains false for online SGD (Arous et al., 2021; Abbe et al., 2022), unless the information exponent of $g$ is $1$.
+The situation is different when reusing batches: Dandi et al.
+(2024) indeed showed that full batch GD on a two-layer neural network implicitly applies transformation $\mathcal{T}$ to the labels.
+The learnable subspace in two iterations with $\mathcal{O}(d)$ sample complexity then exactly coincides with that of $T^{\star}$ in Lemma 3.3, up to the restriction to polynomial transformations.
+A two-layer network with re-use of batches can thus recover the full trivial subspace efficiently for any $\alpha>0$.
+4 Phase transitions ($\alpha_{c}\!>\!0$)
+When the trivial subspace $T^{\star}$ from Def.
+3 is empty, ${{\bm{M}}}=\mathbf{0}$ is a fixed point of state-evolution, thus AMP starting from random initialization fails to recover any subspace of $W^{\star}$ in any finite number of steps.
+However, there may exist directions where recovery requires only an infinitesimal additional side-information.
+To contrast with the trivial subspace, we refer to the directions learnable for arbitrarily small side-information as the easy directions.
+The learnability of easy subspaces is characterized by AMP’s stability, which crucially depends on the sample complexity $\alpha\!=\!n/d$.
+The stability of ${{\bm{M}}}=\mathbf{0}$ is studied by linearising the state evolution (9) around $\delta{{\bm{M}}}$ with $\norm{\delta{{\bm{M}}}}\approx 0$
+$F({{\bm{M}}})\approx\alpha\mathcal{F}(\delta{{\bm{M}}})+\mathcal{O}(\norm{\delta{{\bm{M}}}}^{2})\,,$
+where $\mathcal{F}(\delta{{\bm{M}}})$ is a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$:
+${\!\!\!\!\!\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\mathbb{E}_{Y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,\mathbf{0},{\bm{I}}_{p})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,\mathbf{0},{\bm{I}}_{p})^{\top}\right].}$ (15)
+The above linearization demarcates a threshold on $\alpha$, below which the AMP remains stable:
+Lemma 4.1.
+(Stability of the uninformed fixed point).
+If ${{\bm{M}}}=\mathbf{0}\in\mathbb{R}^{p\times p}$ is a fixed point, then it is an unstable fixed point of Equation 9 if and only if $\norm{\mathcal{F}({{\bm{M}}})}_{F}>0$ and $n>\alpha_{c}d$, where the critical sample complexity $\alpha_{c}$ is:
+$\frac{1}{\alpha_{c}}=\sup_{{{\bm{M}}}\in\mathcal{S}^{+}_{p}}\norm{\mathcal{F}({{\bm{M}}})}_{F},$ (16)
+with $\|\cdot\|_{F}$ the Frobenius norm, and the extremum being achieved at a unique extremizer ${{\bm{M}}}^{\star}\in\mathcal{S}^{+}_{p}$.
+Moreover, if $\norm{\mathcal{F}({{\bm{M}}})}_{F}=0$, then ${{\bm{M}}}=\mathbf{0}$ is a stable fixed point for any $n=\Theta(d)$.
+The above result follows from a generalization of the Perron-Frobenius theorem to the cone $\mathcal{S}^{+}_{p}$; see Section D.3 for a proof.
+Analogously, the operator $\mathcal{F}({{\bm{M}}})$ allows us to define the stability threshold along a specific direction $\mathbf{v}\in\mathbb{R}^{p}$.
+We define the easy subspace as containing the directions where ${{\bm{M}}}=\mathbf{0}$ becomes unstable for some $\alpha<\infty$:
+Definition 4 (Easy subspace $E^{\star}$).
+Let $H^{*}_{E}$ be the subspace spanned by directions $\mathbf{v}\in\mathbb{R}^{p}$ such that:
+$\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}}=\mathbf{0},{\bm{V}}={\bm{I}}_{p})\mathbf{v}=0,$ (17)
+almost surely over $Y$. We define the easy subspace $E^{\star}$ as the orthogonal complement of $H^{\star}_{E}$.
+Modeling side information as additional observations:
+${\bf S}=\sqrt{\lambda}{\bf W^{\star}}+\sqrt{1-\lambda}{\bm{Z}},$ (18)
+where ${\bm{Z}}\!\in\!\mathbb{R}^{p\times d}$ has independent entries $Z_{ij}\!\sim\!\mathcal{N}(0,1)$, we can now state our main theorem:
+Theorem 4.2.
+Let ${{\bm{M}}}_{d}^{t}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ denote the model-target overlap matrix at any finite time $t$.
+Suppose that $T^{\star}={0}$ and consider the AMP algorithm with the Bayes-optimal choice of $f_{t},g_{t}$.
+Then, with high probability as $d\rightarrow\infty$:
+  • (i)
+    For $\alpha\geq\alpha_{c}$, $\exists\delta>0$ such that for any sufficiently small $\lambda$ ,${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}^{\star}$ for $t=\mathcal{O}(\log\nicefrac{{1}}{{\lambda}})$, where ${{\bm{M}}}^{\star}$ is as defined in Lemma 4.1.
+Furthermore, there exists an $\alpha\geq\alpha_{c}$ and a $\delta>0$ such that ${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}_{E^{*}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\lambda}})$ iterations, where ${{\bm{M}}}_{E^{\star}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}$.
+  • (ii)
+    For $\alpha<\alpha_{c}$ however, ${{\bm{M}}}^{t}_{d}=\mathbf{0}$ is asymptotically stable i.e.
+there exist constants $\lambda^{\prime}<1$ and $C>0$ such that for $\lambda<\lambda^{\prime}$, $\sup_{t\geq 0}\norm{{{\bm{M}}}^{t}_{d}}\leq C\sqrt{\lambda}$.
+Concretely, this implies that for $\alpha<\alpha_{c}$, not only does AMP fail to find any pertinent directions, but it also fails to improve on the small side-information.
+For $\alpha>\alpha_{c}$, however, AMP will develop a growing overlap along a non-empty subspace starting with arbitrarily small (but finite) side information.
+Based on the optimality of AMP amongst first-order methods for any finite number of iterations (Celentano et al., 2020; Montanari and Wu, 2022), $\alpha_{c}$ thus marks the onset of computational phase transition (see S.I.
+E) for weak learnability in the first-order methods class.
+Note that in Thm.
+4.2 the side information snr $\lambda>0$ can be arbitrarily small, but must remain $\mathcal{O}(1)$.
+Based on a vast array of evidence from statistical physics literature, we are actually claiming a stronger result, and believe the following conjecture to be exact:
+Conjecture 1.
+AMP initialized randomly will find a finite overlap with the easy directions for $\alpha>\alpha_{c}$ in ${\cal O}(\log d)$ steps, without side information.
+This can be justified by the heuristic $\lambda=\mathcal{O}(\nicefrac{{1}}{{\sqrt{d}}})$ (as this is the scaling of the correlation with the ground truth at initialization) in Thm.
+4.2.
+While is not strictly covered by the state evolution theorem that allow only a finite number of steps, it is however well obeyed in practice in our simulations (see Fig.2) and it is a well known observation in many message-passing works that AMP (or belief propagation) find the fixed point after few iterations from a random start (see e.g.
+Decelle et al.
+(2011); Zdeborová and Krzakala (2016); Barbier et al.
+(2019)).
+A rigorous proof, though, remain a difficult open problem that require a non-asymptotic control of the AMP state evolution, see (Rush and Venkataramanan, 2018; Li and Wei, 2022; Li et al., 2023) for some recent progress in the context of $\mathbb{Z}_{2}$ synchronization where the convergence from random start has been established, corroborating the heuristic identification $\lambda=\mathcal{O}(\nicefrac{{1}}{{\sqrt{d}}})$ in this case.
+An alternative possible strategy to provide a recovery algorithm in $\mathcal{O}(\log(d))$ steps would be to follow (Krzakala et al., 2013; Maillard et al., 2022) and linearize AMP to reach a power iteration method (with a matrix similar to the one discussed in Lemma 3.3) so that the resulting operator can be studied with random matrix theory (Guionnet et al., 2023; Mondelli and Venkataramanan, 2021), as was done for the single-index model in (Mondelli and Montanari, 2019; Lu and Li, 2020; Maillard et al., 2022).
+We leave these, and the proof of conjecture $1$, open for further studies.
+The expression for the weak recovery threshold Eq.(16) in Lemma 3.1 allows to generalise the single-index expression from (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020).
+Indeed, for $p=1$, we have $\partial_{\omega}g_{\rm out}(y,0,1)=\mathbb{E}[{\rm He}_{2}(Z)|Y=y]$, and therefore $\alpha_{c}=\mathbb{E}_{Y}\left[\mathbb{E}[{\rm He}_{2}(Z)|Y]^{2}\right]^{-1}$ or equivalently
+[…]
+which is exactly the threshold in (Barbier et al., 2019; Mondelli and Montanari, 2019; Lu and Li, 2020; Maillard et al., 2022; Damian et al., 2024).
+We now illustrate Lemma 4.1 in a two examples of interest:
+(i) The monomial $g({\bm{z}})=z_{1}\dots z_{p}$ with $p>1$ can always be learned with $\alpha>\alpha_{c}(p)$ large enough (Chen and Meka, 2020).
+For instance, we have $\alpha_{c}(2)\approx 0.5937$, $\alpha_{c}(3)\approx 3.725$ and $\alpha_{c}(4)\approx 4.912$.
+In Appendix F we derive an analytical formula for $\alpha_{c}(p)$ for arbitrary $p$, and show that $\alpha_{c}(p)\sim p^{1.2}$ for large $p$.
+(ii) The embedding of the sparse parity functions:
+$g({\bm{z}})=\prod_{k=1}^{p}{\rm sign}(z_{k})\,.$ (20)
+As this is invariant under permutations of the indices, Lemma 4.1 implies the existence of a computational phase transition.
+In S.I.
+Sec.F.3 we compute analytically the critical value $\alpha_{c}(p)$: $p\!=\!1$ is equivalent to phase retrieval and $\alpha_{c}(1)\!=\!\nicefrac{{1}}{{2}}$.
+For $p\!=\!2$, we show that $\alpha_{c}(2)\!=\!\nicefrac{{\pi^{2}}}{{4}}$, while $\alpha_{c}(p)\!=\!+\infty$ for $p\!\geq\!3$.
+This is illustrated in Fig.1 (left) for 2-sparse parity.
+Figure 1: Weak learnability phase transitions for $g(z_{1},z_{2})\!=\!{\rm sign}(z_{1}z_{2})$ (left) and $g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ (center and right).
+Given the permutation symmetry in the models, we display the optimal permutation of the overlap matrix elements reached by AMP.
+(Left): Overlaps with the two directions $\nicefrac{{1}}{{2}}(M_{11}+M_{22})$ as a function of the sample complexity $\alpha=\nicefrac{{n}}{{d}}$, with the phase transition at $\alpha_{c}=\pi^{2}/4$.
+The solid black line is the asymptotic theory from state evolution while crosses are averages over $72$ AMP runs with $d\!=\!500$.
+(Center): Overlaps with the first direction $|M_{11}|$ ( blue), and with the second and third one $\nicefrac{{1}}{{2}}(M_{22}+M_{33})$ ( red) as a function of the sample complexity $\alpha=\nicefrac{{n}}{{d}}$.
+Solid lines are the state evolution curves Equation 9, and crosses/dots AMP runs with $d=500$ averaged over $72$ seeds.
+All other overlaps are zero (black).
+The two black dots indicate the critical thresholds at $\alpha_{1}\approx 0.575$ and $\alpha_{2}=\pi^{2}/4$.
+(Right) Corresponding generalization error as a function of $\alpha=n/d$.
+The figure can be reproduced using the code provided (See also Appendix G).
+The sparse-parity function is a classic example of a computationally hard function in theoretical computer science (Blum et al., 1994; 2003).
+[^5] Noiseless parity is actually efficiently learnable by Gaussian elimination (Blum et al., 1994).
+Here we consider computational efficiency in a noise-tolerant sense.
+Therefore, it should come as no surprise that $\alpha_{c}(p)=\infty$ for $p\geq 3$, providing a concrete example of what a hard subspace (i.e.
+that cannot be learned with $n=\Theta(d)$ samples) looks like for AMP.
+Similar to the discussion of trivial directions in Lemma 3.3, we can invoke optimality of AMP to translate our result on the optimal denoiser $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$ to a statement for queries on general label pre-processing transformations:
+Lemma 4.3.
+For any $\mathbf{v}\in H^{\star}_{E}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$ we have
+$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0\,.$ (21)
+Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem from eq.(10)).
+Remarks on GD —
+Analogous to Lemma 3.3, Lemma 4.3, implies that the condition $T^{\star},E^{\star}$ being empty (i.e.
+$\alpha_{c}=\infty$ is equivalent to the generative exponent (Damian et al., 2024) being strictly larger than $2$.
+The results in Arous et al.
+(2021); Abbe et al.
+(2023) then imply that even after applying any transformation to the output (as allowed in SQ) online-SGD requires at least $\Omega(d^{2})$ samples to achieve weak-rovery, since online-SGD with squared loss requires $\Omega(d^{\kappa-1})$ samples/iterations for weak-recovery of a target function with leap/information exponent $\kappa$ (Theorems 1.3 and 1.4 in Arous et al.
+(2021)).
+Furthermore, for single-index models satisfying the condition in Lemma 4.3 (i.e $\alpha_{c}\!=\!\infty$), the SQ lower-bound predicts a sample complexity requirement of at-least $\Omega(d^{3/2})$ (Damian et al., 2024).
+Learning such functions with $\mathcal{O}(d)$ samples is thus conjectured to be algorithmically-hard.
+It is again a natural question to ask for non-empy $E^{\star}$, how many samples and iterations do neural networks trained with GD require for achieving weak-recovery.
+In the absence of label transformations, weak-recovery along $E^{\star}$ requires $\Omega(d^{2})$ unless the information exponent is at-most $2$ (Arous et al., 2021; Abbe et al., 2023).
+However, in the presence of transformations, the sample complexity for weak-recovery along $E^{\star}$ can be reduced to $d\operatorname{polylog}(d)$.
+Chen et al.
+(2022) already showed that any polynomial can be learned with $d\operatorname{polylog}(d)$ iterations using an SQ algorithm.
+More-recently, (Arnaboldi et al., 2024; Lee et al., 2024) suggested that $E^{\star}$ can be recovered in $d\operatorname{polylog}(d)$ steps through modified gradient-based algorithms re-using data.
+It remains an open question, however, to see if $\alpha_{d}d$ data are enough for an agnostic GD approach.
+5 The grand staircase
+Figure 2: Trajectories of a single finite-size run of AMP with $d=500$ at $\alpha=4$ for $g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$.
+(Left) Evolution of the overlaps.
+We display $M_{11}$ in blue, $\nicefrac{{1}}{{2}}(M_{22}+M_{33})$ in red, and the off-diagonal overlaps in black.
+(Right) Evolution of the generalisation error.
+The present classification of trivial, easy or hard directions, based upon the notion of weak learnability, does not say what can be learned after a given subspace is learned.
+We now address this question.
+Suppose that the estimator $\hat{{{\bm{{W}}}}}^{t}$ has developed an overlap along a subspace belonging to the span of ${{\bm{{W}}}}^{\star}$, resulting in a non-zero overlap ${{\bm{M}}}^{t}\succ 0$.
+From that point, the main difference with respect to the previous discussion is that the variable ${\bm{\omega}}$ in the linear operator $\mathcal{F}$ defined in eq.(15) becomes non-zero (since it is distributed as ${\bm{\omega}}=\sqrt{{{\bm{M}}}}{\bm{\xi}}$).
+Crucially, this changes the span of $\mathcal{F}({{\bm{M}}})$ and hence the stability condition in Lemma 4.1.
+In particular, as we show next, learning some directions might facilitate learning larger subspaces.
+This is reminiscent of the specialisation transition in committee machines (Saad and Solla, 1996; Aubin et al., 2020b) and of the staircase for SGD (Abbe et al., 2022):
+Definition 5.
+Let $U\subset\mathbb{R}^{p}$. We define $H_{T}^{\star}(U)$ to be the subspace spanned by ${\bm{v}}\in U^{\perp}$ such that
+${\bm{v}}^{\top}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}}_{U}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U})=0$ (22)
+almost surely over ${\bm{\xi}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
+We define the “trivially-coupled” subspace $T^{\star}_{U}$ for $U$ as the orthogonal complement of $H_{T}^{\star}(U)$.
+Analogously, let $H^{\star}_{E}(U)$ be the subspace spanned by directions ${\bm{v}}\in U^{\perp}$ such that
+${\bm{v}}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U}){\bm{v}}=0$ (23)
+almost surely over ${\bm{\xi}}$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
+When ${{\bm{M}}}_{U}$ is additionally a fixed point of $\mathcal{F}_{{{\bm{M}}}}$, one can linearise $\mathcal{F}_{{{\bm{M}}}}$ along the orthogonal complement of $U$.
+We define the easy-coupled subspace $E^{\star}_{U}$ for $U$ as the orthogonal complement of $H^{\star}_{E}(U)$.
+Next, suppose that ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ with $\operatorname{span}({{\bm{M}}}_{U})=U$ is a fixed-point of $\mathcal{F}_{{{\bm{M}}}}$.
+Let $\mathcal{F}_{{{\bm{M}}}_{U}}$ denote the linearization of $F({{\bm{M}}})$ along the orthogonal complement $U^{\perp}$ at ${{\bm{M}}}={{\bm{M}}}_{U}$.
+We define the grand staircase threshold $\alpha_{\text{gst}}({{\bm{M}}}_{U})$ at ${{\bm{M}}}={{\bm{M}}}_{U}$ as $\alpha_{\text{gst}}({{\bm{M}}}_{U})=(\sup_{{{\bm{M}}}^{\perp}\in U^{\perp}}\norm{\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}})}_{F})^{-1}$.
+We denote by ${{\bm{M}}}^{\star}_{U}$ the extremiser achieving the supremum.
+The definitions above generalise the notions of trivial and easy subspaces in Def.
+3 & 4 conditionally on a subspace $U$ that has been previously learned, and characterise the directions whose recovery is enabled upon learning the subspace $U$.
+Concretely, upon developing an initial overlap along $U$, the directions in $T^{\star}_{U},E^{\star}_{U}$ can be recovered analogous to the recovery of $T^{\star},E^{\star}$ in Theorems 3.2, 4.2.
+For brevity, we focus the present discussion on the recovery of $E^{\star}_{U}$ with vanishing side information, and discuss settings involving the recovery of $T^{\star}_{U}$ in Appendix C.
+Proposition 5.1.
+(informal) Let $U\subseteq\mathbb{R}^{p}$ be a subspace such that $E^{\star}_{U}$ is non-empty.
+Consider AMP iterates with the Bayes-optimal choice of $f_{t},g_{t}$ for sufficiently small $\lambda>0$.
+Suppose that ${{\bm{M}}}_{d}^{t}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ is an approximate fixed point of $F({{\bm{M}}})$ in Equation 9 such that ${{\bm{M}}}_{d}^{t}\approx{{\bm{M}}}_{U}$ where $M_{U}$ spans $U$.
+Then:
+  • For $\alpha>\alpha_{\text{gst}}({{\bm{M}}}_{U})$, AMP recovers $M^{*}_{M_{U}}$ in additional $\mathcal{O}(\log\frac{1}{\lambda})$ steps for arbitrarily small $\lambda$, where $M^{*}_{M_{U}}$ is as defined in eq.
+(5).
+  • For $\alpha<\alpha_{\text{gst}}({{\bm{M}}}_{U})$ and sufficiently small $\lambda$, AMP remains at the approximate fixed point ${{\bm{M}}}_{U}$ and fails to gain weak-recovery along $U^{\bot}$.
+When the easy subspace $E^{\star}$ defined by Def.
+4 and $E^{\star}_{E^{\star}_{U}}$ are non-empty, we further show in S.I.
+D.6 that for large enough $\alpha\geq\alpha_{c}$ AMP recovers $E^{\star}_{U}$ by first reaching an approximate fixed point spanning $E^{\star}$ and subsequently “escaping” along $E^{\star}_{E^{\star}_{U}}$ as described above.
+A concrete example of a function displaying this phenomenon is a linear combination between hard parity function and an easy polynomial:
+$g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ (24)
+The sign part is a sparse parity with $p=3$, which cannot be learned with $n=\Theta(d)$ samples, but the quadratic polynomial $z_{1}^{2}$ component in the function allows weak-recovery of the first component i.e.
+$U={(1,0,0)}$ as long as $\alpha>1/2$.
+Hence, conditionally on $U$ the effective multi-index model becomes ${\rm sign}(z_{2}z_{3})$, which as discussed in Sec.4 is an easy function.
+Figure 1 illustrates this: first, $z_{1}$ is learned at $\alpha_{1}\approx 0.575$.
+Then, for larger value when $\alpha>\alpha_{2}$, all directions are learned (see figure 1).
+Knowing $z_{1}$ makes the hard 3-parity an easy 2-parity problem.
+It is easy to construct a multi-index model where AMP will iterate over any number of such plateaus. For instance, consider the model:
+$g({\bm{z}})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})+{\rm sign}(z_{3}z_{4}z_{5})+\ldots\,.$ (25)
+After the first plateau to learn $z_{1}$, there will be one for $z_{2}$ and $z_{3}$, then for $z_{4}$, $z_{5}$.
+Another example (see Fig.
+3 in the S.I.) is the neural net target $g(z_{1},z_{2},z_{3})\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$.
+At first only the one-dimensional direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is learned for any $\alpha>0$.
+This is simply a linear approximation of the function.
+Only at $\alpha\!\approx\!4.3$ does the network learns the other directions.
+Known as the the specialisation transition, this provide a very early example of such sequential learning (Saad and Solla, 1996; Aubin et al., 2020a).
+We can characterize a sequence of growing subspaces that AMP recovers iteratively, with vanishing side information, for large enough $\alpha$, as follows:
+Definition 6.
+Let $E^{\star}_{1}=E^{\star}\cup T^{\star}$, and inductively define for $k\in\mathbb{N}$:
+$E^{\star}_{k+1}=E^{\star}_{k}\cup E^{\star}_{E^{\star}_{k}}\cup T^{\star}_{E^{\star}_{k}}$ (26)
+The dimensionality of the subspaces in the above sequence is non-decreasing and therefore, $\exists k$ such that $E^{\star}_{E^{\star}_{k}}=\emptyset$.
+We denote this maximal subspace by $E^{\star}_{\text{gst}}$ and say that $g$ is grand-staircase learnable if $E^{\star}_{\text{gst}}=\mathbb{R}^{p}$.
+Using proposition 5.1 and certain simplifying properties of the Bayes-AMP dynamics, we can show (see S.I.) that for large enough $\alpha$, AMP recovers $E^{\star}_{\text{gst}}$:
+Theorem 5.2.
+For any $\delta$, there exists $\alpha>\alpha_{c}$ such that with arbitrarily small side-information $\sqrt{\lambda}$, with high probability as $d\rightarrow\infty$, ${{\bm{M}}}_{d}^{t}\succ\delta M_{E^{\star}_{\text{gst}}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\lambda}})$ iterations, where $M_{E^{\star}_{\text{gst}}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}_{\text{gst}}$.
+Final remarks —
+We stress that grand staircase functions are not equivalent to the staircase ones of (Abbe et al., 2022; 2023). Consider for instance
+$g({\bm{z}})=\rm He_{4}(z_{1})+{\rm sign}(z_{1}z_{2}z_{3})\,.$ (27)
+As for (24), it can be AMP-learned with $\mathcal{O}(d)$ samples by first recovering the direction corresponding to $z_{1}$.
+Unlike (24), however, which is “leap 2” under the CSQ staircase classification, $g({\bm{z}})$ is “leap 3”.
+The CSQ lower-bound therefore implies a sample complexity-requirement of $\Omega(d^{1.5})$ with online-SGD learning it in $\mathcal{O}(d^{2})$ samples.
+We expect grand staircase functions to be efficiently learnable by two-layers nets if data reusing is allowed (while only standard staircase ones are learned with single-pass SGD).
+Arnaboldi et al.
+(2024) recently provided strong evidence that they can indeed be learned in $\mathcal{O}(d\log d)$ steps.
+The analysis of Joshi et al.
+(2024) is compatible with this conjecture.
+We hope our study will spark more work in this direction.
+Acknowledgements
+The authors would like to thank Joan Bruna, Theodor Misiakiewicz, Luca Pesce, and Nati Srebro for stimulating discussions.
+This work was supported by the Swiss National Science Foundation under grants SNSF SMArtNet (grant number 212049) and SNSF OperaGOST (grant number 200390) and the Choose France - CNRS AI Rising Talents program.
+Supplemental material
+Appendix A Bayes optimality of AMP
+We present in Alg.1 a detailed version of the AMP algorithm described in the text using the explicit notation of Aubin et al.
+[2019].
+While it was derived by Aubin et al.
+[2019] as a limit of Belief-Propagation, it can also be directly infered from a multi-dimensional version of the standard AMP/GAMP for linear models Rangan [2011], Javanmard and Montanari [2013], Zdeborová and Krzakala [2016].
+As mention in the main, AMP Algorithm 1 is tailored to sample by estimating the posterior marginals Equation 7.
+AMP estimates the mean $\omega$ and variance $\mathbb{V}$ of $\mathbf{W}^{*}\mathbf{x}$ and the mean $\widehat{\mathbf{W}}$ and covariance $\widehat{\mathbf{C}}$ of $\mathbf{W}^{*}$ through a fixed point iteration.
+Note that the Onsager terms ${\bm{V}}_{i}$ and ${\bm{A}}_{k}$ can be made independent of the matrix by noting that the sum concentrate on the average of $\hat{C}$ and $\partial g$ respectively.
 Algorithm 1 Multi-index AMP
@@
   Return: Estimators $\hat{{{\bm{{W}}}}}_{\rm amp}\in\mathbb{R}^{p\times d},\hat{{\bm{C}}}_{\rm amp}\in\mathbb{R}^{d\times p\times p}$
-Our analysis is based on two remarkable properties that make AMP a particularly useful tool for studying high-dimensional estimation.
-The first property is that for any $t=\Theta(1)$, in the high-dimensional limit $d\to\infty$ the performance of AMP can be tracked without actually running the algorithm.
-This result, known as state evolution, makes AMP mathematically tractable in high-dimensions (Bayati and Montanari, 2011).
-The second property is its optimality with respect to Bayesian estimation.
-We discuss this further in Appendix A.
-For multi-index models, the state evolution equations were derived by Aubin et al.
-(2020b) and rigorously proven by Gerbelot and Berthier (2023).
-The state evolution equations, in particular, provide an exact characterization of the asymptotic overlaps and prediction error.
-Lemma 2.1 (State evolution (Aubin et al., 2019; Gerbelot and Berthier, 2023)).
-Let $({\bm{x}}_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d.
-samples from the multi-index model defined in 1.
-Consider running Algorithm 1 from initial condition $\hat{{{\bm{{W}}}}}^{0}\in\mathbb{R}^{p\times d}$, such that the initial overlap $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ converges in probability to a limit ${{\bm{M}}}^{0}$ as $d\rightarrow\infty$.
-Denote by $\hat{{{\bm{{W}}}}}^{t}$ the resulting estimator at time $0\leq t\leq T$.
-Then, in the high-dimensional limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, constant $p$ and any finite time $T$, the limiting overlaps satisfy:
-[…]
-with ${{\bm{M}}}^{t}$ satisfying the following state evolution equations from initial condition ${{\bm{M}}}^{0}$:
-${{\bm{M}}}^{t+1}\!$ $=\!F({{\bm{M}}}^{t})\!\coloneqq{G}\!\left(\!\alpha\,\mathbb{E}\Big{[}{\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}\!\Big{]}\right).$ (7)
-where $G({\bm{X}})=({\bm{I}}_{p}+{\bm{X}})^{-1}{\bm{X}}$ and the expectation is taken over the following effective process
-$Y^{t}=g\left(({\bm{I}}_{p}-{{\bm{M}}}^{t})^{\nicefrac{{1}}{{2}}}{\bm{\xi}}+{{{\bm{M}}}^{t}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}\right),\qquad{\bm{Z}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ (8)
-with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ independently from ${\bm{Z}}$.
-Furthermore, ${{\bm{M}}}^{t}$ is constrained to lie in $\mathcal{S}^{+}_{p}$ due to Nishimori-identity (section G.3 in (Aubin et al., 2019)).
-As a consequence, the asymptotic mean-squared error on the label prediction is given by:
-${\mathbb{E}_{\bm{X},\bm{y}}}\mathbb{E}_{\bm{x},y}\left[\left(y-g\left(\hat{\bm{W}}^{t}(\bm{X},\bm{y})\bm{x}\right)\right)^{2}\right]=\mathbb{E}[(Y^{t}-g({\bm{Z}}))^{2}],$ $0\leq t\leq T$ (9)
-where the expectation is taken over the effective estimation process Equation 8.
-3 The trivial subspace ($\alpha_{c}=0$)
-The state evolution Equation 7 maps the problem of characterising the computational bottlenecks of first-order methods for high-dimensional Gaussian multi-index models to the study of the deterministic, $p$-dimensional dynamical system ${{\bm{M}}}^{t+1}=F({{\bm{M}}}^{t})$.
-A starting point is identifying its fixed points and their basins of attraction.
-In the absence of any prior information on ${{\bm{{W}}}}^{\star}$ aside from its distribution, one cannot do better than taking $\hat{{\bm{w}}}_{k}^{t=0}\sim\mathcal{N}(0,{\bm{I}}_{d})$ with $k\in[p]$ independently at random from the prior.
-With high-probability, at initialisation, the elements of the overlap matrix $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ are $\Theta(d^{-\nicefrac{{1}}{{2}}})$ element-wise.
-The asymptotic overlap for an uninformed initial condition is thus identically zero ${{\bm{M}}}^{0}=0$, a null-rank matrix.
-If ${{\bm{M}}}^{0}=0$ is not a fixed point, then ${{\bm{M}}}^{1}\succ 0$, implying the weak recovery of a subspace of dimension $k={\rm rank}({{\bm{M}}}^{1})>0$ with just a single step of AMP.
-Lemma 3.1 (Existence of uninformed fixed point).
-${{\bm{M}}}=0\in\mathbb{R}^{p\times p}$ is a fixed point of Equation 7 if and only if the following condition holds almost surely over $Y$:
-${\bm{g}}_{\rm out}(Y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})=\mathbb{E}[{\bm{Z}}|Y]=\mathbf{0},$ (10)
-This implies that as long as the conditional expectation above is not zero almost surely, AMP weakly learns a non-empty subspace immediately in the first iteration for any number of samples $n=\Theta(d)$.
-For this reason, we refer to this subspace as a “trivial subspace”:
-Definition 3 (Trivial subspace).
-We define $H^{\star}_{T}\subseteq\mathbb{R}^{p}$ as the subspace spanned by $\mathbf{v}\in\mathbb{R}^{p}$ such that the following holds almost surely over $Y=g({\bm{Z}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
-${\bm{g}}_{\rm out}(Y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}[\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle|y=Y]=0,$ (11)
-where the expectation is w.r.t the joint measure $p(\mathbf{x},y)$ defined in Definition 1. Equation 8. The trivial subspace $T^{\star}$ is defined as the orthogonal complement of $H^{\star}$.
-Note that for single-index models ($p=1$), the condition in Equation 10 reduces to the one derived by (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020).
-Interestingly, this is exactly the same condition appearing in Damian et al.
-(2024) for weak learnability of single-index models in the SQ model; see eq.
-(3) therein.
-To make Lemma 3.1 concrete, let us look at a few examples.
-A detailed derivation of these examples is discussed in the Appendix C.
-  1. 1.
-    For single-index models ($p=1$), $T^{*}$ is one dimensional if and only if $g$ is non-even, e.g.
-$g(z)={\rm He}_{3}(z)$.
-This follows from requiring that $g_{\rm out}(y,0,1)\neq 0$ for at least one value of $y$.
-In particular, on any open interval where $g_{\rm out}$ is invertible we have $g_{\rm out}=g^{-1}$.
-  2. 2.
-    For a linear multi-index model, $g({\bm{z}})=\sum_{i=1}^{p}z_{i}$, $T^{\star}$ is spanned by $\mathbf{1}_{p}\in\mathbb{R}^{p}$ (all-one vector).
-  3. 3.
-    For a committee $g({\bm{z}})\!=\!\sum_{i=1}^{p}{\rm sign}(z_{i})$, the trivial subspace $T^{\star}$ is again 1-d, spanned by $\mathbf{1}\!\in\!\mathbb{R}^{p}$.
-  4. 4.
-    For monomials $g({\bm{z}})=z_{1}\dots z_{p}$, the trivial subspace $T^{\star}$ is non-empty if and only if $p=1$.
-  5. 5.
-    For staircase functions (Abbe et al., 2023):
-    $g({\bm{z}})=z_{1}+z_{1}z_{2}+z_{1}z_{2}z_{3}+\dots+z_{1}\dots z_{p}$ (12)
-    The trivial subspace is $T^{\star}=\mathbb{R}^{p}$ and is spanned by the canonical basis. In other words, AMP learns all the directions with a single step for any $\alpha>0$.
-Interestingly, Lemma 3.1 can also be related to computational models based on queries, such as SQ learning (Kearns, 1998): the denoiser ${\bm{g}}_{\rm out}$ can indeed be interpreted as a non-linear transformation on the labels $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$.
-From this perspective, the statement on the condition for the existence of a non-empty trivial subspace Equation 11 translates to the following condition:
-$\mathbb{E}\left[{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=\mathbb{E}\left[\mathbb{E}[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|Y=y]^{2}\right]\neq 0,$ (13)
-where ${\bm{v}}\in T^{\star}$.
-Indeed, the left-hand side can be interpreted as a statistical query of the type $\mathbb{E}[\varphi(y)\psi(x)]$ with label pre-processing $\varphi={\bm{g}}_{\rm out}$.
-The fact that this linear correlation in the transformed labels is non-vanishing implies that one can weakly recover $\mathbf{v}$ through a tailored spectral method (Mondelli and Montanari, 2019).
-In fact the denoiser ${\bm{g}}_{\rm out}$ is the optimal such transformation in the sense that when ${\bm{g}}_{\rm out}$ fails to obtain a linear correlation along $\mathbf{v}$, i.e when $\mathbf{v}\in H^{\star}$, then no transformation can:
-Lemma 3.2.
-For any $\mathbf{v}\in H^{\star}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
-$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=0$ (14)
-Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem).
-Proof.
-The above is a direct consequence of the Tower law of expectation. Specifically, we have:
-$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right].$
-The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$.
-∎
-This implies that when $T^{\star}$ is empty, running one-pass SGD after applying any transformation $\mathcal{T}$ fails to obtain weak recovery with $\mathcal{O}(d)$ samples, as the absence of a first-order Hermite coefficient implies a sample complexity of $\mathcal{O}(d\log d)$ as shown in Theorems 1.3 and 1.4 in Arous et al.
-(2021).
-The situation can be very different when reusing batches: Dandi et al.
-(2024) showed that full batch GD implicitly applies a transformation $\mathcal{T}$ to the labels, in a mechanism similar to the one discussed above.
-However, this transformation may not be optimal.
-For instance, while AMP learns the trivial subspaces of both $g(z)={\rm He}_{3}(z)$ and staircase functions (12) in a single-step, GD requires 2-steps for the first and $p$ steps for the latter (Dandi et al., 2024).
-4 Computational phase transitions ($\alpha_{c}>0$)
-What happens when the trivial subspace $T^{\star}$ from Definition 3 is empty?
-In this section, we discuss precisely this case, showing that for some link functions $g$ there exists a critical sample complexity threshold $\alpha_{c}>0$ above which some directions become learnable by iterating AMP.
-To contrast with the trivial subspace, we refer to these as the easy directions.
-While the trivial subspace is characterised by the existence of the fixed point ${{\bm{M}}}=0$, the easy subspace will be determined by its stability, which crucially depends on the sample complexity $\alpha=n/d$.
-The stability of ${{\bm{M}}}=\mathbf{0}$ can be obtained by linearising the state evolution (7) around $\delta{{\bm{M}}}$ with $\norm{\delta{{\bm{M}}}}\approx 0$
-$F({{\bm{M}}})\approx\alpha\mathcal{F}(\delta{{\bm{M}}})+\mathcal{O}(\norm{\delta{{\bm{M}}}}^{2})$ (15)
-where $\mathcal{F}(\delta{{\bm{M}}})$ is a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$:
-$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\mathbb{E}_{Y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,0,{\bm{I}}_{p})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,0,{\bm{I}}_{p})^{\top}\right]\,,$ (16)
-This can be summarized as follows:
-Lemma 4.1.
-(Stability of the uninformed fixed point).
-If ${{\bm{M}}}=\mathbf{0}\in\mathbb{R}^{p\times p}$ is a fixed point, then it is an unstable fixed point of Equation 7 if and only if $n>\alpha_{c}d$, where the critical sample complexity $\alpha_{c}$ is:
-$\frac{1}{\alpha_{c}}=\sup_{{{\bm{M}}}\in\mathcal{S}^{+}_{p}}\norm{\mathcal{F}({{\bm{M}}})}_{F},$ (17)
-with $\|\cdot\|_{F}$ the Frobenius norm, and the extremum being achieved at a unique extremizer ${{\bm{M}}}^{\star}\in\mathcal{S}^{+}_{p}$.
-Moreover, if $\norm{\mathcal{F}({{\bm{M}}})}_{F}=0$, then ${{\bm{M}}}=\mathbf{0}$ is a stable fixed point for any $n=\Theta(d)$.
-The above result follows from a generalization of the Perron-Frobenius theorem to the cone $\mathcal{S}^{+}_{p}$; see Section B.2 for a proof.
-A consequence of this result is that for $\alpha<\alpha_{c}$, even with infinitesimally small informed initialisation AMP will converge back to the trivial fixed point and fail to recover any subspace, while for $\alpha>\alpha_{c}$, it develops a growing overlap along a non-empty subspace starting with arbitrarily small initialization.
-This marks the onset of a computational phase transition for weak learnability in the first-order methods class.
-Note that the computational weak learnability threshold $\alpha_{c}$ can be equal to the information-theoretical one (as, for instance, in ”real” phase retrieval (Mondelli and Montanari, 2019; Maillard et al., 2020)), and by no means we imply the existence (or lack of thereof) of a computational-to-statistical gap (Zdeborová and Krzakala, 2016).
-The easy subspace spanned by the directions that can be learned with $n>\alpha_{c}d$ samples can be characterised as follows:
-Definition 4 (Easy subspace $E^{\star}$).
-Let $H^{*}_{E}$ be the subspace spanned by directions $\mathbf{v}\in\mathbb{R}^{p}$ such that:
-$\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})\mathbf{v}=0,$ (18)
-almost surely over $Y$. We define the easy subspace $E^{\star}$ as the orthogonal complement of $H^{\star}_{E}$.
-Lemma 4.1 implies that for $\alpha>\alpha_{c}$ AMP Algorithm 1 initialised from small but non-zero ${{\bm{M}}}^{0}$ will learn some directions in the easy subspace $E^{\star}$.
-Surprisingly, we can show that for $\alpha>\alpha_{c}$ AMP Algorithm 1 spans the full easy subspace $E^{\star}$ starting from an arbitrarily small but extensive initial correlation.
-Theorem 4.2.
-Let ${{\bm{M}}}_{d}^{t}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ denote the model-target overlap matrix at any finite time $t$.
-Suppose that $T^{\star}={0}$ and the AMP algorithm 1 is initialised at $\hat{{{\bm{{W}}}}}^{0}$ such that ${{\bm{M}}}_{d}^{0}=\epsilon{\bm{A}}$, with $\norm{{\bm{A}}}_{F}=1$ and ${\bm{A}}$ being full-rank.
-Then,under Assumption 1, with high probability as $d\rightarrow\infty$, for $\alpha\geq\alpha_{c}$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}^{\star}$ for $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{\star}$ is as defined in Lemma 4.1.
-For $\alpha<\alpha_{c}$ however, ${{\bm{M}}}^{t}_{d}=0$ is asympotically stable i.e.
-there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{d}}\rightarrow 0$ as $t\rightarrow\infty$.
-Furthermore, there exists an $\alpha>0$ and a $\delta>0$ such that ${{\bm{M}}}_{d}^{t}\succ\delta M_{E^{*}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$ iterations, where $M_{E^{\star}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}$.
-Note that in the above the initial overlap scale $\epsilon>0$ can be arbitrarily small but constant with respect to $d$.
-Heuristically, $\epsilon$ is often identified with $\mathcal{O}(\nicefrac{{1}}{{\sqrt{d}}})$, the scale of a random initial condition.
-Although this heuristic is not covered by Theorem 4.2, we have numerically checked it is consistent with the extrapolation of the theory, which suggests that for $\alpha>\alpha_{c}$ easy subspaces can be weakly recovered with $\mathcal{O}(\log d)$ steps, see Figure 2.
-As for Lemma 3.1, the expression for the weak recovery threshold Equation 17 generalises the single-index expression from (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020).
-Indeed, for $p=1$, we have $\partial_{\omega}g_{\rm out}(y,0,1)=\mathbb{E}[{\rm He}_{2}(Z)|Y=y]$, and therefore
-[…]
-which is exactly eq.
-(11) in Barbier et al.
-(2019).
-Optimality of AMP then implies that no first-order method can achieve non-vanishing correlation with ${\bm{w}}^{\star}$ when $n<\alpha_{c}d$ as $d\to\infty$.
-Figure 1: Numerical illustration of the weak learnability phase transition for the 2-sparse parity $g(z_{1},z_{2})\!=\!{\rm sign}(z_{1}z_{2})$ that has a phase transition at $\alpha_{c}(2)=\pi^{2}/4$.
-The overlap shows how well the directions $z_{1}$ and $z_{2}$ are recovered.
-Given the permutation symmetry in (20), we show here and in all the subsequent figures the optimal permutation of the overlap matrix elements reached by AMP.
-The solid black line is the prediction from the theory.
-Crosses are averages over $72$ runs of AMP Algorithm 1 with $d\!=\!500$.
-We now illustrate Lemma 4.1 in a few examples of interest.
-  1. 1.
-    The monomial $g({\bm{z}})=z_{1}\dots z_{p}$ with $p>1$ can always be learned with $\alpha>\alpha_{c}(p)$ large enough (Chen and Meka, 2020).
-For instance, we have $\alpha_{c}(2)\approx 0.5937$, $\alpha_{c}(3)\approx 3.725$ and $\alpha_{c}(4)\approx 4.912$.
-In Appendix C we derive an analytical formula for $\alpha_{c}(p)$ for arbitrary $p$, and show that $\alpha_{c}(p)\sim p^{1.2}$ for large $p$.
-  2. 2.
-    The embedding of the sparse parity functions:
-    $g({\bm{z}})=\prod_{k=1}^{p}{\rm sign}(z_{k})\,.$ (20)
-    As this is invariant under permutations of the indices, Lemma 4.1 implies the existence of a computational phase transition.
-In Section C.3 we compute analytically the critical value $\alpha_{c}(p)$.
-For $p=1$ this problem is equivalent to phase retrieval, for which we have $\alpha_{c}(1)=\nicefrac{{1}}{{2}}$.
-For $p=2$ we show in the appendix that $\alpha_{c}(2)=\nicefrac{{\pi^{2}}}{{4}}$, while $\alpha_{c}(p)=+\infty$ for $p\geq 3$.
-Figure 1 illustrates this discussion for the 2-sparse parity problem, comparing theoretical prediction Equation 7 with finite-size runs of Algorithm 1.
-Figure 2: Hierarchical weak learnability for the staircase function $g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$.
-(Left): Overlaps with the first direction $|M_{11}|$ ( blue), and with the second and third one $\nicefrac{{1}}{{2}}(M_{22}+M_{33})$ ( red) as a function of the sample complexity $\alpha=\nicefrac{{n}}{{d}}$, with solid lines denoting state evolution curves Equation 7, and crosses/dots finite-size runs of AMP Algorithm 1 with $d=500$ and averaged over $72$ seeds.
-All other overlaps are zero (black).
-The two black dots indicate the critical thresholds at $\alpha_{1}\approx 0.575$ and $\alpha_{2}=\pi^{2}/4$.
-(Right) Corresponding generalization error as a function of the sample complexity.
-Details on the numerical implementation are discussed in Appendix D.
-The sparse-parity function is a classic example of a computationally hard function in theoretical computer science (Blum et al., 1994; 2003).
-[^4] Strictly speaking, the noiseless parity is efficiently learnable by Gaussian elimination (Blum et al., 1994).
-Here, we consider computational efficiency in a noise-tolerant sense.
-Therefore, it should come as no surprise that $\alpha_{c}(p)=\infty$ for $p\geq 3$, providing a concrete example of what a hard subspace (i.e.
-that cannot be learned with $n=\Theta(d)$ samples) looks like for AMP.
-Similar to the discussion of trivial directions in Lemma 3.2, we can invoke optimality of AMP to translate our result on the optimal denoiser $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$ to a statement for queries on general label pre-processing transformations:
-Lemma 4.3.
-For any $\mathbf{v}\in H^{\star}_{E}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
-$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0$ (21)
-Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem from Equation 8).
-For an AMP-hard function (that is, when $\alpha_{c}=\infty$), the above Lemma implies that, even after applying any transformation to the output (as allowed in SQ) online-SGD would require at least $\Omega(d^{2})$ data, since online-SGD with squared loss requires $\Omega(d^{\kappa-1})$ samples/iterations for weak-recovery of a target function when the first $\kappa$ Hermite-coefficients are 00 (Theorems 1.3 and 1.4 in Arous et al.
-(2021)).
-Furthermore, for single-index models satisfying the condition in Lemma 4.3 (i.e $\alpha_{c}=\infty$), the SQ lower-bound predicts a sample complexity requirement of at-least $\Omega(d^{3/2})$ (Damian et al., 2024).
-Therefore, learning such functions with $\mathcal{O}(d)$ samples is conjectured to be algorithmically-hard.
-5 Hierarchical Iterative Denoising
-Figure 3: Trajectories of a single finite-size run of AMP with $d=500$ at $\alpha=4$ for $g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$.
-(Left) Evolution of the overlaps.
-We display $M_{11}$ in blue, $\nicefrac{{1}}{{2}}(M_{22}+M_{33})$ in red, and the off-diagonal overlaps in black.
-(Right) Evolution of the generalisation error.
-In Section 3 and 4 we have provided a full classification of what class of multi-index models are trivial, easy or hard to learn for a first-order algorithm.
-This classification was built upon the notion of weak learnability, or the capacity of AMP to learn subspaces starting from a random initial condition.
-However, it says nothing about about what can be learned after a given subspace is learned, for instance, from possible side information about a particular subspace.
-We now address this question.
-Suppose that the estimator $\hat{{{\bm{{W}}}}}^{t}$ has developed an overlap along a subspace belonging to the span of ${{\bm{{W}}}}^{\star}$, resulting in a non-zero overlap ${{\bm{M}}}^{t}\succ 0$.
-From that point, the main difference with respect to the previous discussion is that the variable ${\bm{\omega}}$ in the linear operator $\mathcal{F}$ defined in Equation 16 becomes non-zero (since it is distributed as ${\bm{\omega}}=\sqrt{{{\bm{M}}}}{\bm{\xi}}$).
-Crucially, this changes the span of $\mathcal{F}({{\bm{M}}})$ and hence the stability condition in Lemma 4.1.
-In particular, as we show next, learning some directions might facilitate learning larger subspaces.
-This phenomenon is reminiscent of the specialisation transition in committee machines (Saad and Solla, 1996; Aubin et al., 2020b) and of the staircase phenomenon for one-pass SGD introduced in (Abbe et al., 2022).
-Below, we formalise this discussion:
-Definition 5.
-Let $U\in\mathbb{R}^{p}$ be a subspace of dimensions $k$. We define $H_{E}^{\star}(U)$ to be the subspace spanned by ${\bm{v}}\in U^{\perp}$ such that:
-${\bm{v}}^{\top}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}}_{U}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U})=0,$ (22)
-almost surely over ${\bm{\xi}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
-We define the “trivially-denoising-coupled” subspace $T^{\star}_{U}$ for $U$ as the orthogonal complement of $H_{E}^{\star}(U)$.
-Analogously, let $H^{\star}_{E}(U)$ be the subspace spanned by directions ${\bm{v}}\in U^{\perp}$ such that:
-${\bm{v}}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U}){\bm{v}}=0,$ (23)
-almost surely over ${\bm{\xi}}$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
-When ${{\bm{M}}}_{U}$ is additionally a fixed point of $\mathcal{F}_{{{\bm{M}}}}$, we can linearise $\mathcal{F}_{{{\bm{M}}}}$ along the orthogonal complement of $U$.
-We define the easy-denoising-coupled subspace $E^{\star}_{U}$ for $U$ as the orthogonal complement of $H^{\star}_{E}(U)$.
-Next, suppose that ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ with $\operatorname{span}({{\bm{M}}}_{U})=U$ is further a fixed-point of $\mathcal{F}_{{{\bm{M}}}}$.
-Let $\mathcal{F}_{{{\bm{M}}}_{U}}$ denote the linearization of $F({{\bm{M}}})$ along the orthogonal complement $U^{\perp}$ at ${{\bm{M}}}={{\bm{M}}}_{U}$.
-We define the iterative denoising critical threshold $\alpha_{\text{hid}}({{\bm{M}}}_{U})$ at ${{\bm{M}}}={{\bm{M}}}_{U}$ as:
-$\frac{1}{\alpha_{\text{hid}}({{\bm{M}}}_{U})}=\sup_{{{\bm{M}}}^{\perp}\in U^{\perp}}\norm{\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}})}_{F},$ (24)
-We denote by ${{\bm{M}}}^{\star}_{U}$ the extremiser achieving the above supremum.
-Figure 4: Specialisation as a staircase with $g\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$.
-The direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is first learned for any $\alpha>0$.
-The remaining ones are only learned at the specialisation transition occurring at $\alpha\!\approx\!4.3$.
-We take care of the symmetries as indicated in Appendix D.
-Crosses denote AMP runs with $d\!=\!100$ averaged over $72$ seeds.
-The overlap is shown as a function of the sample complexity.
-Because of the symmetry the elements of $M$ can take one of two values: one on the diagonal and one outside of it.
-We display $(M_{11}+M_{22}+M_{33})/3$ in blue and $(M_{12}+M_{13}+M_{23})/3$ in red.
-The definitions above generalise the notions of trivial and easy subspaces in Definitions 3 and 4 conditionally on a subspace $U$ that has been previously learned, and characterise the directions whose recovery is enabled upon learning the subspace $U$.
-Concretely, upon developing an initial overlap along $U$, the directions in $T^{\star}_{U},E^{\star}_{U}$ can be recovered analogous to the recovery of $T^{\star},E^{\star}$ starting from random initialization.
-We formalise this below:
-Theorem 5.1.
-Let $U\in\mathbb{R}^{p}$ be a fixed subspace.
-Suppose the AMP algorithm 1 is initialised such that ${{\bm{M}}}_{d}^{0}={{\bm{M}}}_{U}+\epsilon{\bm{A}}$, where $\norm{{\bm{A}}}_{F}=1$ and ${{\bm{M}}}_{U}$ is a fixed point of $F({{\bm{M}}})$ in Equation 7 with $\operatorname{span}({{\bm{M}}}_{U})=U$.
-Then, under Assumption 1, the following holds with high probability as $d\rightarrow\infty$: If $\alpha\geq\alpha_{\text{hid}}({{\bm{M}}}_{U})$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}^{(t)}\succ\delta{{\bm{M}}}^{*}_{U}$ for some $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{*}_{U}$ is defined as in Definition 5.
-For $\alpha<\alpha_{\text{hid}}({{\bm{M}}}_{U})$ however, there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{\perp}}=0$ as $t\rightarrow\infty$, where ${{\bm{M}}}^{t}_{\perp}$ denotes the projection of ${{\bm{M}}}^{t}$ orthogonal to $U$.
-Furthermore, suppose that ${\bm{A}}$ is full-rank, then there exists an $\alpha>0$ and a $\delta>0$ such that is ${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}_{E^{\star}_{U}}$, where ${{\bm{M}}}_{E^{\star}_{U}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}_{U}$
-When $E^{\star}_{U}$ is non-empty for some subspace $U\subseteq\mathbb{R}^{p}$, we say that the target $y=g({{\bm{{W}}}}{\bm{x}})$ allows learning through Hierarchical Iterative Denoising.
-A concrete example of a function which displays the hierarchical iterative denoising staircase is a linear combination between hard parity function and an easy polynomial:
-$g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ (25)
-The sign part is a sparse parity with $p=3$, which cannot be learned with $n=\Theta(d)$ samples, but the quadratic polynomial $z_{1}^{2}$ component in the function allows weak-recovery of the first component i.e.
-$U={(1,0,0)}$ as long as $\alpha>1/2$.
-Hence, conditionally on $U$ the effective multi-index model becomes ${\rm sign}(z_{2}z_{3})$, which as discussed in Section 4 is an easy function.
-Figure 2 illustrates this: first, $z_{1}$ is learned at $\alpha_{1}\approx 0.575$.
-Then, for larger value when $\alpha>\alpha_{2}$, all directions are learned (see App.
-(C.4)).
-Knowing $z_{1}$ makes the hard 3-parity an easy 2-parity problem.
-We can see the evolution of the overlaps over time in Figure 3.
-It is easy to construct a multi-index model where AMP will iterate over any number of such plateaus. For instance, consider the model:
-$g({\bm{z}})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})+{\rm sign}(z_{3}z_{4}z_{5})+{\rm sign}(z_{5}z_{6}z_{7})+\ldots$ (26)
-After the first plateau to learn $z_{1}$, there will be one for $z_{2}$, then for $z_{3}$, for $z_{4}$ and so on.
-Another example is provided in Fig.
-4 for the neural net $g(z_{1},z_{2},z_{3})\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$.
-At first only the one-dimensional direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is learned for any $\alpha>0$.
-This is simply a linear approximation of the function.
-Only at $\alpha\!\approx\!4.3$ does the network learns the other directions.
-Known as the the specialisation transition, this has been first discussed by Saad and Solla (1996); Aubin et al.
-(2020a), and provide a very early example of iterative sequential learning.
-Finally, it is important to stress that hierarchical learnable functions are not equivalent to staircase functions (Abbe et al., 2022; 2023). For instance, consider the target:
-$g({\bm{z}})=\rm He_{4}(z_{1})+{\rm sign}(z_{1}z_{2}z_{3}).$ (27)
-Similar to Example 26, the above target is learned sequentially with $\mathcal{O}(d)$ samples by first recovering the direction corresponding to $z_{1}$.
-However, under the CSQ staircase classification, $g({\bm{z}})$ corresponds to leap $3$.
-The CSQ lower-bound therefore implies a sample complexity-requirement of $\Omega(d^{1.5})$ with online-SGD learning it in $\mathcal{O}(d^{2})$ samples (Abbe et al., 2023).
-As we argued in Lemmas 3.2 and 4.3, we believe Hierarchical Iterative Denoising provides the right notion of computational complexity for these models, since it encompasses a larger class of natural algorithms such as batch gradient descent.
-6 Conclusion and limitations
-This work investigated the computational boundaries of learning high-dimensional Gaussian multi-index models with first-order iterative algorithms.
-As we hinted in the text (and provided some partial answers), a natural question is how close neural networks are to achieving this optimal performance.
-Two limitations of our work are: (a) our results hold asymptotically in $d\to\infty$; (b) We assume Gaussian covariates.
-For the former, numerical experiments support that the predictions are already accurate at fairly moderate sizes.
-For the latter, a number of universality results justify the scope of high-dimensional Gaussian covariates and extend AMP to more structured designs (Loureiro et al., 2021; Wang et al., 2022; Dudeja et al., 2023; Dandi et al., 2023).
-Acknowledgements
-The authors would like to thank Joan Bruna, Theodor Misiakiewicz, Luca Pesce, and Nati Srebro for stimulating discussions.
-This work was supported by the Swiss National Science Foundation under grants SNSF SMArtNet (grant number 212049) and SNSF OperaGOST (grant number 200390) and the Choose France - CNRS AI Rising Talents program.
-Appendix / supplemental material
-Appendix A Optimality of AMP
-Recall that AMP Algorithm 1 is tailored to estimate the posterior marginals Equation 5.
-But how efficient is it?
-A priori, making this comparison requires sampling from the posterior distribution, which is computationally prohibitive in the high-dimensional limit of interest here.
+Sampling is, a priori, computationally prohibitive in the high-dimensional limit of interest here.
+The first impressive feat of AMP, a simple iterative first-order algorithm, is thus its efficiency.
 The second remarkable property of AMP is that its optimality with respect to the Bayesian posterior Equation 7 can be exactly characterised in the high-dimensional regime.
-Lemma A.1 (Bayes-optimal correlation).
+Indeed, Aubin et al.
+[2019] —generalizing the earlier results of Barbier et al.
+[2019] on single index models— proved the following (formally, their proof was for the committee machine, but it applies mutatis mutandis to any multi-index models):
+Theorem A.1 (Bayes-optimal correlation, Theorem 3.1 in Aubin et al. [2019], informal).
 Let $(x_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d.
@@
 Indeed, remarking that the update function $F$ in Equation 9 is precisely the gradient of the entropy $H_{Y}$ in Equation 10, one can show that state evolution is equivalent to gradient descent in the objective defined by Equation 29 [Aubin et al., 2019].
-This non-trivial fact implies that whenever Equation 29 has a single minima, AMP optimally estimates the posterior marginals.
+This non-trivial fact implies that whenever Equation 29 has a single minima, AMP optimally estimates the posterior marginals, at least if its state evolution converges to this fixed point.
 Finally, note that by construction $\hat{{{\bm{{W}}}}}_{\rm bo}$ is the optimal estimator of ${{\bm{{W}}}}^{\star}$ given the data $(X,y)$ (in the MMSE sense).
@@
 The fact that AMP follows state evolution is proven for such problems in [Javanmard and Montanari, 2013, Gerbelot and Berthier, 2023].
-Appendix B Proofs of the main results
-B.1 Linear Approximation
+Specifically, we refer to the example 4.4 in Gerbelot and Berthier [2023] on Matrix-valued random variables.
+Appendix B Bayes-AMP in the presence of side-information
+The algorithm is easily generalized to the case with side information of the type eq.(18):
+${\bf S}=\sqrt{\lambda}W^{\star}+\sqrt{1-\lambda}Z,$ (30)
+The effect of the additional side-information is equivalent to a modification of the prior on $W$ to $P^{S}_{W}(W)$, defined as:
+$P^{S}_{W}({{\bm{{W}}}})$ $\propto P({\bm{S}}|{{\bm{{W}}}})P_{W}({{\bm{{W}}}})={\cal N}(\sqrt{\lambda}{\bf S},\sqrt{1-\lambda}\mathbf{I})$
+This is a simple consequence of Bayes theorem:
+$P({{\bm{{W}}}}|{\bm{y}},{\bm{X}},{\bm{S}})\propto P({\bm{y}},{\bm{X}},{\bm{S}}|{{\bm{{W}}}})P_{W}({{\bm{{W}}}})=P({\bm{y}},{\bm{X}}|{{\bm{{W}}}})P({\bm{S}}|{{\bm{{W}}}})P_{W}({{\bm{{W}}}})$ (31)
+In turns, this change updates of $\hat{{{\bm{{W}}}}}$ and $\hat{{\bm{C}}}$ in Algorithm 1 as described in Aubin et al. [2019]:
+$\hat{{{\bm{{W}}}}}_{k}^{t+1}=({\bm{I}}_{p}(1-\lambda)+{\bm{A}}_{k}^{t})^{-1}(\bm{b}_{k}^{t}(1-\lambda)+\sqrt{\lambda}{\bm{A}}_{k}^{t}{\bm{S}})\,,\qquad\hat{{\bm{C}}}_{k}^{t+1}=({\bm{I}}_{p}(1-\lambda)+{\bm{A}}_{k}^{t})^{-1}\,.$ (32)
+It is easily checked that the limit $\lambda\to 0$ gives back the equations in Algorithm 1. Lemma 2.1 is then modified with $G(X)$ in equation (9) being replaced by
+$G(X)=(X(1-\lambda)+\lambda)(1+X(1-\lambda))^{-1}\,.$ (33)
+Appendix C Trivially-coupled Subspaces
+In this section, we discuss and illustrate examples of $g$ where $T^{\star}_{U}\neq\emptyset$ for a learned subspace $U$. Our prototypical example throughout the discussion will be:
+$g(z_{1},z_{2},z_{3})=z^{2}_{1}+z^{2}_{2}+\text{sign}(z_{1}z_{2}z_{3}).$ (34)
+Let $e_{1},e_{2},e_{3}$ denote the directions corresponding to components $z_{1},z_{2},z_{3}$ respectively.
+It is easy to check that for the above example, $E^{\star}=\{e_{1},e_{2}\}$ while $H^{\star}_{E}=\{e_{3}\}$.
+Therefore, by Theorem 4.2, for large enough $\alpha$ the ${{\bm{M}}}^{t}\geq\delta M_{E}^{\star}$ where $M_{E}^{\star}$ spans $z_{1},z_{2}$ after $t=\mathcal{O}(\log\frac{1}{\lambda})$ iterations.
+Subsequently, since $T^{\star}_{E^{\star}}=\{e_{3}\}$, it gains an overlap along $e_{3}$ (independent of $\lambda$).
+Unlike sequential learning of directions in $E^{\star}_{U}$, however, the time-steps for recovery of $\{e_{1},e_{2}\}$ and $\{e_{3}\}$ (upto a threshold independent of $\lambda$) different only by a constant.
+The recovery of $e_{3}$ can therefore be interpreted as instantaneous w.r.t $\{e_{1},e_{2}\}$.
+Appendix D Proofs of the main results
+In this section, we describe the proofs of our main results, starting with a Trivial proof of the learning of Trivial subspaces (Theorem 3.2) and the form of linearization that will be utilized throughout the analysis.
+D.1 Theorem 3.2 and Linearization
+Lemma 3.1 and Theorem 3.2 are direct consequences of the following observation:
+Proposition D.1.
+Let $\mathbf{v}\in\mathbb{R}^{p}$ be arbitrary. Then, starting from ${{\bm{M}}}^{0}=0$, $\mathbf{v}^{\top}{{\bm{M}}}^{1}\mathbf{v}>0$ if and only if:
+$\alpha\mathbb{E}\left[\mathbf{v}^{\top}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right){\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)^{\top}\mathbf{v}\right]>0$ (35)
+Proof.
+The above proposition follows directly from the form of $F$ in Lemma 2.1 and the observation that the mapping $G(X)$ preserves the span of $X$. ∎
+The expression in Equation 35 can be rewritten as:
+$\mathbb{E}\left[(\mathbf{v}^{\top}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right))^{2}\right]$ (36)
+Since the operand inside the expectation is non-negative, the expectation is non-zero if and only if the operand vanishes identically. We therefore obtain Lemma 3.1 and Theorem 3.2.
+D.1.1 Linearization without side-information
 Our analysis relies on the following result:
-Lemma B.1.
+Lemma D.2.
 Let $F({{\bm{M}}})$ be as defined in Lemma 2.1
@@
 where $\norm{}_{F}$ denotes the Frobenius norm and $\mathcal{F}(\delta{{\bm{M}}})$ is a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$:
-$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]\,,$ (31)
+$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p})^{\top}\right]\,,$ (38)
 We proceed through an-entry-wise expansion of each term inside the expection in $F({{\bm{M}}})$ around ${{\bm{M}}}=0$.
@@
 $F({{\bm{M}}})$ $=\alpha\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]+\alpha\mathcal{O}(\norm{M}^{2}_{F})$
-B.2 Proof of Lemma 4.1
+D.1.2 Linearization in the presence of side-information
+Analogous to Lemma D.2, we can linearize state-evolution in the presence of side-information in the limit of small $M,\lambda$.
+Lemma D.3.
+Let $\mathcal{F}(\delta{{\bm{M}}})$ be a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$, defined by:
+$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]$ (39)
+Then the following approximation holds:
+$F({{\bm{M}}})\approx\alpha\mathcal{F}(\delta{{\bm{M}}})+\sqrt{\lambda}\mathbf{I}_{d}+\mathcal{O}(\alpha\norm{\delta{{\bm{M}}}}_{F}^{2})+\mathcal{O}(\lambda),$ (40)
+Proof.
+The above linearization is a direct consequence of Lemma D.2 along with the updated form of $G(X)$ in Equation 33. ∎
+D.2 Proof of Lemma 3.3
+Proof.
+This is a direct consequence of the Tower law of expectation: $\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right]$ The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$.
+The converse for any $v\in T^{\star}$ then follows by setting $\mathcal{T}(y)={\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$.∎
+D.3 Proof of Lemma 4.1
 Suppose that $M\in\mathcal{S}^{+}$, we have, for any $\mathbf{v}\in\mathbb{R}^{p}$:
@@
 Thus, $\mathcal{F}(M)$ is a cone-preserving linear map.
-The generalised Perron-Frobenius theorem for cone-preserving maps [Krein and Rutman, 1948] then implies that the operator $\mathcal{F}({{\bm{M}}})$ admits a unique eigenvector $M^{*}\in\mathcal{S}^{+}_{p}$ corresponding to the largest eigenvalue $\lambda_{\mathcal{F}}$ such that for any $M\in\mathcal{S}^{+}_{p}\perp M^{*}$:
-$\mathcal{F}({{\bm{M}}})<\lambda_{\mathcal{F}}\norm{{{\bm{M}}}}_{F}.$ (34)
+The generalised Perron-Frobenius theorem for cone-preserving maps [Krein and Rutman, 1948] then implies that the operator $\mathcal{F}({{\bm{M}}})$ admits a unique eigenvector $M^{*}\in\mathcal{S}^{+}_{p}$ corresponding to the largest eigenvalue $\nu_{\mathcal{F}}$ such that for any $M\in\mathcal{S}^{+}_{p}\perp M^{*}$:
+$\mathcal{F}({{\bm{M}}})<\nu_{\mathcal{F}}\norm{{{\bm{M}}}}_{F}.$ (43)
 Furthermore, all other eigenvalues are strictly smaller than $\lambda_{\mathcal{F}}$.
-Subsequently, Lemma B.1 implies that $F({{\bm{M}}})$ is stable at ${{\bm{M}}}=0$ if and only if $\alpha\leq\frac{1}{\lambda_{\mathcal{F}}}=\alpha_{c}$.
-B.3 Proof of Theorem 4.2
-Lemma 2.1 allows us to map the behavior of the variable ${{\bm{M}}}^{t}$ in state-evolution to high-probability statements for the limiting overlaps Suppose that $A\in\mathcal{S}^{+}_{p}$ is a full rank matrix as in Theorem 4.2.
-Applying Lemmas B.1 and 3.2, we obtain that for ${{\bm{M}}}^{0}=\epsilon A$:
-$\norm{{{\bm{M}}}^{t+1}}_{F}\leq\alpha\lambda_{\mathcal{F}}\norm{{{\bm{M}}}^{t}}_{F}+\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F},$ (35)
-for some constant $C$.
-Now, suppose that $\alpha<\frac{1}{\lambda_{F}}$ and $\epsilon<\frac{1}{C}(1-\alpha\lambda_{\mathcal{F}}-\kappa)$, for some $0<\kappa<1-\alpha\lambda_{\mathcal{F}}$ then inductively, we obtain:
-$\norm{{{\bm{M}}}^{t+1}}_{F}<(1-\kappa)\norm{{{\bm{M}}}^{t}}_{F},$ (36)
-for all $t>0$. Implying that $\norm{{{\bm{M}}}^{t}}_{F}\rightarrow 0$.
-Now, suppose that $\alpha>\alpha_{c}$ or equivalently $\alpha\lambda_{F}>1$. Since $A$ is full-rank admits a decomposition:
-$\operatorname{tr}(A,{{\bm{M}}}^{\star})=\theta$ (37)
-for some $\theta>0$. From Lemma 4.1 and B.1, we obtain:
-$\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq\alpha\lambda_{\mathcal{F}}\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (38)
-For $\norm{{{\bm{M}}}^{t}}\leq\frac{\theta}{C}(\alpha\lambda_{\mathcal{F}}-1-\kappa)$ for some $\kappa>0$, we inductively obtain:
+Subsequently, Lemma D.2 implies that $F({{\bm{M}}})$ is stable at ${{\bm{M}}}=0$ if and only if $\alpha\leq\frac{1}{\nu_{\mathcal{F}}}=\alpha_{c}$.
+D.4 Proof of Theorem 4.2
+Lemma 2.1 allows us to map the behavior of the variable ${{\bm{M}}}^{t}$ in state-evolution to high-probability statements for the limiting overlaps.
+Applying Lemmas D.3 and 3.3, we obtain that for small enough ${{\bm{M}}}^{t+1},\lambda$:
+$\norm{{{\bm{M}}}^{t+1}}_{F}\leq\alpha\nu_{\mathcal{F}}\norm{{{\bm{M}}}^{t}}_{F}+\lambda+C_{1}\norm{{{\bm{M}}}^{t}}^{2}_{F}+C_{2}\norm{{{\bm{M}}}^{t}}^{2}_{F}$ (44)
+for some constants $C_{1},C_{2}$. Now, suppose that $\alpha<\frac{1}{\nu_{F}}$. Then for $\norm{{{\bm{M}}}^{t}}_{F}\geq\frac{1}{(1-\alpha\nu_{\mathcal{F}})}\lambda$, we have:
+$\alpha\nu_{\mathcal{F}}\norm{{{\bm{M}}}^{t}}_{F}+\lambda\leq\norm{{{\bm{M}}}^{t}}_{F}.$ (45)
+For small enough $\lambda$, we have that:
+${{\bm{M}}}^{0}\approx\lambda<\frac{1}{(1-\alpha\nu_{\mathcal{F}})}\lambda.$ (46)
+Subsequently, by induction, we obtain that:
+$\sup_{t\geq p}\norm{{{\bm{M}}}^{t+1}}_{F}\leq\frac{1}{(1-\alpha\nu_{\mathcal{F}})}\lambda$ (47)
+for all $t>0$. This proves the first part of the Theorem.
+Now, suppose that $\alpha>\alpha_{c}$ or equivalently $\alpha\nu_{F}>1$.
+From Lemma 4.1 and D.3, we obtain:
+$\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq\alpha\nu_{\mathcal{F}}\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})+\lambda-C_{1}\norm{{{\bm{M}}}^{t}}^{2}_{F}-C_{2}\lambda^{2},$ (48)
+for some constants $C_{1},C_{2}>0$.
+Since $\alpha\nu_{F}>1$, we obtain that $\operatorname{tr}({{\bm{M}}}^{1},{{\bm{M}}}^{\star})>\lambda>0$ for small enough $\lambda$.
+Let $10<\kappa<\nu_{F}-$ 1 be fixed.
+Equation 48 implies that for $\frac{\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})}{\norm{{{\bm{M}}}^{t}}^{2}_{F}}<\frac{\alpha\nu_{F}-1-\kappa}{C_{1}}$, and small enough $\lambda$:
 $\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq(1+\kappa)\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star}),$ (49)
-implying that $\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})$ grows as $\omega(e^{\kappa t})$ for sufficiently small $\epsilon$.
+implying that $\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})$ grows as $\omega(e^{\kappa t})$.
+Since at $t=1$, $\frac{\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})}{\norm{{{\bm{M}}}^{t}}^{2}_{F}}=\mathcal{O}(\frac{1}{\lambda})$, Equation 44 implies that the condition $\frac{\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})}{\norm{{{\bm{M}}}^{t}}^{2}_{F}}<\frac{\alpha\nu_{F}-1-\kappa}{C_{1}}$ holds for time $\mathcal{O}(\log\frac{1}{\lambda})$, which along with Equation 49 allows recovery upto an overlap $\delta$ independent of $\lambda$.
 Finally, it remains to show that for large enough $\alpha$, ${{\bm{M}}}^{t}$ spans $E^{\star}$.
 Notice that for any $\mathbf{v}\in E^{\star}$:
-$\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v=\mathbb{E}_{y}\left[(\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p})\mathbf{v})^{2}\right]>0,$ (40)
-while for any $A\in\mathcal{S}^{+}_{p}$:
-$\mathbf{v}^{\top}\mathcal{F}(A)v\geq 0,$ (41)
-since $\mathcal{F}(A)\in\mathcal{S}^{+}_{p}$. Define:
-$\nu_{\mathcal{F}}=\inf_{\mathbf{v}\in\mathbb{E},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v.$ (42)
-Since $\mathbf{v}$ lies in a compact set, $\nu_{\mathcal{F}}>0$. Let $A^{\star}$ denote the projection of $A$ along $E^{\star}$ and consider its eigendecomposition $A^{\star}$:
-$A^{\star}=\sum_{i=1}^{p}\lambda_{i}\mathbf{v}_{i}\mathbf{v}^{\top}_{i}$ (43)
-We
-$\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t+1})\mathbf{v}_{i}\geq\alpha\nu_{\mathcal{F}}\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t})\mathbf{v}_{i}-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (44)
+$\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})\mathbf{v}=\mathbb{E}_{y}\left[(\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p})\mathbf{v})^{2}\right]>0,$ (50)
+Define the smallest eigenvalue of $\mathcal{F}$:
+$\nu^{\star}_{\mathcal{F}}=\inf_{\mathbf{v}\in E^{\star},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})\mathbf{v}.$ (51)
+Since $\mathbf{v}$ lies in a compact set, $\nu^{\star}_{\mathcal{F}}>0$.
+Let $\mathbf{v}_{1},\cdots,\mathbf{v}_{k}$ denote an orthonormal basis for $E^{\star}$
+$\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t+1})\mathbf{v}_{i}\geq\alpha\nu_{\mathcal{F}}\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t})\mathbf{v}_{i}+\lambda-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (52)
 Therefore, for $\alpha\geq\frac{1}{\nu_{\mathcal{F}}}$ and small enough $\epsilon$, ${{\bm{M}}}^{t}$ expands linearly along each $\mathbf{v}_{i}$.
-B.4 Proof of Lemma 4.3
+D.5 Proof of Lemma 4.3
 From the definition of $g_{out}$ in Equation 6 and the dominated-convergence theorem, it is straightforward to see that:
@@
 $\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0.$ (55)
-B.5 Proof of Theorem 5.1
-By assumption, $M_{U}$ is a fixed point of $F(M)$.
-Therefore, Equation 7 implies that ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}-\sqrt{M}_{U})$ almost surely lies in $U$ and thus $T^{\star}_{U}$ is empty.
-The proof of Theorem 5.1 then follows that of 4.2 by considering the following linearized operator along $U^{\perp}$
-$\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})=\mathrel{\mathop{:}}=\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U}){{\bm{M}}}_{\perp}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U})^{\top}\right]\,,$ (48)
+D.6 Staircase Dynamics and the proof of Proposition 5.1
+Before proving Proposition 5.1, we start by proving a monotonicity property of the AMP iterates, which in-turn implies that the AMP iterates gaining overlap along a certain subspace must eventually reach an approximate fixed point.
+D.6.1 Monotonicity of the AMP Iterates
+We begin by showing that the overlaps of the Bayes-AMP iterates is non-decreasing in the following sense:
+Lemma D.4.
+Consider the AMP algorithm with the Bayes-optimal choice of $f_{t},g_{t}$. Then for any $t>0$:
+${{\bm{M}}}^{(t+1)}\succeq{{\bm{M}}}^{(t)}$ (56)
+Proof.
+Our proof relies on a characterization of the state-evolution for the Bayes-AMP iterates as the solution to certain low-dimensional inference problems.
+See for instance Barbier et al.
+[2019], Aubin et al.
+[2019].
+Concretely, consider the following sequence of denoising problems over variables $Z^{\star}\in\mathbb{R}^{k},W^{\star}\in\mathbb{R}^{k}$:
+$Y_{t}=g(\sqrt{{{\bm{M}}}^{(t)}}\xi+\sqrt{\mathbb{I}_{k}-{{\bm{M}}}^{(t)}}Z^{\star})$
+$\tilde{Y}_{t}=\sqrt{\tilde{M}}_{t}W^{\star}+Z$
+where $Z,\xi\sim\mathcal{N}(0,\mathbf{I}_{k})$ and $\tilde{M}_{t}$ is obtained as:
+$\begin{split}\tilde{{{\bm{M}}}}^{(t)}&\mathrel{\mathop{:}}=\mathbb{E}\left[(\mathbb{E}\left[Z^{\star}|Y,\xi\right]\mathbb{E}\left[Z^{\star}|Y,\xi\right]^{\top})\right]\\ &=\mathbb{E}\left[g_{\text{out}}(Y,\sqrt{{{{\bm{M}}}^{(t)}}}\xi)g_{\text{out}}(Y,\sqrt{{{{\bm{M}}}^{(t)}}}\xi)^{\top}\right]\end{split}$ (57)
+We claim that the following statements hold for any $t>0$:
+$\begin{split}{{\bm{M}}}^{(t+1)}\succeq{{\bm{M}}}^{(t)}\\ \tilde{{{\bm{M}}}}^{(t+1)}\succeq\tilde{{{\bm{M}}}}^{(t)}\end{split}$ (58)
+We proceed by induction.
+At $t=0$, ${{\bm{M}}}^{(t)}=0$ and thus ${{\bm{M}}}^{(t+1)}\succeq{{\bm{M}}}^{(t)}$ holds trivially.
+Subsequently, we show via an information-theoretic argument that ${{\bm{M}}}^{(t+1)}\succeq{{\bm{M}}}^{(t)}$ implies $\tilde{{{\bm{M}}}}^{(t+1)}\succeq\tilde{{{\bm{M}}}}^{(t)}$.
+Suppose that ${{\bm{M}}}^{(t+1)}\succeq{{\bm{M}}}^{(t)}$ holds at time-step $t$.
+We decompose the field entering the channel for $Y_{t+1}$ as follows:
+$Y_{t+1}$ $=g(\sqrt{{{\bm{M}}}^{(t+1)}}\xi+\sqrt{\mathbb{I}_{k}-{{\bm{M}}}^{(t+1)}}Z^{\star})$
+$=g(\sqrt{{{\bm{M}}}^{(t)}}\xi_{1}+\sqrt{{{\bm{M}}}^{(t+1)}-{{\bm{M}}}^{(t)}}\xi_{2}+\sqrt{\mathbb{I}_{k}-{{\bm{M}}}^{(t+1)}}Z^{\star}),$
+where $\xi_{1},\xi_{2}$ denote independent variables $\sim\mathcal{N}(0,1)$. Now, compare the following two estimates:
+$\hat{Z}_{1}=\mathbb{E}\left[Z^{\star}|Y,\xi_{1},\xi_{2}\right],\quad\hat{Z}_{2}=\mathbb{E}\left[Z^{\star}|Y,\xi_{1}\right].$ (59)
+Note that by linearity of expectation and independence of $\xi,Z^{\star}$, the estimator $\hat{Z}_{2}$ equals the estimator of the inference problem at time $t$:
+$Y_{t}=g(\sqrt{{{\bm{M}}}^{(t)}}\xi+\sqrt{\mathbb{I}_{k}-{{\bm{M}}}^{(t)}}Z^{\star})$ (60)
+By the optimality of the posterior-mean w.r.t MSE error, we must have for any $v\in\mathbb{R}^{k}$ with $\norm{v}=1$
+$\mathbb{E}\left[(\langle v,\hat{Z}_{1}\rangle-\langle v,\ Z^{\star}\rangle)^{2}\right]\leq\mathbb{E}\left[(\langle v,\hat{Z}_{2}\rangle-\langle v,\ Z^{\star}\rangle)^{2}\right]$ (61)
+, where $\langle v,\hat{Z}_{i}\rangle$ for $i=1,2$ denotes the projection of $\hat{Z}$ along $v$ or equivalently the Bayes-estimator of $\langle v,\ Z^{\star}\rangle$.
+Now, a straightforward consequence of Bayes-rule (Nishimori’s identity) implies that:
+$\mathbb{E}\left[\langle v,\hat{Z}_{i}\rangle\langle v,\ Z^{\star}\rangle\right]=\mathbb{E}\left[(\langle v,\hat{Z}_{i}\rangle)^{2}\right],$ (62)
+for $i=1,2$. Substituting in Equation 61, we obtain:
+$\mathbb{E}\left[\langle v,\hat{Z}_{1}\rangle\langle v,\ Z^{\star}\rangle\right]\geq\mathbb{E}\left[\langle v,\hat{Z}_{2}\rangle\langle v,\ Z^{\star}\rangle\right].$ (63)
+The above terms can be equivalently expressed as $v^{\top}\tilde{{{\bm{M}}}}^{(t+1)}v,v^{\top}\tilde{{{\bm{M}}}}^{(t)}v$ respectively. We obtain:
+$\tilde{{{\bm{M}}}}^{(t+1)}\succeq\tilde{{{\bm{M}}}}^{(t)}.$ (64)
+Similarly, we have that $\tilde{{{\bm{M}}}}^{(t+1)}\succeq\tilde{{{\bm{M}}}}^{(t)}$ implies ${{\bm{M}}}^{(t+2)}\succeq{{\bm{M}}}^{(t+1)}$. The claim therefore follows by induction.
+D.6.2 Sequential visits to Fixed-points
+Now, suppose that $E^{\star}\neq\phi$ and consider the dynamics with side-information provided only along $E^{\star}$ i.e:
+$S=\sqrt{\lambda}P_{E^{\star}}+\sqrt{1-\lambda}\mathbb{I},+$ (65)
+where $P_{E^{\star}}$ denotes the projection along $E^{\star}$.
+Suppose further tha $\alpha>\alpha_{c}$.
+By Theorem 4.2, ${{\bm{M}}}^{(t)}\succ\delta M_{E^{\star}}$ in $\mathcal{O}(\log\frac{1}{\lambda})$ iterations.
+Since $G(X)\prec\mathbf{I}_{k}$ for any $X\in\mathcal{S}^{+}$, we have:
+${{\bm{M}}}^{(t)}\prec\mathbf{I}_{k},$ (66)
+for all $t>0$.
+Considering the topology induced by the Frobenius norm on $\mathbb{R}^{p\times p}$, we obtain that ${{\bm{M}}}^{(t)}$ is a bounded sequence and therefore contains a convergent subsequence.
+By Lemma D.4, we obtain the the convergence of the entire sequence to a fixed point ${{\bm{M}}}^{f}_{E^{\star}}(\lambda,\alpha)$.
+Therefore, for any $\epsilon>0$, $\exists t^{\prime}$ such that for all $t>t^{\prime}(\epsilon,\lambda)$:
+$\norm{{{\bm{M}}}^{(t)}-{{\bm{M}}}^{f}_{E^{\star}}(\lambda,\alpha)}\leq\epsilon.$ (67)
+Next, Lemma D.4 further implies that:
+${{\bm{M}}}^{f}_{E^{\star}}(\lambda,\alpha)\prec{{\bm{M}}}^{f}_{E^{\star}}(\lambda^{\prime},\alpha)$ (68)
+, for $\lambda<\lambda^{\prime}$. Thus the following limit exists:
+${{\bm{M}}}^{f}_{E^{\star}}(\alpha)=\lim_{\lambda\rightarrow 0^{+}}{{\bm{M}}}^{f}_{E^{\star}}(\lambda,\alpha).$ (69)
+We next show that for small enough $\lambda$, even with access to full side-information, the AMP iterates “visit” ${{\bm{M}}}^{f}_{E^{\star}}$
+Proposition D.5.
+Suppose that $E^{\star}\neq\phi$, $T^{\star}=\phi$ and $T^{\star}_{E^{\star}}=\phi$, and $\alpha>\alpha_{c}$.
+Then under side-information $S=\sqrt{\lambda}W^{\star}+\sqrt{1-\lambda}\mathbb{I}$, for any $\epsilon$, for any $\epsilon>0$ and small enough $\lambda$, there exists $t^{\prime}(\epsilon,\lambda)$ such that for
+$\norm{{{\bm{M}}}^{(t^{\prime}(\epsilon,\lambda))}-{{\bm{M}}}^{f}_{E^{\star}}(\alpha)}\leq\epsilon.$ (70)
+Proof.
+Let $U=E^{\star}$ and consider the linearized dynamics along the orthogonal complement $(U)_{\perp}$:
+$\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})\mathrel{\mathop{:}}=(\mathbf{I}+{{\bm{M}}}_{U})^{-1}\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U}){{\bm{M}}}_{\perp}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U})^{\top}\right]\,,$ (71)
+where ${{\bm{M}}}_{U}\subseteq E^{\star}$.
+By assumption, $\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})=0$ for ${{\bm{M}}}_{U}=0$.
+Therefore, by continuity of $\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})$ w.r.t ${{\bm{M}}}_{U}$, for any $\alpha>\alpha_{c}$, $\exists\delta>0$ such that if ${{\bm{M}}}_{U}\prec\delta\mathbf{I}_{p}$, $\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})$ remains contractive.
+For small enough $\delta$, we may therefore split the dynamics into two phases:
+  1. 1.
+    Phase 1: $0<t<t_{\delta}$ where $t_{\delta}=\inf_{t>0}{{\bm{M}}}^{t}_{U}\succ\delta\mathbf{I}_{p}$.
+  2. 2.
+    Phase 2: $t>t_{\delta}$.
+From Theorem 1, we know that $t_{\delta}=\mathcal{O}(\delta\log\frac{1}{\lambda})$.
+By the choice of $\delta$, we have by the end of phase $1$, ${{\bm{M}}}_{\perp}\prec C\lambda\mathbb{I}$ for some constant $C_{1}$.
+By linearization D.3, we obtain that for $t$ additional steps in Phase 2, the overlap along ${{\bm{M}}}_{\perp}$ remains bounded by $e^{Kt\lambda}$ for some constant $K>0$.
+We can further ensure that for small enough $\delta$ we can further ensure that ${{\bm{M}}}^{t_{\delta}}_{U}\prec C_{2}\delta\mathbf{I}_{p}$, for some constant $C_{2}$.
+Subsequently, by considering the limit $\delta\rightarrow 0$ with $\lambda<<\delta$, and the roles of $\lambda,\delta$ interchanged, we obtain that after $\mathcal{O}(t^{\prime}(\epsilon,\delta))$ additional steps, where $t^{\prime}(\epsilon,\delta)$ is defined as in Equation 67:
+$\norm{{{\bm{M}}}^{(t^{\prime}(\epsilon,\delta))}-{{\bm{M}}}^{f}_{E^{\star}}(\alpha)}\leq\epsilon$ (72)
+, since $t^{\prime}(\epsilon,\delta)$ is independent of $\lambda$, the overlap along ${{\bm{M}}}_{\perp}$ at $t^{\prime}(\epsilon,\delta)$ remains $\mathcal{O}(\lambda)$. ∎
+The above proposition showed that, by restricting the side-information to specific subspaces, we can uniquely characterize sequences of fixed points along sequences of growing subspaces.
+We’re now ready to state the detailed version of Proposition 5.1 along with its proof, which states that upon the inclusion of additional side-information along a new subspace:
+Proposition D.6 (Full-version of Proposition 5.1).
+Let $U\subseteq\mathbb{R}^{p}$ be a subspace such that $E^{\star}_{U}$ is non-empty..
+Consider AMP iterates with the Bayes-optimal choice of $f_{t},g_{t}$ with side-information restricted to the subspace $U$ for sufficiently small $\lambda>0$.
+Suppose that ${{\bm{M}}}_{U}$ is a fixed-point under such a dynamics where ${{\bm{M}}}_{U}$ spans $U$.
+Then, under the dynamics with an additional side-information along the subspace $E^{\star}_{U}$:
+  • For $\alpha>\alpha_{\text{gst}}({{\bm{M}}}_{U})$, AMP first reaches arbitrarily close to ${{\bm{M}}}_{U}$ in $\mathcal{O}(\log\frac{1}{\lambda})$ iterations and subsequently recovers $M^{*}_{M_{U}}$ in additional $\mathcal{O}(\log\frac{1}{\lambda})$ steps for arbitrarily small $\lambda$, where $M^{*}_{M_{U}}$ is as defined in eq.
+(5).
+  • For $\alpha<\alpha_{\text{gst}}({{\bm{M}}}_{U})$ and sufficiently small $\lambda$, AMP reaches arbitrarily close to ${{\bm{M}}}_{U}$ in $\mathcal{O}(\log\frac{1}{\lambda})$ and the overlap along $E^{\star}_{U}$ remains bounded by $C\lambda$.
+Proof.
+By assumption, $M_{U}$ is a fixed point of $F(M)$ under the presence of side-information restricted to $U$ Therefore, Equation 9 implies that ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}-\sqrt{M}_{U})$ almost surely lies in $U$ and thus $T^{\star}_{U}$ is empty.
+Therefore, by the proof of Proposition D.5, for sufficiently small $\lambda$, and any $\epsilon>0$ AMP iterates reach $\epsilon$-close to ${{\bm{M}}}_{U}$ in time $\mathcal{O}(\log\frac{1}{\lambda})$ iterations, while the overlap along the orthogonal complement to ${{\bm{M}}}_{U}$ being $\mathcal{O}(\lambda)$.
+Theorem 5.1 then follows that of 4.2 by considering the following linearized operator along $U^{\perp}$
+$\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})\mathrel{\mathop{:}}=(\mathbf{I}+{{\bm{M}}}_{U})^{-1}\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U}){{\bm{M}}}_{\perp}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U})^{\top}\right]\,,$ (73)
 where ${\bm{\omega}}=\sqrt{M}_{U}{\bm{\xi}}$ with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}})$.
-Similar to Lemma B.1, the linearization follows by noting that $y$ is independent of $\sqrt{{{\bm{M}}}_{\perp}}{\bm{\xi}}$ and $\sqrt{{{\bm{M}}}_{U}+{{\bm{M}}}_{\perp}}=\sqrt{{{\bm{M}}}_{U}}+\sqrt{{{\bm{M}}}_{\perp}}$ since ${{\bm{M}}}_{\perp}\perp{{\bm{M}}}_{U}$.
-Appendix C A list of examples of the ${\bm{g}}_{\rm out}$ function
+Similar to Lemma D.2, the linearization follows by noting that $y$ is independent of $\sqrt{{{\bm{M}}}_{\perp}}{\bm{\xi}}$ and $\sqrt{{{\bm{M}}}_{U}+{{\bm{M}}}_{\perp}}=\sqrt{{{\bm{M}}}_{U}}+\sqrt{{{\bm{M}}}_{\perp}}$ since ${{\bm{M}}}_{\perp}\perp{{\bm{M}}}_{U}$.
+∎
+D.7 Proof of Theorem 5.2
+Theorem 5.2 follows by inductively applying Propositions 5.1 and D.5 as we describe below:
+  1. 1.
+    First, suppose that $T^{\star}=\emptyset$ and $T^{\star}_{E^{\star}_{k}}=\emptyset$ for all $k\in\mathbb{N}$.
+By compactness, for any $k$, $\inf_{\mathbf{v}\in E^{\star}_{E^{\star}_{k}}}v^{\top}\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})\mathbf{v}>0$.
+Therefore, as in the proof of Theorem 1, for each $E^{\star}_{k}$, $\exists\alpha_{k}$ such that for $\alpha>\alpha_{k}$, the corresponding fixed-point along $E^{\star}_{k}$ becomes unstable along the full subspace $E^{\star}_{E^{\star}_{k}}$, leading to weak-recovery of $E^{\star}_{k+1}$ in $\mathcal{O}(\log\frac{1}{\lambda})$ additional steps.
+Since $k$ is bounded above by $p$, we obtain that $\alpha^{\star}={\rm max}_{k}\alpha_{k}$ is finite.
+For $\alpha>\alpha^{\star}$, Proposition D.5 implies that AMP iterates sequentially visit $E^{\star}_{k}$ and eventually recover $E^{\star}_{gst}$.
+  2. 2.
+    In the presence of trivial subspaces $T^{\star}$, as shown in the proof of Theorem 3.2 and explained in section C, the trivially-coupled subspaces are learned instantaneously.
+Therefore, the sequence of fixed-points traversed by AMP span the following sequence of subspaces:
+    1. (a)
+      $U^{\star}_{1}=T^{\star\star}_{1}$.
+    2. (b)
+      $U^{\star}_{j+1}=U^{\star}_{j}\cup E^{\star}_{U^{\star}_{j}}\cup T^{\star\star}_{E^{\star}_{U^{\star}_{j}}}$, for $j=1,2,\cdots$,
+    where $T^{\star\star}_{U}$ denotes the maximal chain of trivial subspaces starting from $U$:
+    $T^{\star\star}_{U}=T^{\star}_{U}\cup T^{\star}_{T^{\star}_{U}}\cup\cdots.$ (74)
+    It is easy to check that the set subspaces $U\in\mathbb{R}^{p}$ is totally ordered under the span of $U\cup T^{\star}_{U}\cup E^{\star}_{U}$ and therefore possesses a unique maximal element.
+Therefore, both sequences $U^{\star}_{i}$ defined above and $E^{\star}_{i}$ defined in Definition 6 terminate at the same maximal element $E^{\star}_{gst}$.
+Similar to $(i)$, proposition D.5 then implies that for large enough $\alpha$, and small enough $\lambda$ AMP traverses the sequence of subspaces $U^{\star}_{1},U^{\star}_{2},\cdots$, recovering $E^{\star}_{gst}$ in $\mathcal{O}(\log(\frac{1}{\lambda}))$ iterations.
+Appendix E From AMP thresholds to Computational phase transitions for First-order algorithms
+In this section, we discuss how Theorems 4.2 and 5.2 can be translated to rigorous lower-bounds on the sample complexity of weak-recovery from arbitrarily small initialization.
+For brevity, we omit the full-proof and provide a sketch of the argument basewd on existing results.
+We recall the main results of [Celentano et al., 2020]. [Celentano et al., 2020] consider the following class of general first-order algorithms (GFOM):
+${\bm{v}}^{t}$ $={\bm{X}}f^{(1)}_{t}({\bm{u}}^{1},\cdots,{\bm{u}}^{t},{\bm{y}},{\bm{u}})+f^{(2)}_{t}({\bm{v}}^{1},\cdots,{\bm{v}}^{t},{\bf y})$ (75)
+${\bf u}^{t+1}$ $={\bm{X}}^{T}g^{(1)}_{t}({\bm{v}}^{1},\cdots,{\bm{v}}^{t},{\bm{v}})+g^{(2)}_{t}({\bm{u}}^{1},\cdots,{\bm{u}}^{t},{\bm{y}},{\bm{u}}),$ (76)
+where ${\bm{v}}^{t}\in\mathbb{R}^{n\times r},{\bm{u}}^{t}\in\mathbb{R}^{n\times r}$.
+Note that Algorithm 1 is a special case of the above class.
+The variables ${\bm{u}},{\bm{v}}$ encode additional side-information.
+Subsequently, Celentano et al.
+[2020] show that any algorithm belonging to the above class can be mapped to a message-passing algorithm on an infinite-tree such that the errors of the GFOM algorithm at any finite-time $t$ asymptotically match the errors of the surrogate estimation problem on the tree.
+For the tree-inference problem, the optimality of the Bayes-AMP directly follows from AMP being the dense-limit of the information-theoretically optimal Belief propagation on trees.
+This translates to the optimality of Bayes-optimal AMP at any finite-time $t$ amongst GFOM algorithms.
+Similar to our Theorem 1, Celentano et al.
+[2020] provide thresholds $\alpha_{c}$ for weak-recovery starting from arbitrarily small initialization in the case of sparse Phase retrieval and sparse PCA.
+As we discussed in Section B, the Bayes-optimal AMP in the presence of such side-information corresponds to a modified prior.
+While Celentano et al.
+[2020] prove their results for the specific classes of high-dimensional regression (single-index) problems and low-rank matrix estimation, the underlying argument is general enough to be adapted to our setting of multi-index models.
+We therefore conjecture that the thresholds and class of problems described in Theorems 3.2,1,5.2 fundamentally characterize the learnability of multi-index models through first-order algorithms.
+Appendix F Explicit results for some examples of the ${\bm{g}}_{\rm out}$ function
 In this section, we list several problems, their corresponding ${\bm{g}}_{\rm out}$, and analysis of their weak recoverability.
-C.1 $g(z_{1},z_{2},....,z_{p})=\prod_{j=1}^{p}z_{j}$
+F.1 $g(z_{1},z_{2},....,z_{p})=\prod_{j=1}^{p}z_{j}$
 For all $p$, ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=0$. Otherwise we have two cases We have two cases. If $p=2$:
@@
 The matrix $\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})$ is diagonal, with identical diagonal elements equal to $f(y)$:
-$f(y)=2G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0\end{array}\right)/G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0,1\end{array}\right)-1$ (52)
+$f(y)=2G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0,1\end{array}\right)/G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0\end{array}\right)-1$ (80)
 The alpha critical $\alpha_{c}$ will thus be
 $\alpha_{c}=\left[\int_{-\infty}^{\infty}\,{\rm d}yf(y)^{2}\mathcal{Z}_{\rm out}(y)\right]^{-1}$ (81)
-C.2 $\frac{1}{p}\sum_{i=1}^{p}z_{i}^{2}$
+F.2 $\frac{1}{p}\sum_{i=1}^{p}z_{i}^{2}$
 The function is even, so ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=0$.
@@
 Of course this procedure fail at the first step if $p=1$. On the other hand, one can check that the final result is still correct for all positive integer $p$.
-C.3 $g(z_{1},\ldots,z_{p})=\operatorname{sign}(\prod_{j=1}^{p}z_{j})$
+F.3 $g(z_{1},\ldots,z_{p})=\operatorname{sign}(\prod_{j=1}^{p}z_{j})$
 We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
@@
 We can now focus on studying the stability
-$Z_{\rm out}(y,\bm{0},{\bm{I}})=\frac{1}{(2\pi)^{p/2}}\int{\rm d}{\bm{z}}e^{-\nicefrac{{{\bm{z}}^{\top}{\bm{z}}}}{{2}}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}=\frac{1}{2}$ (62)
-In order to compute $\nabla_{\omega}{\bm{g}}_{\rm out}(y,\bm{0},{\bm{I}})$, we note at first that
+$Z_{\rm out}(y,\mathbf{0},{\bm{I}})=\frac{1}{(2\pi)^{p/2}}\int{\rm d}{\bm{z}}e^{-\nicefrac{{{\bm{z}}^{\top}{\bm{z}}}}{{2}}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}=\frac{1}{2}$ (90)
+In order to compute $\nabla_{\omega}{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}})$, we note at first that
 $\frac{1}{(2\pi)^{p/2}}\int{\rm d}{\bm{z}}(z_{k}^{2}-1)e^{-\nicefrac{{{\bm{z}}^{\top}{\bm{z}}}}{{2}}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}=0,\quad\forall k\in\{1,\ldots,p\}.$ (91)
@@
 Using (16) we obtain that $\alpha_{c}$ is the inverse of the largest eigenvalues of
-$\frac{4}{\pi^{2}}\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix}$ (67)
+$\frac{4}{\pi^{2}}\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix}$ (95)
 So $\alpha_{c}=\pi^{2}/4$. For $p\geq 3$, defining (±)$\mathbb{R}^{p}_{(\pm)}=\{{\bm{z}}\in\mathbb{R}^{p}|\operatorname{sign}(z_{1}\ldots z_{p})=\pm 1\}$, we need to compute integrals of the type
@@
 This model cannot be learned with $n=\mathcal{O}(d)$ samples for $p\geq 3$.
-C.4 $g(z_{1},\ldots,z_{p})=z_{1}^{2}+\operatorname{sign}\left(\prod_{j=1}^{p}z_{j}\right)$
+F.4 $g(z_{1},\ldots,z_{p})=z_{1}^{2}+\operatorname{sign}\left(\prod_{j=1}^{p}z_{j}\right)$
 We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
@@
 The result is dependent on a number of integrals which we can compute analytically. First we have
-$Z_{\rm out}(y,\bm{0},{\bm{I}})$ $=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (73)
+$Z_{\rm out}(y,\mathbf{0},{\bm{I}})$ $=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (101)
 $=\frac{1}{2(2\pi)^{\nicefrac{{p}}{{2}}}}\sum_{s_{1},\ldots,s_{p}=\pm 1}\frac{\bm{1}_{y>s}}{\sqrt{y-\prod_{j}s_{j}}}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}e^{-\frac{y-s_{1}\ldots s_{p}}{2}-\frac{z_{2}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (102)
@@
 $\alpha_{c}=\left[\int\,{\rm d}yC(y)^{2}Z_{\rm out}(y)\right]^{-1}\approx 0.575166$ (114)
-C.5 $g(z_{1},\ldots,z_{p})=\sum_{j=1}^{p}\operatorname{sign}(z_{p})$
+F.5 $g(z_{1},\ldots,z_{p})=\sum_{j=1}^{p}\operatorname{sign}(z_{p})$
+Figure 3: Specialisation as an example of grand staircase with $g\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$.
+The direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is first learned for any $\alpha>0$.
+The remaining ones are only learned at the specialisation transition occurring at $\alpha\!\approx\!4.3$.
+We take care of the symmetries as indicated in S.I.
+G.
+Crosses denote AMP runs with $d\!=\!100$ averaged over $72$ seeds.
+The overlap is shown as a function of the sample complexity.
+Because of the symmetry the elements of $M$ can take one of two values: one on the diagonal and one outside of it.
+We display $(M_{11}+M_{22}+M_{33})/3$ in blue and $(M_{12}+M_{13}+M_{23})/3$ in red.
 We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
@@
 $\left({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\right)_{k}=\frac{1}{Z_{\rm out}}\sum_{a=1}^{p}\frac{\delta_{y,2a-p}}{2}\sum_{\bm{s}\in S_{a}}s_{k}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\delta(z_{k}-0)$
-Appendix D Further numerical observations and details
+Appendix G Further numerical observations
+Figure 4: Trajectories of $16$ AMP runs with random initialisation at $\alpha=4$ for $g(z_{1},z_{2})={\rm sign}(z_{1}z_{2})$.
+The shaded areas represent the error on the mean.
+We can see that as the dimension increases the algorithm gets increasingly slower
 Here we give more details about the numerical implementation of State Evolution and AMP.
 Both approaches require to compute ${\bm{g}}_{\rm out}$.
-All the examples we implemented are detailed in Appendix C.
+All the examples we implemented are detailed in S.I sec.
+F.
 The integrals are doing with the quadrature package in Scipy.
@@
 We do these integrals using a simple Monte Carlo approach.
-For Figures 1 and 2 we used a total $72000$ samples, for Figure 4 $7200$.
+For Figure 1 we used a total $72000$ samples, for Figure 3 $7200$.
 Computing such integrals is the numerical bottleneck.
@@
 We can see how AMP has a saddle-to-saddle dynamic, where the algorithm alternates plateaus for $\mathcal{O}(\log{d})$ ) iterations to fast drops in generalisation error, which are associated with new directions being learned.
+We probe experimentally the dependence on size in Figure 4.
+All the error bars present represent the error on the mean.
+We used the local computing cluster for our experiments.
+In particular we needed around $40$ CPUs with $72$ cores each and approximately $3$ Gb per core.
 As stated in the main text, models non-trivial subspaces are associated with symmetries.